A Bibliography of Publications in Theoretical Computer Science: 1990–1994

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

30 October 2017 Version 1.16

Title word cross-reference

#P [553]. #PH [553].

(n − 2) [1095]. (p, q) [458]. (x_{i} + y_{j}) [595]. ∗ [996]. 0 [1160, 890, 1017]. 1 [826, 800, 1011, 128, 646, 553]. 12 [931]. 2 [744, 1095, 1043, 18, 16, 17, 414, 1097, 793, 224]. 2 × 2 [587]. 3 [526]. ^{0} [161]. A^{2}CCS [53]. AC^{0} [758]. α_{0} [328, 186]. B [1178, 1178]. C [664]. CCS [53]. co − NL_{e} [310]. Δ_{2}^{P} [337]. F [1179, 1035]. G(u)(u^{n}) [306]. H_{2} [446]. h_{m} [937]. ∞ [598]. k [520, 818, 274, 700, 1097, 891]. K < A > [841]. L_{e} [310]. λ [719, 716, 904, 1036, 909, 6, 360, 910, 1190]. ⌊N + V_{k}, V_{l}⌋ [647]. log log n [82]. log n [82]. μ [996]. N [145, 487, 850]. N^{k} [428]. NL_{e} [310]. ≠ [1120]. n × n [946]. ν [254]. ν_{1} [186]. O(n^{2}) [595]. ω [706]. p [506, 283, 267]. P_{1}^{1} [869]. P_{e} [310]. π [1165]. Π_{2}^{P} [857, 777]. Π_{2} [945]. q [834, 693, 412]. R [1120]. s − 1 − 1 [887]. Σ_{2}^{p} [939]. T [945, 810, 506]. ↑ [825]. v_{n} [937]. X [706]. x^{m}y^{n} = z^{p} [409]. z [684].

* [164].

-automata [706]. -automatic [693]. -calculi [716]. -calculus [1165, 996, 719, 904, 1036, 909, 6, 360, 910, 254]. -codes [684]. -coloring [744]. -competitive [526]. -complete [857, 777, 337]. -completely [283]. -connected [700]. -creative [283]. -degrees [506]. -element [487]. -expressions [664]. -functions [810]. -grams [412]. -inkdot [1011]. -isomorphism [267]. -modules [841].

-order [807]. -products [328, 186]. -property [1095]. -Regular [598, 520]. -resolution [818]. -semantics [1035]. -separated [891]. -series [834, 834]. -sparse [274]. -structure [1097]. -structures [1095, 1043, 18, 16, 17, 414, 1097, 810]. -tape [128]. -truth-table-hard [800]. -Turing [553]. -type [458]. -universality [869]. -words [706].

10th [924]. 13th [46]. 1988 [99]. 1990 [400]. 1992 [913].

2 [982]. 2-D [982]. 2D [1152]. 2DPDA [527]. 2nd [26, 476].

3 [392]. 3rd [767].

5th [858].

6th [300].

’88 [46].

’91 [858].

= [1114].

A. [1200]. abduction [922]. abductive [880]. abelian [1022, 210]. absence [327]. absolute [7]. absorbing [878]. Abstract [218, 133, 613, 207, 483, 528, 958, 902, 920, 245, 1156, 1185]. abstraction [1178]. abstractions [214]. ACC [161]. accepted [260, 292]. acceptors [982]. access [661, 546]. accessibility [528]. Accumulators [214]. action [135, 775]. actions [847, 53, 766, 765]. actor [702]. Acyclic [303, 411]. adaptive [227]. Adding [67, 915]. addition [1116, 521]. Additions [963]. Additive [877]. addresses [1105]. Adequacy [1189, 443, 1188]. Adequacy-preserving [443]. adjacency [309]. adversary [692]. advice [540]. Affine [816]. after [976]. agent [864]. Aggregate [1130, 854]. aggregation [806]. ahead [422]. Algebra [46, 306, 960, 92, 373, 1008, 174, 135, 631, 369, 798, 167, 604, 562, 518, 866]. algebra-based [135]. Algebraic [131, 821, 805, 92, 509, 1182, 311, 703, 287, 380, 316, 376, 352, 248, 271, 592, 593, 519, 134, 330, 612, 177, 13, 1049, 1172, 865, 209, 877, 1156, 729, 185, 396, 323]. algebraically [139]. Algebras [1, 852, 674, 1137, 34, 52, 788, 175, 838, 439, 908, 441, 557, 533, 76, 316, 10, 764]. Algèbre [46]. algèbres [838]. algorithm [1088, 402, 1109, 1136, 338, 1158, 932, 581, 1162, 93, 226, 969, 612, 545, 649, 1046, 407, 1027, 408, 981, 236, 684, 989, 937, 96, 32, 342, 552, 686, 1194, 892, 795, 801, 1052]. algorithme [795]. Algorithmic [835]. Algorithms [1112, 390, 752, 688, 744, 306, 551, 572, 92, 889, 250, 1108, 1006, 420, 896, 815, 1059, 166, 51, 78, 406, 432, 1016, 23, 430, 1118, 469, 386, 98, 471]. allocation [972]. allowed [498]. Almost [791, 1197, 887, 1195, 202]. Almost-everywhere [791]. ALOHA [686]. alphabet [210, 263, 1079, 1091]. Alphabetic [535, 434]. alphabets [403, 71, 467]. also [82]. Alternating [504, 351, 754, 694, 1102, 542, 538, 350, 1011, 982, 492]. alternation [1155]. alternative [27, 911]. AMAST [131]. ambiguity [353]. among [95]. Analog [1080]. Analogical [762]. Analogy [760]. analyse [451]. analyses [555]. Analysis [146, 51, 365, 752, 302, 92, 829, 902, 642, 166, 451, 974, 903, 370, 286, 430, 1118, 1032, 523]. analytic [1068]. Andersson [653]. angels [561]. Angular [414]. animals [1020, 827]. Animaux [827]. annealing [534]. answering [144]. anti [840]. anti-excédances [840]. anti-exceedance [840]. antirandomization [1083]. apart [1039]. aperiodicity [340]. application [930, 339, 866, 523]. applications [784, 528, 271, 593, 519, 1115, 843, 750, 735]. approach [1170, 620, 860, 204, 701, 1026, 493, 1129, 424, 1187, 366, 1049, 1172, 171, 703, 1121, 295, 734]. approaches [29]. Approximate [412, 206, 988, 194, 407, 555, 142]. Approximating [585, 1157]. Approximation [169, 446, 643, 78, 662]. APS [865]. arbitrary [1158, 503]. Arbres [46, 827, 835]. Area [235, 507]. arising [814]. arithmetic [435, 58, 468, 229, 741]. arithmetical [181]. Arrangements [420]. array [206, 285, 333]. arrays [338]. Artin [453]. ascent [622]. Askey [836]. Askey-Gasper [836]. Aspects [509, 24]. assertions [885]. assignment [325, 1035]. associated [826, 951, 521]. associative [644]. associativity [611]. associée [826]. assuming [346]. assumptions [915]. asymétrie [823]. asymmetry [823]. Asymptotic [1162, 974, 588]. asymptotics [944]. Asynchronous [1096, 1024, 83]. ATMs [128]. Atomic [942, 53]. Attribute [515, 31]. Attributed [859]. Auto [833, 833]. Auto-similarité [833]. Auto-similarity [833]. autoepistemic [605]. Automata [84, 168, 1000, 488, 1133, 201, 431, 1173, 888, 792, 291, 1007, 816, 281, 1152, 706, 328, 260, 677, 186, 707, 986, 575, 160, 42, 190, 694, 793, 982, 1150, 678, 1015, 680, 1117, 261, 289, 870, 209, 877, 566, 266, 504, 349, 126, 1096, 1191, 411, 1100, 382, 1005, 275, 998, 353, 1145, 354, 539, 471, 413]. automata-like [354]. Automatic [166, 445, 30, 693, 397]. Automatizing [1174]. automaton [340, 940, 596, 801]. automaton-definable [596]. automorphism [471]. availability [977]. Average [752, 402, 1109, 829, 932, 166, 59, 410, 686, 850]. Average-case [752, 829, 166, 410]. AVL [1014, 55]. axiom [607]. axiomatic [244]. axiomatization [576, 849]. axiomatizations [852]. axioms [376, 771]. ayant [931].

B [1200, 363]. B-rational [363]. backtracking [1104]. bad [1195]. bagdomain [1180]. Balance [1014, 652]. balanced [829]. balancing [1058]. Barrington [656]. bars [1154]. base [145]. based [64, 218, 149, 889, 135, 762, 955, 917, 253, 1029, 341, 475]. Bases [838, 175, 345, 1074, 841, 1141, 867, 1195]. Basic [237, 697, 901, 16, 143, 985]. BCK [665]. BCK-lambda-terms [665]. be [89, 257]. beating [850]. behaved [498]. behavior [527]. behaviour [961]. behavioural [24]. behaviours [604, 371]. Bendix [736]. beta [220]. beta-cycles [220]. between [70, 192, 624, 330, 605, 1054, 379, 1003, 1144]. beyond [619]. Bezout [1122]. Bi [69, 284, 1197]. Bi-immunity [69]. bicartesian [1041]. bidirectional [59]. Bieberbach [836]. biinfinite [311]. bijection [1144]. bilinear [306]. Binary [584, 1198, 260, 114, 286, 965]. binding [370, 735]. binomial [948]. biography [897]. bipartite [1016, 797]. Birkhoff [441]. bishop [490]. bisimilarity [939, 1085]. Bisimulation [775, 580]. bisimulations [771]. bit [226]. bit-strings [226]. Blair [500]. block [797]. blocking [976]. blocking-after-service [976]. board [946]. Bohm [897, 903]. Boolean [1201, 993, 94, 277, 1135, 187, 507, 217, 586, 568, 819, 1195]. Borel [749]. born [392]. Bottleneck [451]. Bottom [958, 1007, 635, 92, 850]. Bottom-up [958, 1007, 635, 92]. BOTTOM-UP-HEAPSORT [850]. bound [216, 932, 1199, 55]. boundary [738]. Bounded [277, 493, 280, 270, 832, 331, 581, 538, 617, 1092, 227, 935, 196, 492, 750, 1123, 560]. boundedness [357, 560]. Bounds [1143, 1064, 94, 1106, 1115, 1063, 507, 949, 255, 524, 73, 1146]. boxes [201]. Boyer [402, 408]. brackets [329]. branching [479]. Branges [836]. breadth [291]. breadth-first [291]. breakdowns [975]. bridge [379]. broadcast [968]. brother [1014]. Buchi [525]. Buchi-McNaughton [525]. Burstall [885]. Byzantine [466].

C [378]. C-MOS [378]. CAAP [46]. cacti [276]. calculi [669, 728, 716, 1185]. calculus [899, 664, 1165, 996, 781, 1089, 719, 904, 1187, 184, 1036, 908, 909, 441, 325, 100, 6, 360, 910, 423, 668, 361, 607, 254, 760]. call [1036]. call-by-name [1036]. can [319, 89]. Cancellation [1010]. cancellative [416]. cannot [257]. Canonical [1044, 1053, 406, 536]. Cantor [224]. Capturing [565, 668]. Carlo [934]. carrier [441]. cartes [826]. Cartesian [3, 62, 13, 1184]. case [752, 948, 550, 829, 166, 410]. Catalan [936]. Categorical [5, 1185, 728, 705]. categories [3, 786, 663, 13, 183, 1184, 396]. Category [135, 730, 62, 11]. Category-sorted [135]. category-theoretic [11]. catenative [259]. Causal [575, 847, 961]. caused [62]. Cayley [677, 1100]. CCS [108, 761]. cellular [1152, 707, 1015, 877, 940, 566, 1096, 1100]. certain [896, 486]. CFL [954]. CFL’s [820]. chain [219]. change [949]. change-making [949]. channels [1022, 756]. characterisation [1030, 12, 856]. characterization [343, 719, 542, 11, 37, 479, 333, 288]. characterizations [265, 737, 197]. characterize [571]. characterized [89]. Characterizing [547, 560]. characters [843]. CHARM [920]. cheatable [69]. check [25]. checking [993, 302, 481, 598, 359, 361, 254]. checks [610]. chemical [483]. chemins [837]. Chern [843]. chip [744, 417, 1199]. choice [571, 772, 581, 531, 607]. Church [184, 423]. Circuit [659, 378, 819]. circuits [556, 435, 507, 658, 568, 313, 1104, 229, 290, 984]. circular [406]. circumscription [857, 777]. Clans [1043, 16]. class [655, 883, 1148, 260, 32, 749]. classes [540, 351, 590, 620, 234, 1113, 648, 548, 200, 547, 565, 265, 419, 95, 297, 72, 310, 739, 843]. classical [582]. Classification [826, 306, 41, 1007, 200, 587, 826]. classifiers [1184]. clause [460, 358, 574]. clauses [498, 1037, 969, 880, 103]. clique [386]. cliques [1144]. Clock [1024]. Clocks [442, 461]. closed [978, 3, 857, 777, 62, 13, 1184, 915, 807, 122]. closed-form [807]. closed-world [857, 777]. closing [192]. Closure [238, 894, 163, 260, 187, 124]. closures [308]. cms [955]. cms-based [955]. co [637, 964]. co- [637]. co-induction [964]. coalgebras [780, 963]. coarsest [1046]. Code [66, 219, 189]. Codes [1142, 44, 97, 270, 983, 487, 1196, 679, 684, 280, 1143, 681]. coding [453]. codings [1153]. coefficient [948]. coefficients [306, 717, 513]. Coexistence [938]. Coherence [299, 783]. cohesive [1196]. coinductive [918]. col [451]. collapses [577]. collection [1069]. collision [686]. Colloque [46]. Colloquium [671, 1131, 46]. colonies [1140]. coloring [744, 645, 1063, 93, 346]. colors [93]. colouring [36, 545]. combination [833]. combinational [451, 378]. combinatoire [833, 451]. combinatoires [823]. Combinatorial [400, 842, 1089, 823, 171, 295, 720, 874, 1193, 399]. combinatorics [515, 420, 821, 671, 1131]. combinators [779, 327, 912]. Combinatory [49, 907, 1178, 558, 6, 499]. Combining [561, 34]. Comments [652]. commitment [213]. common [401, 689, 691, 980]. communicating [1151]. Communication [19, 991, 1009, 552]. commutation [517]. commutations [675]. commutative [673, 491, 308, 405, 159, 167, 838, 1141, 76, 541, 513]. commutatives [838]. commutativity [611, 990]. compact [536]. compaction [64, 309]. compaction-based [64]. Comparative [101]. Comparing [728, 120, 574]. comparison [262, 330, 711, 955, 469]. comparisons [595]. Compatibility [134]. compatible [711]. Competitive [1059, 1064, 526, 1061, 1027]. Compilation [26, 858, 27, 29, 318, 1074]. compilers [380]. complementation [1091]. complements [143]. Complete [852, 452, 363, 727, 130, 579, 1098, 883, 960, 1134, 1198, 1012, 1203, 777, 896, 200, 864, 104, 486, 946, 512, 391, 1079, 114, 1169, 917, 337, 340, 857, 211, 292, 691]. completely [1143, 283]. Completeness [88, 498, 610, 303, 50, 11, 441, 325, 6, 853, 181, 502]. Completing [252]. Completion [294, 356, 52, 611, 736]. Completions [529]. complex [816]. Complexité [835]. complexities [82, 985]. Complexity [714, 644, 72, 641, 1065, 1087, 189, 1122, 19, 791, 94, 352, 556, 351, 278, 312, 872, 645, 115, 620, 70, 991, 1037, 806, 740, 1113, 648, 1101, 107, 231, 685, 862, 158, 835, 1152, 1111, 754, 339, 547, 565, 1115, 22, 657, 265, 419, 1176, 507, 1009, 486, 1085, 128, 879, 297, 1116, 23, 804, 258, 345, 59, 217, 794, 973, 255, 660, 1119, 223, 567, 980, 586, 504, 235, 205, 469, 651, 1048, 552, 98, 290, 224, 984, 1042, 1112]. Composition [887, 769, 1072]. compositional [914, 318, 576, 336, 1028, 1029]. Compositionality [476, 767]. compounds [874]. Comprehension [663]. Compression [1203, 1012, 848, 413]. COMPUGRAPH [697]. Comput [1201, 1203, 1202]. Computability [1093, 813, 759, 493, 867, 140]. Computable [518, 73, 287, 759]. Computation [172, 474, 470, 435, 29, 367, 368, 369, 57, 613, 1119, 1065, 1080, 396]. Computational [723, 901, 1188, 1090, 258, 973, 31, 1037, 862, 158, 1021, 660, 223, 741, 224, 199]. computationally [269]. Computations [278, 865, 401, 1008, 338, 1045, 329, 102, 1026, 21, 702, 196, 233]. Computer [99, 858, 924, 509, 913, 92, 364, 379, 1194]. Computer-Theoretic [509]. Computing [695, 697, 1116, 790, 306, 556, 312, 991, 406, 127, 471]. concatenation [389, 753, 1079]. concavity [404]. concept [1129, 9]. concepts [143, 881]. Conceptual [459]. concerning [982, 616]. Concrete [905, 906, 572]. concretization [366]. Concurrency [476, 767, 1173, 733, 120, 480, 761, 459, 372]. Concurrent [485, 101, 956, 373, 920, 102, 627, 136, 454, 100, 704, 195, 919, 869, 523, 301]. condition [1077]. Conditional [104, 480, 105, 577]. conditions [429, 332]. Conduction [326]. cone [472]. Conference [131, 99, 300, 913, 821]. configurations [748]. Confluence [1002]. Confluent [640, 675, 757, 921]. congruence [124]. Congruences [1196, 1160, 12]. congruential [951, 801]. conjecture [927, 836, 1047, 1110, 467, 616, 267]. conjugate [778]. conjunctive [1148, 217, 766]. connected [700, 1023, 892, 471]. Connectedness [374, 219, 1156]. connection [1007, 602]. connections [1183]. connectivity [1062]. conservative [377]. conserves [248]. Constant [269, 1169, 988, 128, 878, 895]. Constant-only [1169]. Constant-round [269]. constants [197]. constraint [923, 113, 627, 625, 496, 704]. constraints [971, 384, 1157, 118, 44, 807]. Constructing [1061, 536, 801]. Construction [147, 1024, 376, 743, 183, 968, 380, 957, 26, 841]. Constructions [841, 141, 812, 607, 323]. Constructive [899, 716, 30, 583, 450, 379]. constructively [919]. constructors [601]. Context [830, 312, 222, 155, 740, 187, 437, 939, 542, 1110, 1010, 812, 236, 987, 256, 524, 472, 121, 521, 347, 12, 341, 475, 619]. Context-free [830, 312, 222, 155, 740, 187, 437, 939, 542, 1110, 1010, 812, 236, 987, 256, 472, 121, 347, 12, 341, 475, 619]. context-freeness [521]. Continuations [296]. Continuous [1112, 168, 258, 776]. Contractions [120, 947]. Control [318, 600, 459]. controlled [344, 686, 33]. convergence [986, 1118]. convex [1121]. convexity [404]. convincing [269]. convolution [468]. Conway [489]. cooperating [1009, 522]. cooperation [221]. coprimality [641]. Corrado [897]. corrections [963]. Correctness [195, 1038, 667]. Corrigendum [1202, 954]. cosemicomputable [557]. cost [1105, 1109, 1014, 998]. COSY [443, 86]. counter [737, 694, 123]. counterpart [1183]. Counting [590, 203, 673, 655, 739, 205, 997, 555, 1121]. Coupling [717]. cover [797]. covering [276]. covers [465]. cpo [1170, 955]. cpo-based [955]. CPOs [13]. CRCW [688, 953]. CRCW-PRAM [688]. creative [283]. CRI [686]. CRI-length [686]. criteria [260]. criterion [436]. cryptography [339]. cryptosystems [814]. CSP [955]. CTL [996]. cubic [700]. CUCH [911]. curved [738]. curves [738, 420, 453]. Cut [397, 886, 28, 1042]. cut-formulas [1042]. cycle [1087]. cycles [220, 38]. Cyclic [15, 925, 978, 772].

D [905, 1160, 890, 982]. D. [430]. D0L [225, 257]. Dactl [54]. Daisies [489]. d’algorithmes [835]. dans [833, 447, 840]. Data [474, 1203, 322, 1062, 138, 632, 413, 960, 572, 38, 1012, 34, 272, 393, 653, 804, 630, 533, 287, 121, 867, 615, 139]. Data-movement-intensive [474]. Database [148, 149, 220, 1128, 809, 718, 116, 560]. databases [387, 163, 156, 1125, 1129, 330, 151, 425, 808]. dataflow [442, 375, 957]. Datalog [152, 115, 119, 806, 807]. Days [392]. DCP [320]. Decidability [994, 80, 232, 97, 20, 1004, 1139, 72, 541]. Decidable [357, 1084, 1054, 521, 639, 197]. decide [581, 530, 720]. Deciding [939, 742, 1085, 32]. decipherability [1136]. deciphering [270]. decision [463, 648, 519, 657, 987, 33, 1151]. declarative [851, 103, 574]. Decoders [314]. decomposable [699]. decomposition [550, 889, 119, 670, 489]. Decompositions [681]. decreasing [1111, 1002]. Dedekind [708]. deduction [631, 917]. Deductive [148, 156, 1125, 718, 425]. default [150, 605]. definability [192]. definable [995, 343, 908, 596]. define [620]. defined [1099, 1013, 1174, 428, 964, 638]. Defining [577, 919]. definite [851, 1026]. definitions [240, 547, 321, 1071, 918, 569]. Deforestation [68]. degree [309, 125, 353, 354]. degrees [945, 497, 158, 557, 506, 503]. Dejean [467]. Delaunay [743]. delay [270, 554, 280]. deletion [990]. Delta [93, 454]. demons [561]. Denotational [1170, 28, 7, 957, 574]. density [1111]. dependable [1031]. dependencies [560]. dependency [928, 663, 207]. dependent [183]. Depth [1042, 941, 1090, 229, 616, 290]. derandomization [1083]. derivation [378, 628, 1004, 762]. derivations [701, 798]. derivatives [687]. Derived [256]. Deriving [889]. derivors [66]. describe [297]. Descriptional [740]. design [133, 253, 362, 139]. designs [378]. d’espèces [831]. detection [871]. Determining [978, 234, 390]. Determinism [178, 324, 1155]. Deterministic [281, 642, 1102, 85, 152, 351, 110, 200, 1149, 511, 820, 954, 411, 342, 12, 341, 475, 82]. d’Euclide [795]. Development [858, 377, 134, 450, 67, 258, 454, 365, 1127]. developments [22, 345]. Deviation [1115]. DFA [793]. diagrams [147, 1002]. dictionary [1078]. difference [70]. differences [893]. different [220]. differentiable [991]. Differential [519, 172, 830, 49, 1115]. Differential-algebraic [519]. digital [944, 288]. digraphs [968, 876]. dimension [569]. dimensional [784, 1108, 1093, 981]. dinaturality [783, 856]. Diophantine [96]. dipping [948]. direct [463]. directed [48, 1020, 139]. director [668]. discipline [579]. discrepancy [930]. disjoint [994, 276]. disjointness [651]. disjunctive [304, 425]. disk [1032]. dissolution [1068]. distance [42, 209, 1145]. distances [497]. distinct [549]. distinguishing [226]. Distributed [554, 1064, 466, 216, 717, 701, 1026, 83, 59, 1127, 552, 1029, 892, 362]. Distribution [126, 978, 840]. distributional [651]. distributions [315, 1082, 974]. distributive [1001]. divergence [736]. division [778, 1175, 546, 257]. divisor [416]. domain [786, 1123, 323]. Domains [440, 8, 1173, 905, 594, 863, 110, 725, 1182, 724, 1181, 183, 906, 439, 726, 964, 372]. dominating [386]. Domination [1103]. Domineering [875]. domino [1076, 523]. dot [941, 616]. dot-depth [941, 616]. Double [948]. down [1149, 422]. Downward [112]. Drawing [1081]. droite [841]. DT [1017]. duality [922]. d’une [544]. Duval [1109]. Dyck [837]. Dynamic [48, 1057, 815, 404, 1078, 430, 81, 876, 1137, 51, 169, 649, 394, 118, 1052]. dynamical [1093]. Dynamics [382]. d’Yves [927].

E0L [333, 232]. E0L-E0L [333]. Earley [622]. easy [900, 836, 87, 90]. ECTL [996]. Edge [750, 745, 36, 93, 1062]. edge-colouring [36]. Editorial [589, 751, 47, 563, 570, 715, 132, 732]. Effective [550, 141, 882, 318]. effects [1045]. efficacy [741]. efficiency [199]. Efficient [688, 871, 144, 812, 809, 386, 539, 635, 388, 1158, 156, 896, 1128, 93, 187, 1046, 23]. efficiently [226]. election [59]. ELECTRE [495]. element [487]. Elementary [478, 462, 836, 1176, 679, 910]. elements [872, 931, 931]. eliminate [68]. elimination [1158, 397]. else [37]. embedding [344, 643]. embeddings [1198, 506]. empty [441, 873]. eNCE [891]. encodings [314, 1136]. end [1055]. engendrant [929]. enriching [736]. enrichment [1010]. entailment [923]. enumerable [221, 193]. Enumeration [832, 173, 445]. environment [245]. environments [67]. equal [82]. Equality [880, 112, 922, 588]. equation [936, 409, 127, 936]. Equational [52, 631, 137, 664, 784, 629, 105, 104, 441]. Equations [941, 678, 597, 951, 861, 786, 1115, 1094, 1079, 1091, 96, 813, 316]. Equivalence [702, 349, 626, 573, 20, 320, 160, 678, 106, 630, 1003, 32, 382, 232, 383, 178, 541, 1003]. Equivalence-preserving [626]. equivalent [435]. eraser [310]. ergodic [938]. Erratum [1201, 1203, 857]. error [1118]. errors [74, 1104, 632]. ESPRIT [697]. Essence [368]. estimation [415]. Euclid [795]. Euclidean [988]. Evaluation [170, 367, 854, 807]. evaluations [699]. Event [398, 926, 847, 479, 178]. events [1127]. every [236, 1197]. everywhere [791]. evolving [57]. Exact [976, 59]. example [171]. examples [710, 167, 912, 618]. excédances [840]. exceedance [840]. exception [198]. exceptional [1106]. exceptions [631, 632]. exchange [588]. exclusion [1034]. execution [717]. Executions [395]. existence [239, 223, 313]. existent [912]. exit [978]. expansions [171]. expected [552]. expedient [85]. Experimenting [573]. Experiments [190]. explicit [965]. Exponential [799, 1146, 271, 542, 453, 1197]. Exponential-time [799, 542]. Exponentiation [1053]. exposition [614]. expressibility [115]. expression [151]. expressions [664, 1134, 888, 485, 125, 391, 354, 503]. expressive [264, 247, 421, 773]. Extended [125, 354, 503, 772, 857, 777, 1147]. Extending [460]. Extension [499, 1088, 979, 452, 423]. extensional [393]. Extensions [656, 152, 377, 506]. exterior [281, 843]. External [652, 682, 1032]. extraction [1167]. extremal [986]. extreme [727].

facilities [67]. factor [828, 413]. Factorization [45]. Factorizations [60, 1144]. factorizing [1006]. factors [1044, 673, 585, 35, 1117, 228]. failure [1078, 574]. failures [102]. Fair [1177]. Fairness [497, 188, 243, 727, 667, 532]. families [829, 17]. fan [290]. fan-in [290]. Fast [206, 401, 643, 61, 820, 954, 410, 1136, 801]. Fatou [437]. fault [1028, 1029, 868]. faults [1033, 1030]. Feasibility [846]. feature [923]. fewness [590]. Fibonacci [355]. field [741]. fields [274, 49]. fifo [756, 291]. Fifth [99, 913]. Filter [464]. Final [557]. finding [216, 38, 689, 981]. Fine [465]. Finite [926, 340, 803, 102, 961, 1150, 947, 998, 197, 1145, 1073, 1133, 278, 888, 792, 274, 676, 711, 160, 42, 285, 793, 982, 678, 679, 1049, 1117, 289, 209, 742, 1142, 349, 997, 1143, 1048, 279, 867, 45, 741, 525, 353]. Finite-automaton [340]. finite-field [741]. Finite-memory [1150]. Finite-model [803]. finite-state [285, 1048]. Finite-valued [1145]. Finitely [284, 421]. Finiteness [332, 1077]. firing [417, 388, 1199, 1019]. First [1168, 928, 291, 293, 79, 916, 337, 626]. First-order [1168, 928, 293, 916, 626]. FIX [1188]. FIX-logic [1188]. Fixed [3, 10, 315, 145, 1017, 393, 286, 327, 544, 5, 856, 473, 316, 1055]. fixed-point [393, 286]. fixes [544]. fixpoint [1070, 627, 1038]. fixpoints [62]. Flajolet [555]. flat [330, 103]. flow [113, 621, 393, 114]. flowchart [528]. fold [106, 626, 305]. folk [474]. following [1118]. fonctionnelle [826]. forest [432]. forests [45]. Foreword [802, 400, 63, 672, 609, 722, 40, 654]. form [974, 406, 940, 807]. Formal [898, 821, 1099, 509, 825, 462, 165, 141, 830, 737, 177, 1004, 512, 167, 261, 241, 518, 31]. formalisms [531]. formally [255]. formation [1129]. formed [215]. forming [714]. forms [331, 740, 586]. formula [517]. formulae [215, 719, 657]. Formulas [568, 818, 543, 175, 266, 886, 1042]. Forward [822]. Foundations [721, 1163, 559, 253, 901, 731]. founded [780, 963, 624, 917, 915]. Four [372]. fraction [258]. fragment [582]. fragments [996, 565]. framework [245, 922, 786, 1171, 297, 241, 1028]. free [1132, 664, 312, 571, 792, 516, 830, 222, 210, 155, 740, 447, 772, 174, 581, 272, 1001, 494, 508, 308, 405, 416, 187, 437, 939, 542, 1110, 1010, 1139, 812, 159, 175, 90, 838, 908, 236, 987, 256, 537, 524, 472, 121, 76, 347, 12, 341, 475, 1144, 10, 60, 619]. free-choice [772]. freeness [521]. French [26, 858, 46, 826, 927, 827, 929, 516, 931, 831, 833, 834, 447, 835, 449, 451, 823, 837, 838, 840, 841, 936, 544, 1003, 795]. French-Soviet [26]. Friedberg [90]. friends [1178]. frontier [200, 1054]. fully [1024]. Fun [249]. function [514, 828, 111, 708, 889, 901, 1179, 1046, 290]. Functional [951, 6, 826, 1189, 623, 171, 11, 630, 207, 214, 921, 32, 569]. Functions [908, 94, 556, 930, 951, 234, 334, 991, 675, 669, 810, 781, 682, 1047, 79, 238, 1153, 211, 42, 177, 1078, 904, 217, 209, 1174, 223, 468, 842, 998, 758, 15, 224]. Functorial [2, 1171]. Fundamental [152, 396]. further [365].

G. [905]. Gale [1051]. Galmiche [1202]. Galois [1183]. game [111, 1084, 946, 1199, 1051, 473, 712]. Games [796, 744, 687, 417, 645, 714, 591, 1103, 490, 713, 966, 175, 947, 392, 489, 1083, 874, 875]. gap [192, 807]. garbage [1069]. Gasper [836]. General [904, 528, 517, 1171, 238, 452, 1036, 236, 605]. generalised [646]. generalization [1097, 684, 325, 9]. Generalizations [501]. generalize [387]. Generalized [416, 275, 818, 832, 863, 368, 177, 1017, 391, 1071, 748, 168, 726, 1118]. Generalizing [610]. generated [284, 357, 1143, 257]. Generating [765, 766, 1104, 1109, 176, 842]. Generation [99, 913, 1159, 1020, 816, 922, 1089, 61, 189]. generative [522]. generators [472]. genericity [909]. genre [826]. Gentzen [849]. Gentzen-type [849]. genus [826, 750]. genus- [826]. Geography [713, 745]. geometric [643, 811, 348, 1121, 619, 795]. geométrique [795]. geometry [766]. gerbes [447]. given [312]. Global [956, 188]. goals [460]. Gödel [1066, 1067]. good [710, 887]. gossiping [1023]. grammar [331, 740, 644, 1009, 236, 1151]. grammars [344, 218, 859, 830, 291, 155, 515, 230, 71, 542, 702, 143, 256, 121, 232, 522, 705, 347]. grammatical [983]. grams [412]. Graph [344, 695, 666, 697, 704, 130, 1069, 155, 995, 1158, 230, 343, 78, 711, 1063, 346, 602, 891, 790, 703, 1087, 705, 341, 475]. graph-based [341, 475]. Graphical [634]. GraphLog [806]. Graphs [159, 993, 417, 645, 192, 564, 699, 585, 107, 1103, 1081, 815, 36, 796, 357, 93, 1062, 545, 1023, 621, 61, 1199, 965, 536, 1191, 1100, 1144, 797, 386, 750, 564]. Grassmannians [843]. grates [950]. Grigorchuk [339]. Gröbner [1141]. ground [637]. Group [697, 737, 516, 190, 840, 471]. groupe [516, 840]. Groups [679, 447, 339, 437, 737, 840]. growth [748]. Grundy [111]. Grzegorczyk [989]. GSOS [1073]. Guarded [1031, 246, 4, 103]. guards [136]. guingois [827].

Hagino [728]. HAL [864]. halting

[107, 1054]. Hamiltonian [38, 1087, 1104]. handling [198]. Hanoi [925]. Hard [794, 655, 893, 800, 194, 1168, 739]. Hardy [446]. HARMONIC [526]. hashing [660]. Hausdorff [488]. Hausdorff-Kuratowski [488]. having [931]. Head [1167]. heaps [272]. HEAPSORT [1162, 850]. height [832, 390, 646, 45]. Hennessy [56]. Herbrand [145, 864]. Hereditarily [867]. Hereditarily-finite [867]. hereditary [18]. Heterogeneous [975, 642]. Hierarchies [266, 791, 389, 1009]. hierarchly [758]. hierarchy [488, 277, 583, 1147, 501, 989, 524, 205, 73, 749, 619]. Higher [784, 779, 375, 547, 562, 1185]. Higher-dimensional [784]. Higher-order [779, 1185]. Hindley [325]. Historical [905]. histories [384]. Hoare [247, 198]. Hoare-like [198]. home [571]. homomorphic [328]. Homotopies [1123]. honest [158]. hook [842]. Hopf [76]. Horn [358, 818, 1037, 156, 460, 969, 880, 304, 103, 242, 574]. Horspool [402]. hp [945]. hp- [945]. hulls [594]. Hurwitz [261]. Hyperedge [698, 1022, 357]. hyperedge-replacement [357]. hypergraphs [744, 48, 237, 597, 700]. Hypermap [295]. Hypothetical [115, 555].

I-categories [786]. ICLP [300]. idempotent [943]. Identities [824, 948, 1134, 363, 518, 76]. identity [611, 824]. If [82, 873, 37, 850]. if-then-else [37]. II [306, 597, 766, 936, 1051, 10]. II. [17]. images [816]. immigration [890]. immune [583, 1197]. immunity [69]. impact [522]. implementation [463, 612, 614]. Implementing [54]. implication [460]. implicit [773]. implies [1120]. imply [628]. Improved [42]. improvement [534]. Improvements [817]. inclusion [390]. incomplete [591, 1125, 151]. incompleteness [1066, 1067]. inconsistencies [62]. incremental [1129, 1065]. indefinite [808]. Independence [607, 1139, 577, 241, 1156]. independent [144, 971, 418, 313]. Indexed [396]. indexes [472]. Indicator [834]. indicatrices [834]. induction [885, 1128, 1177, 964, 607, 383]. inductive [710, 711, 321, 1071, 569]. inequality [836]. inevitability [885]. Inf [1179]. inference [902, 860, 710, 65, 1049, 142, 959]. Inferring [630]. infinies [929]. infinitary [884, 439]. Infinite [237, 597, 596, 929, 431, 481, 264, 753, 1103, 676, 979, 275, 749]. infix [1196]. Informatika [858, 26]. Information [776, 8, 591, 1125, 110, 1061, 1130]. Informatique’ [924]. inherently [989]. inheritance [631]. initial [960, 314]. injective [952]. inkdot [1011]. inner [1088]. insertion [988]. instances [203]. instructions [546]. Integer [496, 807]. integers [1036, 1018]. integrating [705]. integrity [118]. intensive [474]. Interactive [754, 1161]. interconnection [974]. Interdefinability [182]. Interleaving [109]. intermittent [885]. International [300, 1131, 131, 99, 671, 695]. interpolation [274]. interpretability [58]. interpretable [904]. Interpretation [795, 958, 902, 1125, 1071, 409, 918]. Interpretations [638, 723, 331, 882]. Interpreting [100]. Intersection [558, 579, 1052]. interval [930, 938]. introducing [715]. Introduction [444, 845, 905, 907, 171, 1141]. Intuitionistic [1125, 785, 582]. invariance [314]. invariants [41, 789]. inverse [943, 316]. Inversion [1152, 66]. invertible [1094]. Investigation [965]. invisible [690]. involution [837]. involving [466, 641]. Irrational [748]. isomorphic [931]. isomorphie [931]. isomorphism [1047, 1186, 267, 471]. isomorphism=step [178]. Issue [400, 913]. ISWIM [911]. iterates [238, 75]. iteration [352, 193, 332]. iterations [932, 935]. iterative [989].

job [1064]. jobs [1055]. Join [1082]. ‘Journées [924]. July [400]. June [913]. jungle [698].

Kahn [905]. Karp [338].

Karp-Miller-Rosenberg [338]. Kayles [489]. kernel [778, 495]. Khrapchenko [817]. kind [437, 1074]. Kleene [674, 979, 279]. knapsack [342]. knight [490]. knowledge [269, 1101, 1074, 385]. Knuth [51, 430, 555, 736]. Koiran [1113]. Kolmogorov [265]. Kriegspiel [490]. Kunen [500]. Kuratowski [488].

L [1160, 890, 1017, 280, 259]. labeled [17]. Labelled [1173, 1073, 1089]. labelling [926]. labels [705]. Lagois [1183]. lambda [899, 900, 728, 665, 1187, 184, 908, 325, 1003, 423, 1185]. lambda-calculi [728]. lambda-calculus [325, 423]. lambda-termes [1003]. lambda-terms [900, 1003]. Language [814, 1079, 1189, 133, 728, 864, 1094, 263, 1091, 987, 292, 495, 921, 465, 198, 801, 139, 1146]. Language-theoretic [814]. Languages [671, 1131, 268, 487, 1140, 161, 488, 805, 312, 149, 872, 1099, 442, 222, 592, 221, 219, 1101, 284, 343, 260, 102, 317, 311, 508, 330, 1160, 187, 737, 800, 486, 542, 1110, 1010, 906, 812, 54, 72, 617, 891, 1092, 1049, 168, 1172, 333, 524, 894, 568, 31, 472, 214, 646, 521, 12, 341, 475, 616, 122, 124, 123, 619, 985, 301]. large [585, 974, 873]. lark [720]. lattice [981, 506]. lattices [1001]. laws [451]. layers [1081]. layouts [536]. lazy [1189, 692]. leader [216]. Learnability [315]. Learning [462, 121, 618, 1129, 1021, 879]. least [1070, 856]. least-fixed-point [856]. Left [962, 644, 416]. left-associative [644]. left-cancellative [416]. Legality [881]. length [652, 470, 312, 55, 555, 686]. lengths [628]. letter [1134, 210, 71, 653, 1079, 1091]. letters [467]. level [459]. lexicographically [337]. libre [516]. libres [447, 838]. Lie [174, 175, 838]. lies [473]. like [1040, 198, 354]. limit [451, 707, 1015]. Limitedness [209, 42]. limités [451]. limiting [1082]. limits [316]. line [744, 470, 1058, 231, 685, 796, 1063, 128, 982, 1016, 1027, 895, 1055]. linéaires [449]. Linear [783, 723, 527, 401, 463, 1165, 930, 1198, 1008, 77, 582, 640, 1166, 493, 969, 285, 236, 1169, 1168, 940, 1167, 765, 766, 256, 566, 292, 866, 1202, 937, 411, 96, 449]. linear-space [401]. lines [690]. list [85, 878, 423, 692, 1060]. list-oriented [423]. lists [262]. liveness [581]. LLL [932]. load [1058]. Local [481, 361, 1050, 188, 1061, 599, 44]. localities [770]. Localizable [384]. Locally [673, 14, 259]. Locating [1147]. location [81]. log [655, 233, 140]. log-space [655]. Logarithmic [540, 1105]. logarithmic-cost [1105]. Logic [300, 916, 723, 923, 664, 463, 1165, 150, 783, 302, 610, 1124, 914, 967, 784, 303, 958, 264, 1070, 593, 698, 666, 192, 564, 1188, 629, 154, 107, 558, 582, 145, 624, 343, 102, 440, 627, 1166, 79, 452, 730, 460, 493, 565, 598, 529, 179, 247, 243, 80, 625, 109, 106, 394, 58, 1169, 1168, 304, 137, 252, 605, 765, 766, 480, 947, 496, 729, 379, 6, 881, 599, 1202, 214, 1039, 720, 499, 25, 127, 957, 886, 301, 574, 56]. Logical [731, 769, 634, 804, 1074, 641, 479, 921, 853]. Logically [428]. logics [484, 1037, 716, 774, 1040, 7, 138, 461, 181, 385, 869]. logspace [114, 661]. lois [451]. longest [401]. look [422]. look-ahead [422]. lookahead [236]. loop [302, 610, 554]. lop [827]. lop-sided [827]. loss [973]. Low [806, 945, 309, 1147, 1093]. Low-complexity [806]. low-dimensional [1093]. Lower [1064, 94, 1063, 507, 216, 1115, 255, 1146].

LR [623, 236]. Lyapunov [986]. Lyndon [1109, 838, 650].

m [940, 656]. M-program [656]. Machine [333, 483, 920, 87, 310, 292]. machines [245, 608, 231, 685, 1057, 318, 1102, 348, 755, 285, 538, 350, 1011, 227, 1025, 1155, 1054, 1185, 492, 546]. Macsyma [173, 172]. maintenance [48]. making [949]. Malcev [516]. Malcev-Neumann [516]. many [262, 356, 441, 502]. many-one [502]. many-ordered [262]. many-sorted [356, 441]. Map [578, 1184, 273]. mapped [75]. mapping [938]. mappings [239, 119]. maps [826, 147, 307]. Marcus [622]. Markov [1032, 1051]. marriages [1016]. matches [412]. Matching [399, 400, 206, 74, 635, 403, 1108, 669, 405, 1078, 432, 1016, 407, 409, 410, 1194, 734, 412]. Mate [490]. Mathematical [721, 1163, 490, 713, 613]. Mathematics [924]. Mathématiques [924]. Mathématiques-Informatique’ [924]. Matos [1200]. matrices [309, 518, 456]. matrix [1088, 231, 685]. max [272]. Maximal [270, 470, 337, 412, 681]. Maximality [580]. maximize [977]. McNaughton [275, 525]. mean [113]. meanders [839]. means [12]. measure [180, 208]. measures [1009, 917]. mechanical [91]. mechanism [198]. mechanisms [302, 1128]. Mellin [944]. memoization [527]. memory [1150, 878, 712]. MEO [1158]. Merging [307, 262, 272, 1032, 235]. merits [1068]. mesh [1033, 892]. mesh-connected [892]. message [59, 1025]. Metacircularity [360]. method [451, 602, 1068, 1130, 91]. Méthode [451]. Methodology [131]. Methods [26, 858, 1025, 853, 988, 851, 519, 934, 613]. Métivier [927]. metric [1170, 759, 101]. metrics [532]. Miller [338]. Milner [56]. min [272]. min-max [272]. Minimal [1098, 263, 172, 1034, 470, 306, 872, 1158, 587, 968, 287]. Minimisation [411]. minimization [528, 971, 349]. Minimizing [113]. Minimum [594, 554, 545, 386]. miracles [561]. misere [489]. missing [948]. Mixed [29, 367, 603, 369, 6]. MLOG [921]. mob [388]. mobile [1064, 774]. Möbius [514, 828, 675]. MOCA [1027]. mod [95]. mod-classes [95]. Modal [774, 729, 385, 1037, 293, 996, 864, 1040, 361, 254]. modalities [1168]. modally [89]. Model [993, 656, 435, 481, 829, 1113, 922, 803, 320, 51, 83, 598, 318, 57, 804, 430, 241, 718, 480, 761, 361, 254]. modeling [653]. models [1082, 961, 247, 464, 394, 667, 1119, 325, 1065, 31, 910, 301]. mods [590]. Modular [157, 764, 920, 493, 981, 253, 342]. Modularity [601, 1192]. module [134]. modules [718, 841]. Modulo [997, 306, 611, 1142]. Modulo-counting [997]. moléculaires [831]. molecular [831]. Monadic [699, 995, 192, 564, 343, 757, 266, 504, 1142]. Monads [533]. monoid [1132, 941, 537]. monoids [1133, 278, 943, 308, 405, 1139, 159, 1143, 398, 76, 1144, 525, 60]. monotone [229]. Monte [934]. Monte-Carlo [934]. Moore [402, 408]. Morphic [1160]. morphism [75]. morphisms [16, 355]. Morse [448, 1050]. MOS [378]. most [125, 742, 354]. mots [447, 544]. Motzkin [1159]. move [85]. move-to-rear [85]. movement [474]. mu [361]. mu-calculus [361]. Muller [488, 275]. Muller-McNaughton [275]. multi [635, 1161]. multi-pattern [635]. multi-prover [1161]. multicomputer [991]. multicounter [350]. Multidimensional [282, 934, 206, 1155]. Multihead [289]. Multimodal [629]. multiparameter [523]. Multiple [935, 631, 661, 1032, 138, 347]. multiple-disk [1032]. multiple-valued [138]. multiples [693]. multiplexing [410]. Multiplication [546]. multiplicative [1169, 586]. multiplicatively [416]. Multiplicities [511, 1134]. multiprocessor [975, 1027]. multiset [628]. multisets [708]. multitape [160]. multivalued [314, 1136]. multivariate [274]. mutate [630]. mutual [1034]. mutual-exclusion [1034].

N [427, 940]. N-resolution [427]. NAF [883]. name [1036]. narrowing [50]. nationality [516]. natural [928, 1182]. NC [36]. nearly [931]. necessity [551]. necklaces [588]. need [461]. Negation [150, 149]. negative [855]. neighbourhoods [426]. Nested [534, 330]. net [482, 1176, 1167, 24]. nets [1135, 1201, 1138, 581, 146, 485, 961, 429, 243, 80, 530, 1177, 185, 395]. network [978, 991, 554, 1026, 114, 588]. networks [976, 889, 129, 587, 974, 973, 819, 382, 1080, 892, 199]. Neumann [516]. neural [1135, 1201, 1080, 199]. Newton [1118]. NEXPTIME [1168]. NEXPTIME-hard [1168]. Nivat [505]. NLC [71]. No [1199]. node [281]. nodes [1033]. noetherianity [436]. noise [1045]. nombre [931]. Non [43, 110, 167, 255, 1141, 524, 73, 912, 156]. non-commutative [167, 1141]. non-computable [73]. non-context-free [524]. Non-deterministic [110]. non-existent [912]. non-Horn [156]. Non-repetitive [43]. non-trivial [255]. Nonacceptability [260]. nonadaptive [227]. Nonclairvoyant [1056]. Noncommutative [172]. noncounting [548]. Nondeterminacy [966]. nondeterminism [511, 324]. Nondeterministic [984, 152, 791, 351, 373, 1102, 393, 82]. Nonerasing [1054]. nonisomorphism [211]. nonlinear [169, 171]. Nonmodal [582]. Nonrecursiveness [39]. nonstandard [298, 1120]. nonsubsequence [691]. Nonterminal [230]. Nonuniform [486, 297, 794]. normal [666, 1082, 625, 1040]. normalization [248, 1166, 298, 1202]. normed [939, 1085]. Norms [967]. not=NP [346]. Note [925, 746, 1023, 161, 4, 1111, 694, 62, 1085, 747, 350, 457, 254]. Notes [463]. NP [1114, 896, 893, 211, 194, 1169, 292, 691, 873, 1120, 1120]. NP-complete [211, 292, 691, 896, 1169]. NP-hard [893, 194]. NP-P [873]. NU [454]. number [312, 163, 931, 210, 932, 203, 415, 935, 258, 1119, 228, 936, 820, 954, 588, 522, 473]. numbers [1084, 549, 345, 39]. numeration [640, 449, 449]. numerical [1008, 91].

O [233]. object [805, 1148, 384]. object-oriented [805, 1148]. objects [298, 116]. observable [1135, 1201]. Observational [484]. Observing [770]. obtaining [527]. Occam [551]. occur [25]. Ochmanski [979]. octal [489]. off [231, 685, 128]. off-line [231, 685, 128]. offs [229]. old [29]. omega [488, 960, 284, 168, 596, 122, 648]. omega-complete [960]. omega-continuous [168]. Omega-decision [648]. omega-languages [284, 122]. omega-regular [488]. omega-transduction [122]. omega-words [596]. On-line [744, 470, 1058, 1016, 1055, 1063, 982, 1027, 895]. once [657, 661]. One [1047, 211, 1100, 846, 943, 941, 1134, 70, 231, 685, 737, 71, 694, 285, 350, 289, 870, 1079, 1091, 1107, 292, 100, 876, 502]. one-counter [737, 694]. one-letter [1134, 71, 1079, 1091]. one-tape [231, 685, 292]. One-way [1047, 211, 1100, 694, 285, 350, 289, 870]. only [1169, 492]. onto [174]. operation [825, 843]. Operational [495, 28]. operations [389, 163, 835, 631, 1094, 809, 709, 1130, 705, 39, 182, 985, 835]. operator [164, 240, 778, 856]. operators [830, 986]. opfibrations [1069]. Opt [501]. Optimal [406, 432, 949, 276, 797, 144, 975, 1108, 329, 129, 1162, 1023, 85, 878, 189, 81, 892]. optimality [434]. Optimally [36, 1147]. optimistic [1026]. Optimization [972, 805, 621, 365]. optimizing [1048]. Oracle [1008, 1111, 348, 202]. oracles [583, 419, 418, 659]. orbits [1086]. Order [356, 631, 928, 514, 828, 293, 192, 564, 699, 995, 629, 779, 854, 343, 317, 79, 66, 52, 565, 724, 916, 1116, 263, 1168, 999, 807, 1185, 626, 375, 395, 438, 1146]. Order-sorted [356, 629, 66, 52, 438]. ordered [262, 403, 1114]. ordering [1158, 108, 704]. orderings [8, 628, 1175]. orders [872]. organization [878]. organizing [85]. orientations [309]. oriented [805, 1148, 866, 423, 1075]. Orthogonal [174, 833]. orthogonaux [833]. Othello [946]. other [216, 1114, 910]. out-degree [309]. outerplanar [36]. output [685]. Overlap [792]. Overlap-free [792]. overloading [631]. overview [1124, 1126]. OWHY [911].

P [1203, 1012, 1114, 893, 346, 1147, 501, 114, 1197, 1120, 130, 873, 1120]. P-bi-immune [1197]. P-complete [1203, 130, 1012]. P-selective [893]. P/poly [1147]. Padding [213]. pairs [1098]. PAL [849]. paper [1200]. paperfolding [1044]. Papers [400, 913]. Paraconsistent [425]. Parallel [262, 417, 1006, 701, 285, 117, 895, 474, 206, 1024, 572, 222, 1008, 338, 1108, 1158, 329, 1057, 78, 1115, 545, 649, 57, 1046, 432, 1199, 23, 1025, 182, 1151, 546, 91]. parallelism [318]. parallelization [64, 129]. parameterized [1171]. parameters [287]. parametric [898]. Parametricity [1186, 909]. parametrization [933]. Paris [400]. parity [590, 290]. parser [623]. parsing [285, 622, 236, 895]. Part [716]. Partial [527, 872, 321, 336, 1184, 367, 234, 334, 108, 1138, 781, 102, 368, 79, 631, 788, 263, 917, 704, 533, 999, 959, 395]. partial-order [395]. partially [210, 491, 308, 405, 159, 838, 76, 541]. partiellement [838]. Partition [130, 1046]. partitioning [146]. partitions [842]. Pascal [748]. path [344, 652, 714, 587, 628, 55, 443, 1118, 1068, 555, 797]. path-controlled [344]. path-following [1118]. path-forming [714]. paths [470, 815, 837, 276, 273]. Patricia [652]. Pattern [445, 399, 400, 27, 635, 1108, 669, 410, 734, 257]. Patterson [684]. pavings [307]. PCF [182]. PDL [324]. perfect [645, 269, 718]. Performance [975]. period [1199]. Periodic [1018, 584, 257]. Periodicity [225]. permanent [585]. permutation [797, 386]. permutations [1086]. perpetual [529, 103]. Perrier [1202]. persistency [429]. personal [803]. perspective [803]. Petri [482, 1138, 146, 485, 961, 429, 243, 80, 530, 1177, 185, 395]. Picture [690, 219, 72, 617, 1092, 252]. pictures [219]. piecewise [930]. piles [796]. Pinwheel [549]. Planar [309, 346, 738, 815, 93, 1062, 536, 81, 876]. Plane [839, 988, 420, 147, 194, 842]. plane-topology [420]. plans [765, 766]. plasma [27]. plethysm [843]. Plotkin [905]. point [393, 286, 327, 5, 81, 856, 316]. pointed [826]. pointées [826]. pointer [272]. pointer-free [272]. pointers [1105]. points [3, 145, 1017, 544, 10]. policies [972]. poly [1147]. polyhedral [594]. Polymorphic [248, 184, 1189, 358, 464, 421, 909, 360]. polymorphism [898, 2, 787]. polynômes [833]. Polynomial [389, 202, 553, 471, 306, 583, 158, 581, 335, 493, 1199, 501, 965, 205, 1121, 96, 121, 867, 639, 758, 433, 1123, 502, 1122]. Polynomial-time [202, 471, 581, 335, 493, 867, 758, 502]. polynomially [193]. Polynomials [172, 306, 274, 833, 758, 15, 1]. Polyominoes [1013, 173]. poset [884]. posets [931, 14, 776]. positions [1044]. Positive [140, 456]. possibilities [365]. possible [153, 296]. possible-world [296]. possibly [441]. Power [821, 509, 724, 673, 164, 825, 165, 830, 264, 221, 517, 1161, 710, 21, 1181, 177, 348, 755, 512, 421, 261, 773, 544, 522, 541]. powerdomain [603]. powerdomains [387]. powers [843]. practice [364]. PRAM [688, 1024]. PRAMs [19, 953]. predicate [582, 322, 947]. predicates [724]. Preface [782, 633, 768, 696, 477, 913, 970, 992]. preferences [151]. prefetching [1032]. prefix [636, 1153, 1196, 1143]. prefixes [415]. premises [855]. près [931]. Prescribing [371]. presence [327]. presentations [737]. presented [943, 14]. Preservation [106]. preserving [580, 443, 626]. Primitive [1095, 250, 669, 547, 628]. Primitivity [18]. Principal [665]. principle [964, 607]. Principles [27, 885]. priorities [790]. priority [482, 90, 86]. priority-free [90]. Probabilistic [204, 415, 653, 758, 297, 430, 433]. probability [829]. Problem [866, 130, 1034, 926, 446, 943, 971, 388, 107, 204, 1059, 598, 160, 1046, 80, 949, 114, 987, 1054, 691, 1087, 797, 386, 541, 692, 1060, 1019]. Problem-oriented [866]. Problème [447]. Problems [1203, 474, 1098, 814, 846, 466, 643, 216, 811, 41, 29, 784, 20, 1012, 447, 77, 134, 1076, 896, 357, 405, 243, 649, 678, 794, 980, 337, 342, 33, 734, 1151]. procédés [929]. procedure [463, 602, 1040]. procedures [352, 929, 494, 421, 397]. process [240, 573, 34, 10, 764]. Processes [531, 1164, 466, 373, 939, 1085, 1177, 774, 670, 103, 195, 919]. Processor [621, 846, 971]. processors [216, 59, 820, 954]. product [1088, 306, 1086, 863, 491, 974, 183, 261, 176]. product-form [974]. production [929]. Products [123, 328, 84, 186, 680]. Professor [653]. Program [26, 858, 1072, 450, 1021, 381, 561, 656, 967, 119, 30, 882, 366, 106, 916, 762, 365, 32, 615, 918]. Programmation [46]. Programming [46, 300, 721, 1163, 154, 923, 664, 610, 698, 629, 105, 107, 728, 440, 404, 731, 649, 906, 1172, 865, 1119, 704, 379, 455, 1039]. Programs [788, 498, 883, 956, 150, 302, 914, 303, 958, 593, 666, 885, 851, 862, 145, 624, 627, 452, 358, 529, 747, 625, 443, 304, 605, 630, 454, 496, 103, 866, 881, 599, 7, 305, 142, 181, 869, 68, 1, 25, 127, 957]. projection [174, 34]. projective [839]. Prolog [244, 101, 28, 860, 862, 614, 1190]. promise [794]. Proof [851, 298, 56, 467, 785, 836, 754, 716, 1166, 864, 297, 1126, 602, 90, 209, 255, 727, 606, 1167, 1040, 1202, 1029, 397]. Proofs [1164, 191, 899, 612, 628, 1174, 765, 455, 842, 1042]. Properties [983, 152, 237, 666, 885, 851, 260, 1050, 1149, 882, 711, 894, 999, 492, 124, 1193]. property [652, 899, 1095, 437, 184, 637]. Propositional [777, 1037, 394, 869, 886, 857]. propositions [255]. protocols [269, 1161]. provable [534]. prove [383]. prover [1161, 614]. Proving [999, 967, 885, 20, 436, 882, 853, 91]. pseudorandom [659]. PSPACE [340, 946, 502]. PSPACE-complete [340, 946]. Puissances [544]. pure [244, 1167, 607]. pushdown [1007, 505]. Pushout [143, 701, 703]. Pushout-complements [143]. pyramid [1194].

QRT [291]. quadratic [1101, 217, 1118, 1119, 586]. quadtree [1194]. quantifier [506]. Quantifiers [662, 997]. Quantitative [439]. Quasi [952, 8, 934, 940]. Quasi-injective [952]. quasi-linear [940]. quasigroups [984]. quasiperiodicities [871]. Quelques [929]. queries [277, 1148, 155, 804, 286, 807, 618]. query [144, 498, 805, 330, 67, 227, 81]. querying [153]. quest [1034]. question [1114]. questions [516]. queueing [642]. queues [976, 755]. QUICKSORT [850]. quite [887]. quotient [852].

R [471]. R-strongly [471]. Rabin [1133]. RAM [1006, 233]. Ramsey [785]. random [1020, 746, 1089, 418, 546]. Randomization [1083]. randomized [743, 657]. randomness [486, 661]. range [899]. rank [1050, 537, 742]. Ranking [165]. Rapid [376]. ratio [1064]. Rational [510, 472, 513, 446, 352, 1134, 676, 1153, 169, 363, 693, 15, 122, 541]. rationale [105]. rationalité [516]. Rationality [514]. Reachability [772, 1084, 994, 530, 876]. read [657, 661]. read-once [657, 661]. real [446, 271, 350, 1027, 258, 345, 1119, 1127, 1075, 39]. real-time [350, 1027, 1127, 1075]. Realizability [1071, 918, 381]. realization [328]. reals [1114, 1116, 1120]. realtime [694]. rear [85]. reasoning [857, 777, 494]. recoding [1053]. Recognising [700]. Recognition [1191, 222, 820, 954, 524, 795]. recognizability [192, 544]. Recognizable [308, 405, 681]. recognizer [187]. Recognizing [220, 892, 801]. reconfigurable [889]. reconnaissabilité [544]. reconnaissance [795]. rectilinear [536]. recurrence [813]. recurring [846]. Recursion [127, 246, 424, 966, 789, 56]. Recursive [240, 249, 155, 622, 7, 455, 597, 1189, 1106, 234, 334, 250, 901, 669, 34, 547, 961, 628, 904, 421, 506, 98]. Recursively [808, 221, 1174, 964, 638]. reducibilities [335, 202]. reducibility [1090, 213, 637]. reducible [346]. reduction [528, 1010, 907, 1167, 349, 233, 668]. reductions [952, 793, 853, 553]. refined [1187]. Refinement [847, 1127, 239, 322, 1075, 775]. reflectors [238]. regions [738, 1043]. Regular [1041, 888, 598, 1117, 520, 488, 872, 636, 221, 608, 548, 343, 485, 125, 61, 812, 391, 263, 968, 568, 1087, 1143, 749, 124, 1146, 354, 503, 985]. regularities [60]. Regularity [180, 341, 475]. related [816, 648, 974, 255, 734]. Relation [323, 826, 436, 241]. relational [387, 564, 778, 67, 809, 326, 459]. Relations [95, 1041, 535, 291, 510, 624, 676, 180, 382, 596]. relationship [485, 605]. relative [43, 390, 659, 1068]. Relativizations [1114, 140, 433]. Relativized [334]. relator [943]. release [113]. remark [1200]. removal [424]. renaming [969]. renewal [457]. repair [975, 1031]. repetitive [43]. replacement [698, 357]. replicated [1030]. reply [653]. report [600]. representable [669]. Representation [17, 1153, 345, 172, 831, 153, 930, 831, 221, 449, 536, 385, 499]. représentations [449]. represented [84]. requirements [878]. Research [697]. residua [880]. residuosity [1101]. Resolution [293, 883, 88, 215, 818, 303, 427, 358, 686]. respect [498, 315, 70, 263, 241]. Restricted [953, 433, 72]. restrictions [1021, 579]. result [498, 528, 348, 325]. Results [328, 71, 69, 900, 610, 1198, 1152, 1176, 982, 196]. Retreat [1092]. retrieval [1060]. returns [935]. reversal [70, 617]. reversal-bounded [617]. reversals [128]. Reversible [57]. review [196]. revised [600]. revisited [474, 463, 88, 543]. Rewrite [251, 811, 1007, 608, 54, 1107]. rewriting [1069, 927, 248, 994, 43, 636, 698, 107, 637, 601, 790, 757, 480, 704, 1192, 1142, 33, 295, 1077, 735, 383]. Rice [1015]. Richelieu [814]. right [841, 12]. right-congruences [12]. ring [520, 513]. rings [216, 517, 1114, 59, 1177]. Robust [87]. robustly [335]. role [1039]. root [200]. root-to-frontier [200]. roots [1121]. Rosenberg [338]. Rosser [184, 423]. Rotating [687]. Rotating-table [687]. round [269]. routing [1025]. row [1033]. rule [883, 215, 149, 608, 1107, 253]. rule-based [149, 253]. rules [811, 727]. running [402, 128, 236].

S [940]. SA [1004]. saccept [87]. samples [376, 391, 1049]. sampling [809]. sand [796]. Sardinas [684]. SAT [277]. satisfaction [496]. satisfiability [1148, 1037, 203, 204, 56, 233]. satisfying [423]. scanners [268]. Scattered [990]. scheduling [1064, 1057, 549, 1027, 1056, 1055]. schemas [375]. scheme [271, 220]. schemes [144, 528, 32, 259, 560]. School [400, 399]. Schreier [453]. Sci [1201, 1203, 1202]. Science [858, 924, 364]. scientific [897]. Scott [8]. search [288, 824]. Searching [162, 688, 473]. second [192, 564, 699, 995, 343, 317, 565, 724, 1131, 509]. second-order [192, 564, 699, 995, 343, 317, 565, 724]. Section [715]. segments [795]. Selected [400, 913]. selection [66, 552]. selective [893]. self [213, 346]. self-reducibility [213]. self-reducible [346]. Semantic [667, 301, 719, 1030]. Semantics [1163, 494, 304, 116, 438, 25, 1170, 319, 928, 482, 28, 914, 593, 108, 244, 1138, 624, 135, 627, 731, 358, 663, 370, 906, 393, 120, 1172, 718, 955, 605, 326, 704, 5, 533, 103, 7, 915, 495, 151, 761, 1075, 396, 296, 395, 1190, 957, 1035, 101, 574, 721]. Semi [212, 498, 271, 675, 612]. semi-algebraic [271]. semi-algorithm [612]. semi-commutations [675]. semi-strict [498]. Semi-unification [212]. Semicommutations [592, 942]. semidirect [680]. semigroup [416]. semigroups [673, 680, 279, 332]. semilinear [618]. semilinearity [429]. Semiretracts [1132]. semiring [1134]. semirings [416, 512, 168, 349]. sensitivity [870]. separated [891]. Separating [648, 419, 891, 310]. Separation [351, 1113, 230, 208, 196]. separations [583]. separators [750]. sequence [448, 1050, 469, 178]. sequences [445, 520, 1044, 929, 329, 35, 879, 693, 584, 458]. sequent [781]. Sequential [874, 600, 136]. sequentiality [763]. serial [206]. serializability [956]. Series [821, 164, 825, 165, 141, 516, 830, 510, 834, 517, 169, 177, 512, 167, 261, 877, 176, 541, 513, 834, 509, 516, 834]. server [1059]. servers [526, 642]. service [976]. ses [840]. Set [910, 780, 963, 707, 75, 969, 109, 1197, 742, 709, 197]. Set-theoretical [910]. sets [153, 945, 69, 583, 234, 334, 893, 335, 357, 87, 211, 193, 104, 812, 678, 1015, 391, 441, 1018, 739, 867, 506, 749, 618, 202, 799, 283, 127, 1060]. Seven [1019]. Seven-state [1019]. several [755, 175]. Shallow [950]. shared [601]. sheaf [734]. sheaf-theoretic [734]. shells [708]. shift [546]. shifting [819]. short [897, 836]. Shortest [738, 273, 815, 1157, 981, 691]. Shuffle [176, 491, 798, 588]. shuffle-exchange [588]. Sibert [489]. Sibert-Conway [489]. sided [827]. sigmoids [1115]. signatures [736]. similarité [833]. similarity [833]. Simple [28, 317, 319, 583, 829, 781, 747, 601, 606]. simplicity [539]. simplification [92]. simply [904, 908]. simulated [534]. Simulation [608, 1024]. Single [639, 281, 701, 1046, 703]. single-pushout [701, 703]. Single-valuedness [639]. singly [271]. sink [876]. Sixth [721]. Size [229, 334, 1021, 659]. Size-depth [229]. sizes [1082]. sketch [9]. Sketches [933]. Skew [517, 842]. SLDNF [883, 303]. SLDNF-resolution [303]. slender [1110, 894]. slight [500]. small [1011, 850]. smaller [820, 954]. snake [1076]. Snakes [875]. SnS [89]. Software [131, 977, 67]. soliton [281, 84]. solution [388, 1079]. solutions [1034, 976, 203, 1115, 1019]. solvability [1076]. solvable [77]. Solving [736, 811, 786, 96, 1123]. Some [900, 1176, 1009, 982, 167, 33, 492, 912, 396, 24, 1151, 458, 1193, 161, 234, 293, 852, 519, 1152, 707, 1023, 649, 469, 869, 985, 929]. sorted [629, 135, 356, 66, 52, 631, 441, 438]. Sorting [595, 844, 129]. Sound [917].

Soundness [441]. source [1130, 876].

sousgroupes [840]. Soviet [26, 858]. Space [196, 655, 401, 446, 1014, 1111, 896, 754, 1102, 538, 1011, 57, 567, 709, 469, 492, 82, 140, 224, 984]. space-bounded [538, 492]. space-efficient [896]. space-model [57]. spaces [299, 481, 726, 759]. spare [1033]. sparse [274, 335, 739, 410]. sparsity [404]. Special [400, 913]. species [831]. specifiability [287]. Specification [362, 1171, 613, 1028, 1075, 615, 764]. specification-oriented [1075]. specifications [561, 960, 634, 34, 134, 1171, 855, 318, 136, 916, 336, 773, 118, 729, 185, 438]. specified [139]. spectra [445, 965, 242]. speed [1106, 1045, 848, 566, 313]. speed-independent [313]. speed-up [1106, 848, 566]. speeds [748]. Speedup [1155]. splitting [146]. Sprague [111]. Sprague-Grundy [111]. Squad [1019]. square [210]. square-free [210]. squares [1041]. Stability [683, 190]. Stable [1181, 726, 426, 725, 624, 1016, 605]. Stack [870, 844]. stacks [966]. Standard [680, 244, 587, 841, 259]. standard-path [587]. standardising [1039]. Star [508, 125, 390, 1079, 646, 354, 503]. Star-free [508]. start [1055]. state [240, 872, 481, 314, 600, 188, 285, 1048, 985, 1019]. Statecharts [576]. states [571, 492]. Static [370, 972]. stationary [234]. Statman [606]. status [882]. Stiefel [843]. Stiefel-Whitney [843]. Stochastic [1156]. storage [1036]. strategies [975, 50, 809, 85, 1032, 1051]. strategy [156, 878]. stratification [271, 599]. Stratified [1070, 305]. streams [918]. strength [191]. strengthening [500]. strict [498]. strictly [181]. strictness [902]. String [74, 403, 405, 406, 407, 757, 409, 1142, 668, 412]. String-matching [403, 405, 407, 412]. string-rewriting [757, 1142]. strings [871, 312, 746, 338, 75, 226, 879, 263, 1175]. Strong [583, 335, 532, 73, 248, 727, 668, 471, 681]. stronger [106]. strongly [921, 471]. Structural [787, 22, 429, 1177, 121, 232, 705]. Structure [733, 27, 746, 636, 38, 1097, 272, 21, 530, 178]. structured [593]. Structures [810, 926, 484, 1095, 572, 564, 847, 1043, 18, 16, 17, 414, 1097, 1089, 1062, 653, 823, 613, 207, 479, 398, 735]. studies [458]. study [550, 528, 986]. studying [107]. Sturmian [35, 228, 355, 1193]. style [555]. styles [362]. sub [840]. sub-groups [840]. subclasses [16]. subdirect [550]. subexponential [799]. subexponential-time [799]. subfamilies [122]. subgraph [337]. sublinear [649]. submonoid [75]. submonoids [308, 683]. subsequences [401, 162, 226]. subset [1052]. subsets [487, 537, 428]. substitution [544]. substitutional [544]. substitutions [1099]. substring [445, 688]. subsumption [602]. subsystems [938]. subtrees [689]. subtypes [65]. Subtyping [319]. subword [514]. success [574]. successful [288]. Successive [78]. suffix [1153]. suites [929]. Sums [693, 392, 944, 453, 595]. supersequence [980]. superstring [980]. superstrings [1157]. supported [605]. survey [149, 995, 697, 907, 567]. switching [587]. Symbolic [930, 613, 831]. symbolically [811]. symbolique [831]. symétrie [823]. symétrique [840]. symmetric [556, 893, 843, 199]. symmetrical [840]. symmetry [94, 194, 823]. Symposium [858]. Synchronization [1019, 1026, 136, 374]. Synchronized [676, 793, 1005, 694, 1102, 538, 126, 492]. synchronizing [584]. synchronous [559]. syntactic [600, 1143]. syntax [377, 868]. Synthesis [21, 30, 916, 381, 918]. system [923, 927, 377, 1134, 43, 1179, 519, 30, 855, 864, 1126, 1027, 865, 718, 454, 866, 1121, 198]. Systèmes [449]. Systems [99, 913, 1073, 462, 1024, 559, 482, 571, 994, 975, 293, 1007, 593, 698, 920, 772, 107, 861, 110, 642, 754, 640, 449, 716, 83, 977, 1128, 169, 1176, 890, 1017, 1009, 297, 907, 1093, 637, 1030, 363, 1117, 601, 225, 56, 1107, 790, 757, 256, 478, 1192, 1142, 253, 1127, 457, 96, 626, 1029, 1075, 521, 33, 1077, 1151, 383, 385, 1123, 776, 362, 523, 1035, 257, 1031, 597]. Systolic [468, 260, 21].

T [233]. table [687, 800]. tableau [1040]. tableau-like [1040]. tableaux [832, 1068]. tables [410]. Tail [789]. tally [419]. tape [70, 231, 685, 848, 128, 292]. tapes [70]. tasks [846]. taxonomy [243]. techniques [643, 216, 20, 816]. Technology [131, 614]. Tecton [1126]. template [1194]. Temporal [461, 191, 264, 598, 243, 80, 109, 58, 252]. Tentative [1026]. Term [1069, 994, 852, 698, 107, 432, 54, 637, 601, 1192, 383]. term-rewriting [698, 637]. termes [1003]. Terminal [780, 963, 71]. Termination [628, 611, 967, 962, 601, 1107, 1174, 1192]. terms [900, 967, 665, 1102, 1003]. tessellation [982]. test [278, 812, 341, 475]. Testing [1148, 194, 756, 637, 599, 434, 274, 1052, 539, 471]. tests [977]. texts [810]. their [97, 484, 278, 967, 291, 110, 485, 226, 318, 657, 1017, 461]. theme [1180]. theorem [785, 500, 848, 903, 1015, 505, 817, 1036, 90, 209, 909, 208, 606, 566, 1066, 423, 73, 275, 1122, 614, 1067, 525, 91]. theorems [474, 979, 1115, 42, 255, 223]. Theoret [1201, 1203, 1202]. Theoretic [509, 814, 11, 734]. Theoretical [858, 910, 911]. theories [945, 4, 600, 104, 641]. Theory [16, 17, 364, 64, 1000, 780, 963, 559, 201, 572, 763, 901, 519, 1097, 1171, 803, 450, 730, 578, 22, 378, 179, 1004, 1187, 678, 1116, 23, 453, 762, 936, 504, 1029, 506, 615, 735, 24, 124, 647, 499, 148]. there [244]. thesis [8]. theta [449]. theta-representation [449]. theta-représentations [449]. three [210, 154, 881, 1083, 122]. three-letter [210]. three-valued [154, 881]. threshold [556, 435, 657]. thresholds [590]. tight [55]. tile [937]. Tiling [448, 1050, 1154, 523]. Time [709, 791, 527, 978, 402, 971, 107, 77, 700, 113, 1111, 581, 754, 335, 493, 542, 350, 370, 1027, 236, 940, 1197, 292, 81, 1127, 411, 73, 121, 867, 1075, 639, 1122, 895, 202, 799, 758, 553, 433, 502, 471]. Time-space [709, 1111, 754]. timed [1000, 604]. times [1055]. Tokyo [913]. tolerance [1028, 868]. tolerant [1029]. Tolerating [1033]. tool [107]. Tools [858, 396]. Top [422, 1149]. Top-down [422, 1149]. Topological [41, 334, 746, 726]. topologically [873]. topology [420]. toposes [143]. torus [1154]. total [667, 223, 1175]. tours [988, 1061]. Towers [925]. trace [1138, 415, 508, 398, 1029]. trace-based [1029]. traces [77, 753, 884, 979]. tracking [1064, 1130]. trade [229]. trade-offs [229]. tradeoffs [709]. traffic [972]. transaction [1124, 118]. transactions [117, 241, 459]. Transducers [1139, 157, 682, 1149, 505, 1048, 639, 422]. transduction [122]. transductions [218, 995, 512, 15]. transform [170]. Transformation [697, 118, 200, 106, 680, 703, 718, 1028, 626, 305, 695]. Transformational [366, 37]. Transformations [136, 1182, 104, 443, 965, 705, 1038]. transformers [322]. Transforming [625, 68]. transforms [944]. Transition [855, 1073, 593, 1176, 478]. transitions [146]. translations [112]. transmission [554]. transposition [231, 685]. trapezes [937]. Traps [571]. treatment [854]. Tree [682, 859, 1133, 743, 535, 1007, 699, 17, 157, 260, 200, 1149, 657, 1049, 1005, 639, 998, 124, 422, 801]. tree-decomposable [699]. Trees [46, 282, 827, 431, 141, 635, 1014, 829, 1198, 648, 847, 835, 983, 55, 286, 504, 997, 434, 275, 749, 650, 288, 98, 596, 68]. triangles [748]. triangulated [545]. tries [652]. trivial [255]. tropical [1134]. true [761]. truth [800]. Tseitin [543]. Turing [608, 231, 685, 755, 538, 1011, 310, 1155, 1054, 292, 492, 553, 502]. turned [962]. tutorial [716, 715]. twice [844]. Two [466, 1198, 405, 474, 164, 306, 1086, 941, 70, 526, 1108, 1013, 1081, 125, 1062, 549, 747, 981, 742, 1154, 96, 506, 616, 422]. two-dimensional [1108, 981]. two-edge [1062]. two-quantifier [506]. two-way [164, 422]. Type [902, 860, 65, 359, 959, 224, 928, 34, 450, 547, 663, 1187, 762, 137, 325, 911, 615, 579, 849, 458, 1035]. type-theoretical [911]. typed [728, 719, 716, 904, 421, 908, 921, 1185]. types [960, 1189, 249, 220, 558, 358, 665, 464, 370, 562, 533, 116, 138, 287, 638, 959, 632].

Ulam [473]. ultimate [225]. ultra [532]. Unambiguity [658]. unambiguous [222]. unary [1085]. unavoidable [60]. unbounded [531, 290]. unboundedness [756]. undecidability [669]. Undecidable [1149, 317, 255, 647]. Undirected [745]. Unfold [305, 106, 626]. Unfold/fold [305, 106, 626]. Unfolding [627, 917]. Unification [1001, 752, 294, 779, 317, 1072, 212]. unified [1187, 480]. Uniform [1159, 161, 1099, 620, 854, 61]. uniformity [180]. unimodal [1117]. union [994, 1079]. Unique [670, 1136]. uniquely [938]. Unit [1105, 930]. Unit-cost [1105]. Universal [771, 562, 688, 967, 707, 789, 241, 660, 940, 492]. universality [1054, 869]. universally [421]. universes [359]. Unrestricted [427, 1091]. unsolvability [557]. unstratifiable [150]. untopological [299]. update [854, 117, 459, 692, 1060]. Upper [524, 932, 55, 1146]. upward [208]. urn [1082]. use [967]. used [836, 167]. useful [38, 515]. Usefulness [338]. user [1064]. Using [811, 387, 173, 1069, 556, 829, 629, 1053, 34, 147, 577, 236, 536, 381, 1178]. Uspensky [1066].

V [192, 1122]. V. [1066]. valued [154, 881, 138, 1145]. valuedness [639]. values [1106, 59]. variable [664, 214]. variable-free [664]. variables [747, 921]. variant [850]. variation [408]. Variations [1180, 1051]. Varieties [525, 941, 271]. variety [1143]. various [183]. VAS [465]. vector [981, 189, 521]. vectors [1013]. verification [336, 866, 198, 362]. version [164]. versus [1105, 634, 427, 1102, 227, 870, 1096]. vertex [594, 276]. vertex-disjoint [276]. very [655, 904, 850]. via [744, 779, 1072, 966, 1080, 853, 618]. view [769, 511]. viewpoint [652, 288]. views [188]. VII [564]. Virus [977]. Voronoi [147]. vs [661].

waiting [971]. waiting-time [971]. walking [422]. was [836]. way [164, 1047, 356, 211, 694, 285, 350, 289, 870, 1100, 422]. Weak [179, 325, 1113, 429, 1023, 266, 504]. weak-connected [1023]. Weakly [247, 893, 757]. Weighted [107, 861, 74, 1059, 545, 1016, 386, 692]. Weights [1175]. well [498, 780, 963, 215, 8, 624, 719, 917, 479, 915]. well-behaved [498]. well-formed [215]. well-founded [780, 963, 624, 917, 915]. well-quasi-orderings [8]. well-typed [719]. where [669]. whether [742]. Which [960, 279, 883, 77, 317, 416]. Whitney [843]. whole [345]. whose [487]. Winograd [1088]. within [1123]. without [848, 236, 1168, 947, 273, 1104, 383, 25]. Word [77, 175, 1159, 943, 784, 895, 447, 544]. Words [671, 1131, 673, 687, 1109, 872, 43, 792, 264, 210, 1006, 706, 676, 311, 678, 653, 228, 742, 355, 690, 596, 1193]. Working [697]. Workshop [26, 476, 695, 721, 767, 509, 148, 1112]. world [857, 777, 915, 296]. worlds [153]. Wythoff [111]. WZ [836].

yield [988]. Young [832]. Yves [927].

zag [164]. zero [269, 274, 416]. zero-divisor-free [416]. zero-knowledge [269]. zerotesting [274]. zeta [177]. zig [164]. zig-zag [164]. zigzag [97, 683].

References

Wagner:1990:APP

[1] E. G. Wagner. Algebras, polynomials and programs. Theoretical Computer Science, 70(1):3–34, January 15, 1990. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bainbridge:1990:FP

[2] E. S. Bainbridge, P. J. Freyd, A. Scedrov, and P. J. Scott. Functorial polymorphism. Theoretical Computer Science, 70(1):35–64, January 15, 1990. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Barr:1990:FPC

[3] M. Barr. Fixed points in Cartesian closed categories. Theoretical Computer Science, 70(1):65–72, January 15, 1990. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bloom:1990:NGT

[4] S. L. Bloom. A note on guarded theories. Theoretical Computer Science, 70(1):73–83, January 15, 1990.

CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Mulry:1990:CFP

[5] P. S. Mulry. Categorical fixed point semantics. Theoretical Computer Science, 70(1):85–97, January 15, 1990. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Nielson:1990:FCM

[6] H. R. Nielson and F. Nielson. Functional completeness of the mixed λcalculus and combinatory logic. Theoretical Computer Science, 70(1):99– 126, January 15, 1990. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Pasztor:1990:RPD

[7] A. Pasztor. Recursive programs and denotational semantics in absolute logics of programs. Theoretical Computer Science, 70(1):127–150, January 15, 1990. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bottinger:1990:STD

[8] C. Bottinger. On Scott’s thesis for domains of information and well-quasi-orderings. Theoretical Computer Science, 70(1):151–158, January 15, 1990. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Wells:1990:GCS

[9] C. Wells. A generalization of the concept of sketch. Theoretical Computer Science, 70(1):159–178, January 15, 1990. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Tiuryn:1990:FPF

[10] J. Tiuryn and D. B. Benson. Fixed points in free process algebras. II. Theoretical Computer Science, 70(2):179– 192, January 26, 1990. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Longo:1990:CTC

[11] G. Longo and E. Moggi. A categorytheoretic characterization of functional completeness. Theoretical Computer Science, 70(2):193–211, January 26, 1990. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Senizergues:1990:CDC

[12] G. Senizergues. A characterisation of deterministic context-free languages by means of right-congruences. Theoretical Computer Science, 70(2):213– 232, January 26, 1990. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Jung:1990:CCC

[13] A. Jung. Cartesian closed categories of algebraic CPOs. Theoretical Computer Science, 70(2):233–250, January 26, 1990. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gambosi:1990:LPP

[14] G. Gambosi, J. Nesetril, and M. Talamo. On locally presented posets. Theoretical Computer Science, 70(2):251– 260, January 26, 1990. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Terlutte:1990:CRT

[15] A. Terlutte. Cyclic rational transductions and polynomials of rational functions. Theoretical Computer Science, 70(2):261–271, January 26, 1990. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ehrenfeucht:1990:TSC

[16] A. Ehrenfeucht and G. Rozenberg. Theory of 2-structures. I. clans, basic subclasses, and morphisms. Theoretical Computer Science, 70(3):277– 303, February 15, 1990. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ehrenfeucht:1990:TSI

[17] A. Ehrenfeucht and G. Rozenberg. Theory of 2-structures. II. representation through labeled tree families. Theoretical Computer Science, 70(3):305– 342, February 15, 1990. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ehrenfeucht:1990:PHS

[18] A. Ehrenfeucht and G. Rozenberg. Primitivity is hereditary for 2-structures. Theoretical Computer Science, 70(3):343–358, February 15, 1990. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Aggarwal:1990:CCP

[19] A. Aggarwal, A. K. Chandra, and M. Snir. Communication complexity of PRAMs. Theoretical Computer Science, 71(1):3–28, March 13, 1990. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Culik:1990:NTP

[20] K. Culik, II. New techniques for proving the decidability of equivalence problems. Theoretical Computer Science, 71(1):29–45, March 13, 1990. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gruska:1990:SSP

[21] J. Gruska. Synthesis, structure and power of systolic computations. Theoretical Computer Science, 71(1):47– 77, March 13, 1990. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hartmanis:1990:NDS

[22] J. Hartmanis. New developments in structural complexity theory. Theoretical Computer Science, 71(1):79– 93, March 13, 1990. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kruskal:1990:CTE

[23] C. P. Kruskal, L. Rudolph, and M. Snir. A complexity theory of efficient parallel algorithms. Theoretical Computer Science, 71(1):95–132, March 13, 1990. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Thiagarajan:1990:SBA

[24] P. S. Thiagarajan. Some behavioural aspects of net theory. Theoretical Computer Science, 71(1):133–153, March 13, 1990. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Weijland:1990:SLP

[25] W. P. Weijland. Semantics for logic programs without occur check. Theoretical Computer Science, 71(1):155– 174, March 13, 1990. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:1990:IFS

[26] Anonymous. INFORMATIKA 88. 2nd French-Soviet Workshop on Methods of Compilation and Program Construction. Theoretical Computer Science, 71 (2):??, March 30, 1990. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Arcangeli:1990:PPP

[27] J.-P. Arcangeli and C. Pomian. Principles of plasma pattern and alternative structure compilation. Theoretical Computer Science, 71(2):177– 191, March 30, 1990. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Billaud:1990:SOD

[28] M. Billaud. Simple operational and denotational semantics for Prolog with cut. Theoretical Computer Science, 71 (2):193–208, March 30, 1990. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bulyonkov:1990:MCC

[29] M. A. Bulyonkov. Mixed computation and compilation: new approaches to old problems. Theoretical Computer Science, 71(2):209–226, March 30, 1990. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Galmiche:1990:CSA

[30] D. Galmiche. Constructive system for automatic program synthesis. Theoretical Computer Science, 71(2):227– 239, March 30, 1990. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Penjam:1990:CAM

[31] J. Penjam. Computational and attribute models of formal languages. Theoretical Computer Science, 71(2): 241–264, March 30, 1990. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Sabelfeld:1990:ADF

[32] V. K. Sabelfeld. An algorithm deciding functional equivalence in a new class of program schemes. Theoretical Computer Science, 71(2):265– 279, March 30, 1990. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Senizergues:1990:SDP

[33] G. Senizergues. Some decision problems about controlled rewriting systems. Theoretical Computer Science, 71(3):281–346, April 10, 1990. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ehrig:1990:CDT

[34] H. Ehrig, F. Parisi-Presicce, P. Boehm, C. Rieckhoff, C. Dimitrovici, and M. Grosse-Rhode. Combining data type and recursive process specifications using projection algebras. Theoretical Computer Science, 71(3):347– 380, April 10, 1990. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dulucq:1990:FSS

[35] S. Dulucq and D. Gouyou-Beauchamps. On the factors of the Sturmian sequences. Theoretical Computer Science, 71(3):381–400, April 10, 1990. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gibbons:1990:OEC

[36] A. Gibbons and W. Rytter. Optimally edge-colouring outerplanar graphs is in NC. Theoretical Computer Science, 71 (3):401–411, April 10, 1990. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Manes:1990:TCI

[37] E. G. Manes. A transformational characterization of if-then-else. Theoretical Computer Science, 71(3):413– 417, April 10, 1990. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chrobak:1990:DSU

[38] M. Chrobak, T. Szymacha, and A. Krawczyk. A data structure useful for finding Hamiltonian cycles. Theoretical Computer Science, 71(3):419– 424, April 10, 1990. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Skandalis:1990:NOR

[39] K. Skandalis. Nonrecursiveness of the operations on real numbers. Theoretical Computer Science, 71(3):425– 429, April 10, 1990. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Sakarovitch:1990:F

[40] J. Sakarovitch. Foreword. Theoretical Computer Science, 72(1):1– ??, April 25, 1990. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Budach:1990:TIC

[41] L. Budach. Topological invariants of classification problems. Theoretical Computer Science, 72(1):3–26, April 25, 1990. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hashiguchi:1990:ILT

[42] K. Hashiguchi. Improved limitedness theorems on finite automata with distance functions. Theoretical Computer Science, 72(1):27–38, April 25, 1990. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Carpi:1990:NRW

[43] A. Carpi and A. de Luca. Nonrepetitive words relative to a rewriting system. Theoretical Computer Science, 72(1):39–53, April 25, 1990. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Restivo:1990:CLC

[44] A. Restivo. Codes and local constraints. Theoretical Computer Science, 72(1):55–64, April 25, 1990. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Simon:1990:FFF

[45] I. Simon. Factorization forests of finite height. Theoretical Computer Science, 72(1):65–94, April 25, 1990. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:1990:CCA

[46] Anonymous. CAAP ’88 — 13th colloque sur les arbres en algèbre et en programmation. (French) [Colloquium on trees in algebra and programming]. Theoretical Computer Science, 72(2-3):??, May 23, 1990. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dauchet:1990:E

[47] M. Dauchet and J. L. Remy. Editorial. Theoretical Computer Science, 72(2-3): 95–??, May 23, 1990. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ausiello:1990:DMD

[48] G. Ausiello, U. Nanni, and G. F. Italiano. Dynamic maintenance of directed hypergraphs. Theoretical Computer Science, 72(2-3):97–117, May 23, 1990. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Engeler:1990:CDF

[49] E. Engeler. Combinatory differential fields. Theoretical Computer Science, 72(2-3):119–131, May 23, 1990. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Echahed:1990:CNS

[50] R. Echahed. On completeness of narrowing strategies. Theoretical Computer Science, 72(2-3):133–146, May 23, 1990. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Francon:1990:ADA

[51] J. Francon, B. Randrianarimanana, and R. Schott. Analysis of dynamic algorithms in Knuth’s model. Theoretical Computer Science, 72(2-3):147– 167, May 23, 1990. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gnaedig:1990:ECO

[52] I. Gnaedig, C. Kirchner, and H. Kirchner. Equational completion in ordersorted algebras. Theoretical Computer Science, 72(2-3):169–202, May 23, 1990. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gorrieri:1990:AA

[53] R. Gorrieri, S. Marchetti, and U. Montanari. A^{2}CCS: atomic actions for CCS. Theoretical Computer Science, 72(2-3):203–223, May 23, 1990. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kennaway:1990:ITR

[54] R. Kennaway. Implementing term rewrite languages in Dactl. Theoretical Computer Science, 72(2-3):225– 249, May 23, 1990. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Klein:1990:TUB

[55] R. Klein and D. Wood. A tight upper bound for the path length of AVL trees. Theoretical Computer Science, 72(2-3): 251–264, May 23, 1990. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Larsen:1990:PSS

[56] K. G. Larsen. Proof systems for satisfiability in Hennessy–Milner Logic with recursion. Theoretical Computer Science, 72(2-3):265–288, May 23, 1990. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Jacopini:1990:RPC

[57] G. Jacopini and G. Sontacchi. Reversible parallel computation: an evolving space-model. Theoretical Computer Science, 73(1):1–46, June 08, 1990. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kroger:1990:IAT

[58] F. Kroger. On the interpretability of arithmetic in temporal logic. Theoretical Computer Science, 73(1):47– 60, June 08, 1990. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lavault:1990:EAM

[59] C. Lavault. Exact average message complexity values for distributed election on bidirectional rings of processors. Theoretical Computer Science, 73 (1):61–79, June 08, 1990. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Varricchio:1990:FFM

[60] S. Varricchio. Factorizations of free monoids and unavoidable regularities. Theoretical Computer Science, 73(1): 81–89, June 08, 1990. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Jerrum:1990:FUG

[61] M. Jerrum and A. Sinclair. Fast uniform generation of regular graphs. Theoretical Computer Science, 73(1):91– 100, June 08, 1990. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Huwig:1990:NIC

[62] H. Huwig and A. Poigne. A note on inconsistencies caused by fixpoints in a Cartesian closed category. Theoretical Computer Science, 73(1):101– 112, June 08, 1990. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ganzinger:1990:F

[63] H. Ganzinger. Foreword. Theoretical Computer Science, 73(2):119– ??, June 22, 1990. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Aiken:1990:TCB

[64] A. Aiken. A theory of compactionbased parallelization. Theoretical Computer Science, 73(2):121–154, June 22, 1990. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fuh:1990:TIS

[65] You-Chin C. Fuh and P. Mishra. Type inference with subtypes. Theoretical Computer Science, 73(2):155– 175, June 22, 1990. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Giegerich:1990:CSI

[66] R. Giegerich. Code selection by inversion of order-sorted derivors. Theoretical Computer Science, 73(2):177– 211, June 22, 1990. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Horwitz:1990:ARQ

[67] S. Horwitz. Adding relational query facilities to software development environments. Theoretical Computer Science, 73(2):213–230, June 22, 1990. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Wadler:1990:DTP

[68] P. Wadler. Deforestation: transforming programs to eliminate trees. Theoretical Computer Science, 73(2):231– 248, June 22, 1990. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Beigel:1990:BIR

[69] R. Beigel. Bi-immunity results for cheatable sets. Theoretical Computer Science, 73(3):249–263, July 22, 1990. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chen:1990:DBO

[70] Jian er Chen. The difference between one tape and two tapes: with respect to reversal complexity. Theoretical Computer Science, 73(3):265–278, July 22, 1990. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hoffmann:1990:RNG

[71] J. Hoffmann and M. G. Main. Results on NLC grammars with one-letter terminal alphabets. Theoretical Computer Science, 73(3):279–294, July 22, 1990. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kim:1990:CDR

[72] Changwook Kim. Complexity and decidability for restricted classes of picture languages. Theoretical Computer Science, 73(3):295–311, July 22, 1990. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Robson:1990:STB

[73] J. M. Robson. Strong time bounds: non-computable bounds and a hierarchy theorem. Theoretical Computer Science, 73(3):313–317, July 22, 1990. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bertossi:1990:SMW

[74] A. A. Bertossi, F. Luccio, E. Lodi, and L. Pagli. String matching with weighted errors. Theoretical Computer Science, 73(3):319–328, July 22, 1990. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Head:1990:SSM

[75] T. Head. The set of strings mapped into a submonoid by iterates of a morphism. Theoretical Computer Science, 73(3):329–333, July 22, 1990. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Schmitt:1990:HAI

[76] W. Schmitt. Hopf algebras and identities in free partially commutative monoids. Theoretical Computer Science, 73(3):335–340, July 22, 1990. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Diekert:1990:WPT

[77] V. Diekert. Word problems over traces which are solvable in linear time. Theoretical Computer Science, 74(1): 3–18, July 30, 1990. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fussell:1990:SAP

[78] D. Fussell and R. Thurimella. Successive approximation in parallel graph algorithms. Theoretical Computer Science, 74(1):19–35, July 30, 1990. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gavilanes-Franco:1990:FOL

[79] A. Gavilanes-Franco and F. Lucio-Carrasco. A first order logic for partial functions. Theoretical Computer Science, 74(1):37–69, July 30, 1990. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Jancar:1990:DTL

[80] P. Jancar. Decidability of a temporal logic problem for Petri nets. Theoretical Computer Science, 74(1):71– 93, July 30, 1990. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Preparata:1990:DPP

[81] F. P. Preparata and R. Tamassia. Dynamic planar point location with optimal query time. Theoretical Computer Science, 74(1):95–114, July 30, 1990. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Szepietowski:1990:IDN

[82] A. Szepietowski. If deterministic and nondeterministic space complexities are equal for log log n, then they are also equal for log n. Theoretical Computer Science, 74(1):115–119, July 30, 1990. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gastin:1990:AMD

[83] P. Gastin. An asynchronous model for distributed systems. Theoretical Computer Science, 74(2):121–162, August 21, 1990. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gecseg:1990:ARP

[84] F. Gecseg and H. Jurgensen. Automata represented by products of soliton automata. Theoretical Computer Science, 74(2):163–181, August 21, 1990. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Oommen:1990:DOE

[85] B. J. Oommen, E. R. Hansen, and J. I. Munro. Deterministic optimal and expedient move-to-rear list organizing strategies. Theoretical Computer Science, 74(2):183–197, August 21, 1990. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Okulicka:1990:PC

[86] F. Okulicka. On priority in COSY. Theoretical Computer Science, 74(2): 199–216, August 21, 1990. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hartmanis:1990:RMS

[87] J. Hartmanis and L. A. Hemachandra. Robust machine saccept easy sets. Theoretical Computer Science, 74(2):217– 225, August 21, 1990. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bezem:1990:CRR

[88] M. Bezem. Completeness of resolution revisited. Theoretical Computer Science, 74(2):227–237, August 21, 1990. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Huttel:1990:SCM

[89] H. Huttel. SnS can be modally characterized. Theoretical Computer Science, 74(2):239–248, August 21, 1990. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kummer:1990:EPF

[90] M. Kummer. An easy priority-free proof of a theorem of friedberg. Theoretical Computer Science, 74(2):249– 251, August 21, 1990. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Zhang:1990:PNM

[91] Jingzhong Zhang, Lu Yang, and M. Deng. The parallel numerical method of mechanical theorem proving. Theoretical Computer Science, 74 (3):253–271, August 28, 1990. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Casas:1990:ASC

[92] R. Casas, M.-I. Fernandez-Camacho, and J.-M. Steyaert. Algebraic simplification in computer algebra: an analysis of bottom-up algorithms. Theoretical Computer Science, 74(3):273– 298, August 28, 1990. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

He:1990:EAE

[93] Xin He. An efficient algorithm for edge coloring planar graphs with delta colors. Theoretical Computer Science, 74 (3):299–312, August 28, 1990. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Babai:1990:LBC

[94] L. Babai, P. Pudlak, V. Rodl, and E. Szemeredi. Lower bounds to the complexity of symmetry Boolean functions. Theoretical Computer Science, 74(3):313–323, August 28, 1990. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hertrampf:1990:RAM

[95] U. Hertrampf. Relations among modclasses. Theoretical Computer Science, 74(3):325–328, August 28, 1990. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Romeuf:1990:PAS

[96] J.-F. Romeuf. A polynomial algorithm for solving systems of two linear Diophantine equations. Theoretical Computer Science, 74(3):329– 340, August 28, 1990. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anselmo:1990:ZCT

[97] M. Anselmo. On zigzag codes and their decidability. Theoretical Computer Science, 74(3):341–354, August 28, 1990. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Szymanski:1990:CAR

[98] J. Szymanski. On the complexity of algorithms on recursive trees. Theoretical Computer Science, 74(3):355– 361, August 28, 1990. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:1990:ICF

[99] Anonymous. International Conference on Fifth Generation Computer Systems 1988. Theoretical Computer Science, 75(1-2):??, September 25, 1990. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Milner:1990:IOC

[100] R. Milner. Interpreting one concurrent calculus in another. Theoretical Computer Science, 75(1-2):3–13, September 25, 1990. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

deBakker:1990:CMS

[101] J. W. de Bakker and J. N. Kok. Comparative metric semantics for concurrent prolog. Theoretical Computer Science, 75(1-2):15–43, September 25, 1990. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Falaschi:1990:FFP

[102] M. Falaschi and G. Levi. Finite failures and partial computations in concurrent logic languages. Theoretical Computer Science, 75(1-2):45–66, September 25, 1990. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Murakami:1990:DSF

[103] M. Murakami. A declarative semantics of flat guarded Horn clauses for programs with perpetual processes. Theoretical Computer Science, 75(1-2):67– 83, September 25, 1990. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Holldobler:1990:CET

[104] S. Holldobler. Conditional equational theories and complete sets of transformations. Theoretical Computer Science, 75(1-2):85–110, September 25, 1990. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dershowitz:1990:RCE

[105] N. Dershowitz and M. Okada. A rationale for conditional equational programming. Theoretical Computer Science, 75(1-2):111–138, September 25, 1990. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kawamura:1990:PSE

[106] T. Kawamura and T. Kanamori. Preservation of stronger equivalence in unfold/fold logic program transformation. Theoretical Computer Science, 75(1-2):139–156, September 25, 1990. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Devienne:1990:WGT

[107] P. Devienne. Weighted graphs: a tool for studying the halting problem and time complexity in term rewriting systems and logic programming. Theoretical Computer Science, 75(1-2):157– 215, September 25, 1990. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Degano:1990:POS

[108] P. Degano, R. De Nicola, and U. Montanari. A partial ordering semantics for CCS. Theoretical Computer Science, 75(3):223–262, October 01, 1990. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Katz:1990:IST

[109] S. Katz and D. Peled. Interleaving set temporal logic. Theoretical Computer Science, 75(3):263–287, October 01, 1990. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Droste:1990:NDI

[110] M. Droste and R. Gobel. Nondeterministic information systems and their domains. Theoretical Computer Science, 75(3):289–309, October 01, 1990. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Blass:1990:SGF

[111] U. Blass and A. S. Fraenkel. The Sprague-Grundy function for Wythoff’s game. Theoretical Computer Science, 75(3):311–333, October 01, 1990. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Allender:1990:DTE

[112] E. Allender and C. Wilson. Downward translations of equality. Theoretical Computer Science, 75(3):335– 346, October 01, 1990. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Du:1990:MMF

[113] J. Du, J. Y.-T. Leung, and G. H. Young. Minimizing mean flow time with release time constraint. Theoretical Computer Science, 75(3):347– 355, October 01, 1990. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lengauer:1990:BNF

[114] T. Lengauer and K. W. Wagner. The binary network flow problem is logspace complete for P. Theoretical Computer Science, 75(3):357–363, October 01, 1990. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bonner:1990:HDC

[115] A. J. Bonner. Hypothetical datalog: complexity and expressibility. Theoretical Computer Science, 76(1):3– 51, October 31, 1990. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ohori:1990:STD

[116] A. Ohori. Semantics of types for database objects. Theoretical Computer Science, 76(1):53–91, October 31, 1990. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Karabeg:1990:PUT

[117] D. Karabeg and V. Vianu. Parallel update transactions. Theoretical Computer Science, 76(1):93–114, October 31, 1990. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lipeck:1990:TDI

[118] U. W. Lipeck. Transformation of dynamic integrity constraints into transaction specifications. Theoretical Computer Science, 76(1):115–142, October 31, 1990. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dong:1990:DDP

[119] Guozhu Dong and S. Ginsburg. On the decomposition of datalog program mappings. Theoretical Computer Science, 76(1):143–177, October 31, 1990. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kok:1990:CCC

[120] J. N. Kok and J. J. M. M. Rutten. Contractions in comparing concurrency semantics. Theoretical Computer Science, 76(2-3):179–222, November 21, 1990. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Sakakibara:1990:LCF

[121] Y. Sakakibara. Learning context-free grammars from structural data in polynomial time. Theoretical Computer Science, 76(2-3):223–242, November 21, 1990. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Timmerman:1990:TSR

[122] E. Timmerman. The three subfamilies of rational omega-languages closed under omega-transduction. Theoretical Computer Science, 76(2-3):243–250, November 21, 1990. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Weil:1990:PLC

[123] P. Weil. Products of languages with counter. Theoretical Computer Science, 76(2-3):251–260, November 21, 1990.

CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Tirri:1990:CTC

[124] S. Tirri. The congruence theory of closure properties of regular tree languages. Theoretical Computer Science, 76(2-3):261–271, November 21, 1990. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hashiguchi:1990:ERE

[125] K. Hashiguchi and H. Yoo. Extended regular expressions of star degree at most two. Theoretical Computer Science, 76(2-3):273–284, November 21, 1990. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Petit:1990:DSA

[126] A. Petit. Distribution and synchronized automata. Theoretical Computer Science, 76(2-3):285–308, November 21, 1990. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Yamasaki:1990:RES

[127] S. Yamasaki. Recursion equation sets computing logic programs. Theoretical Computer Science, 76(2-3):309–322, November 21, 1990. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Jiang:1990:CTA

[128] T. Jiang. On the complexity of 1tape ATMs and off-line 1-tape ATMs running in constant reversals. Theoretical Computer Science, 76(2-3):323– 330, November 21, 1990. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gannett:1990:OPS

[129] E. Gannett, S. C. Kothari, and H.- C. Yen. On optimal parallelization of sorting networks. Theoretical Computer Science, 76(2-3):331–341, November 21, 1990. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Sarnath:1990:PCG

[130] R. Sarnath and Xin He. A P-complete graph partition problem. Theoretical Computer Science, 76(2-3):343–351, November 21, 1990. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:1990:ICA

[131] Anonymous. International Conference on Algebraic Methodology and Software Technology, AMAST. Theoretical Computer Science, 77(1-2):??, December 07, 1990. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Rus:1990:E

[132] T. Rus. Editorial. Theoretical Computer Science, 77(1-2):1–??, December 07, 1990. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bradley:1990:ALD

[133] L. Bradley. Abstract language design. Theoretical Computer Science, 77(1-2): 5–26, December 07, 1990. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ehrig:1990:CPD

[134] H. Ehrig, W. Fey, H. Hansen, M. Lowe, D. Jacobs, and F. Parisi-Presicce. Compatibility problems in the development of algebraic module specifications. Theoretical Computer Science, 77(1-2):27–71, December 07, 1990. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Even:1990:CSA

[135] S. Even and D. A. Schmidt. Categorysorted algebra-based action semantics. Theoretical Computer Science, 77(1-2): 73–95, December 07, 1990. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Janicki:1990:TSS

[136] R. Janicki and T. Muldner. Transformations of sequential specifications into concurrent specifications by synchronization guards. Theoretical Computer Science, 77(1-2):97–129, December 07, 1990. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Manca:1990:ETL

[137] V. Manca, A. Salibra, and G. Scollo. Equational type logic. Theoretical Computer Science, 77(1-2):131–159, December 07, 1990. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Pigozzi:1990:DTM

[138] D. Pigozzi. Data types over multiplevalued logics. Theoretical Computer Science, 77(1-2):161–194, December 07, 1990. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Wagner:1990:ASL

[139] E. G. Wagner. An algebraically specified language for data directed design. Theoretical Computer Science, 77(1-2):195–219, December 07, 1990. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Toda:1990:PRL

[140] S. Toda. Positive relativizations for log space computability. Theoretical Computer Science, 77(3):221–235, December 15, 1990. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bozapalidis:1990:ECF

[141] S. Bozapalidis. Effective constructions on formal series of trees. Theoretical Computer Science, 77(3):237–247, December 15, 1990. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Smith:1990:IAP

[142] C. H. Smith and M. Velauthapillai. On the inference of approximate programs. Theoretical Computer Science, 77(3): 249–266, December 15, 1990. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kawahara:1990:PCB

[143] Y. Kawahara. Pushout-complements and basic concepts of grammars in toposes. Theoretical Computer Science, 77(3):267–289, December 15, 1990. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Atzeni:1990:EOQ

[144] P. Atzeni and E. P. F. Chan. Efficient and optimal query answering on independent schemes. Theoretical Computer Science, 77(3):291–308, December 15, 1990. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dowling:1990:FPL

[145] W. F. Dowling and R. Kline. The fixed points of logic programs with Herbrand base N. Theoretical Computer Science, 77(3):309–319, December 15, 1990. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fahmy:1990:APN

[146] H. M. A. Fahmy. Analysis of Petri nets by partitioning: splitting transitions. Theoretical Computer Science, 77(3): 321–330, December 15, 1990. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Elbaz:1990:CVD

[147] M. Elbaz and J.-C. Spehner. Construction of Voronoi diagrams in the plane by using maps. Theoretical Computer Science, 77(3):331–343, December 15, 1990. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:1991:WDD

[148] Anonymous. Workshop on Deductive Database Theory. Theoretical Computer Science, 78(1):??, January 21, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bidoit:1991:NRB

[149] N. Bidoit. Negation in rule-based database languages: a survey. Theoretical Computer Science, 78(1):3– 83, January 21, 1991. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bidoit:1991:NDU

[150] N. Bidoit and C. Froidevaux. Negation by default and unstratifiable logic programs. Theoretical Computer Science, 78(1):85–112, January 21, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Royer:1991:SID

[151] V. Royer. The semantics of incomplete databases as an expression of preferences. Theoretical Computer Science, 78(1):113–136, January 21, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Abiteboul:1991:FPD

[152] S. Abiteboul and E. Simon. Fundamental properties of deterministic and nondeterministic extensions of datalog. Theoretical Computer Science, 78(1): 137–158, January 21, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Abiteboul:1991:RQS

[153] S. Abiteboul, P. Kanellakis, and G. Grahne. On the representation and querying of sets of possible worlds. Theoretical Computer Science, 78(1): 159–187, January 21, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Delahaye:1991:PTV

[154] J. P. Delahaye and V. Thibau. Programming in three-valued logic. Theoretical Computer Science, 78(1):189– 216, January 21, 1991. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Courcelle:1991:RQC

[155] B. Courcelle. Recursive queries and context-free graph grammars. Theoretical Computer Science, 78(1):217– 244, January 21, 1991. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Demolombe:1991:ESN

[156] R. Demolombe. An efficient strategy for non-Horn deductive databases. Theoretical Computer Science, 78(1): 245–259, January 21, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Engelfriet:1991:MTT

[157] J. Engelfriet and H. Vogler. Modular tree transducers. Theoretical Computer Science, 78(2):267–303, January 31, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Downey:1991:CCH

[158] R. Downey. On computational complexity and honest polynomial degrees. Theoretical Computer Science, 78(2): 305–317, January 31, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Konig:1991:GFP

[159] R. Konig. Graphs and free partially commutative monoids. Theoretical Computer Science, 78(2):319– 346, January 31, 1991. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Harju:1991:EPM

[160] T. Harju and J. Karhumaki. The equivalence problem of multitape finite automata. Theoretical Computer Science, 78(2):347–355, January 31, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Barrington:1991:NSL

[161] D. A. M. Barrington and J. Corbett. A note on some languages in uniform ACC^{0}. Theoretical Computer Science, 78(2):357–362, January 31, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Baeza-Yates:1991:SS

[162] R. A. Baeza-Yates. Searching subsequences. Theoretical Computer Science, 78(2):363–376, January 31, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Burosch:1991:NDC

[163] G. Burosch, J. Demetrovics, G. O. H. Katona, Kleitman, D. J., and A. A. Sapozhenko. On the number of databases and closure operations. Theoretical Computer Science, 78(2):377– 381, January 31, 1991. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anselmo:1991:ZZP

[164] M. Anselmo. The zig-zag power series: a two-way version of the * operator. Theoretical Computer Science, 79(1):3– 24, February 21, 1991. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bertoni:1991:RFP

[165] A. Bertoni, D. Bruschi, and M. Goldwurm. Ranking and formal power series. Theoretical Computer Science, 79 (1):25–35, February 21, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Flajolet:1991:AAC

[166] P. Flajolet, B. Salvy, and P. Zimmermann. Automatic average-case analysis of algorithms. Theoretical Computer Science, 79(1):37–109, February 21, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Krob:1991:SEF

[167] D. Krob. Some examples of formal series used in non-commutative algebra. Theoretical Computer Science, 79(1): 111–135, February 21, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kuich:1991:ALG

[168] W. Kuich. Automata and languages generalized to omega-continuous semirings. Theoretical Computer Science, 79 (1):137–150, February 21, 1991. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hespel:1991:AND

[169] C. Hespel and G. Jacob. Approximation of nonlinear dynamic systems by rational series. Theoretical Computer Science, 79(1):151–162, February 21, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Minh:1991:ET

[170] V. Hoang Ngoc Minh. Evaluation transform. Theoretical Computer Science, 79(1):163–177, February 21, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Leroux:1991:CAN

[171] P. Leroux and X. G. Viennot. A combinatorial approach to nonlinear functional expansions: an introduction with an example. Theoretical Computer Science, 79(1):179–193, February 21, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Oussous:1991:CMM

[172] N. E. Oussous. Computation, on Macsyma, of the minimal differential representation of noncommutative polynomials. Theoretical Computer Science, 79(1):195–207, February 21, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Delest:1991:EPU

[173] M. Delest. Enumeration of polyominoes using Macsyma. Theoretical Computer Science, 79(1):209–226, February 21, 1991. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Duchamp:1991:OPF

[174] G. Duchamp. Orthogonal projection onto the free Lie algebra. Theoretical Computer Science, 79(1):227–239, February 21, 1991. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Koseleff:1991:WGF

[175] P.-V. Koseleff. Word games in free Lie algebras: several bases and formulas. Theoretical Computer Science, 79(1): 241–256, February 21, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Rotella:1991:SPG

[176] F. Rotella. Shuffle product of generating series. Theoretical Computer Science, 79(1):257–261, February 21, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Honkala:1991:AGZ

[177] J. Honkala. On algebraic generalized zeta functions of formal power series. Theoretical Computer Science, 79(1): 263–273, February 21, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Vaandrager:1991:DES

[178] F. W. Vaandrager. Determinism to (event structure isomorphism=step sequence equivalence). Theoretical Computer Science, 79(2):275–294, February 27, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Holden:1991:WLT

[179] M. Holden. Weak logic theory. Theoretical Computer Science, 79(2):295– 321, February 27, 1991. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Jaoua:1991:RRM

[180] A. Jaoua, A. Mili, N. Boudriga, and J. L. Durieux. Regularity of relations: a measure of uniformity. Theoretical Computer Science, 79(2):323– 339, February 27, 1991. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Szalas:1991:SAC

[181] A. Szalas. On strictly arithmetical completeness in logics of programs. Theoretical Computer Science, 79(2): 341–355, February 27, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Stoughton:1991:IPO

[182] A. Stoughton. Interdefinability of parallel operations in PCF. Theoretical Computer Science, 79(2):357–358, February 27, 1991. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Jung:1991:DPC

[183] A. Jung. The dependent product construction in various categories of domains. Theoretical Computer Science, 79(2):359–363, February 27, 1991. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Koletsos:1991:PLC

[184] G. Koletsos. Polymorphic lambda calculus: the Church–Rosser property. Theoretical Computer Science, 79(2): 365–371, February 27, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Reisig:1991:PNA

[185] W. Reisig. Petri nets and algebraic specifications. Theoretical Computer Science, 80(1):1–34, March 21, 1991.

CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gecseg:1991:PA

[186] F. Gecseg and H. Jurgensen. On α_{0}-ν_{1}-products of automata. Theoretical Computer Science, 80(1):35– 51, March 21, 1991. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Heilbrunner:1991:ERB

[187] S. Heilbrunner and L. Schmitz. An efficient recognizer for the Boolean closure of context-free languages. Theoretical Computer Science, 80(1):53– 75, March 21, 1991. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Howell:1991:GLV

[188] R. R. Howell, L. E. Rosier, and Hsu-Chun Yen. Global and local views of state fairness. Theoretical Computer Science, 80(1):77–104, March 21, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Powell:1991:COV

[189] P. Powell and Viet Ngo. Complexity of optimal vector code generation. Theoretical Computer Science, 80(1):105– 115, March 21, 1991. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hofer:1991:ESG

[190] G. Hofer. Experiments and stability in group automata. Theoretical Computer Science, 80(1):117–120, March 21, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Andreka:1991:STP

[191] H. Andreka, I. Nemeti, and I. Sain. On the strength of temporal proofs. Theoretical Computer Science, 80(2):125– 151, March 29, 1991. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Courcelle:1991:MSO

[192] B. Courcelle. The monadic secondorder logic of graphs. V. on closing the gap between definability and recognizability. Theoretical Computer Science, 80(2):153–202, March 29, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hemachandra:1991:SPE

[193] L. A. Hemachandra, A. Hoene, D. Siefkes, and P. Young. On sets polynomially enumerable by iteration. Theoretical Computer Science, 80(2):203– 225, March 29, 1991. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Iwanowski:1991:TAS

[194] S. Iwanowski. Testing approximate symmetry in the plane is NP-hard. Theoretical Computer Science, 80(2): 227–262, March 29, 1991. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Olderog:1991:CCP

[195] E.-R. Olderog. Correctness of concurrent processes. Theoretical Computer Science, 80(2):263–288, March 29, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ranjan:1991:SBC

[196] D. Ranjan, R. Chang, and J. Hartmanis. Space bounded computations: review and new separation results. Theoretical Computer Science, 80(2):289– 302, March 29, 1991. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Steffen:1991:FCC

[197] B. Steffen and J. Knoop. Finite constants: characterizations of a new decidable set of constants. Theoretical Computer Science, 80(2):303– 318, March 29, 1991. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Szczepanska:1991:HLV

[198] D. Szczepanska. A Hoare-like verification system for a language with an exception handling mechanism. Theoretical Computer Science, 80(2):319– 335, March 29, 1991. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Wiedermann:1991:CES

[199] J. Wiedermann. On the computational efficiency of symmetric neural networks. Theoretical Computer Science, 80(2):337–345, March 29, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fulop:1991:CCD

[200] Z. Fulop and S. Vagvolgyi. A complete classification of deterministic root-to-frontier tree transformation classes. Theoretical Computer Science, 81(1): 1–15, April 22, 1991. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Boe:1991:BAT

[201] J.-M. Boe. The boxes (automata theory). Theoretical Computer Science, 81 (1):17–34, April 22, 1991. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Tang:1991:PTR

[202] Shouwen Tang and R. V. Book. Polynomial-time reducibilities and ‘almost all’ oracle sets. Theoretical Computer Science, 81(1):35–47, April 22, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dubois:1991:CNS

[203] O. Dubois. Counting the number of solutions for instances of satisfiability. Theoretical Computer Science, 81(1): 49–64, April 22, 1991. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dubois:1991:PAS

[204] O. Dubois and J. Carlier. Probabilistic approach to the satisfiability problem. Theoretical Computer Science, 81 (1):65–75, April 22, 1991. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Piotrow:1991:CCP

[205] M. Piotrow. On the complexity of counting in the polynomial hierarchy. Theoretical Computer Science, 81(1): 77–95, April 22, 1991. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Amir:1991:FPS

[206] A. Amir and G. M. Landau. Fast parallel and serial multidimensional approximate array matching. Theoretical Computer Science, 81(1):97–115, April 22, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Matus:1991:AFD

[207] F. Matus. Abstract functional dependency structures. Theoretical Computer Science, 81(1):117–126, April 22, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lutz:1991:UMS

[208] J. H. Lutz. An upward measure separation theorem. Theoretical Computer Science, 81(1):127–135, April 22, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Leung:1991:LTF

[209] H. Leung. Limitedness theorem on finite automata with distance functions: an algebraic proof. Theoretical Computer Science, 81(1):137–145, April 22, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cori:1991:NPA

[210] R. Cori and M. R. Formisano. On the number of partially Abelian squarefree words on a three-letter alphabet. Theoretical Computer Science, 81(1): 147–153, April 22, 1991. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hartmanis:1991:OWF

[211] J. Hartmanis. One-way functions and the nonisomorphism of NP-complete sets. Theoretical Computer Science, 81 (1):155–163, April 22, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kapur:1991:SU

[212] D. Kapur, D. Musser, P. Narendran, and J. Stillman. Semi-unification. Theoretical Computer Science, 81(2):169– 187, April 30, 1991. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Khadilkar:1991:PCS

[213] S. N. Khadilkar and S. Biswas. Padding, commitment and selfreducibility. Theoretical Computer Science, 81(2):189–199, April 30, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Pingali:1991:ANL

[214] K. Pingali and K. Ekanadham. Accumulators: new logic variable abstractions for functional languages. Theoretical Computer Science, 81(2):201– 221, April 30, 1991. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bhatta:1991:RRW

[215] K. S. H. S. R. Bhatta and H. Karnick. A resolution rule for well-formed formulae. Theoretical Computer Science, 81(2):223–235, April 30, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bodlaender:1991:NLB

[216] H. L. Bodlaender. New lower bound techniques for distributed leader finding and other problems on rings of processors. Theoretical Computer Science, 81(2):237–256, April 30, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lenz:1991:CCQ

[217] K. Lenz and I. Wegener. The conjunctive complexity of quadratic Boolean functions. Theoretical Computer Science, 81(2):257–268, April 30, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Asveld:1991:AGB

[218] P. R. J. Asveld. Abstract grammars based on transductions. Theoretical Computer Science, 81(2):269– 288, April 30, 1991. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dassow:1991:CPC

[219] J. Dassow. On the connectedness of pictures in chain code picture languages. Theoretical Computer Science, 81(2):289–294, April 30, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cheung:1991:RDT

[220] To-Yat Cheung and Yunzhou Zhu. Recognizing different types of beta-cycles in a database scheme. Theoretical Computer Science, 81(2):295–304, April 30, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Csuhaj-Varju:1991:PCR

[221] E. Csuhaj-Varju and J. Kelemen. On the power of cooperation: a regular representation of recursively enumerable languages. Theoretical Computer Science, 81(2):305–310, April 30, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chytil:1991:PRU

[222] M. Chytil, M. Crochemore, B. Monien, and W. Rytter. On the parallel recognition of unambiguous context-free languages. Theoretical Computer Science, 81(2):311–316, April 30, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Megiddo:1991:TFE

[223] N. Megiddo and C. H. Papadimitriou. On total functions, existence theorems and computational complexity. Theoretical Computer Science, 81(2):317– 324, April 30, 1991. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Weihrauch:1991:TCC

[224] K. Weihrauch and C. Kreitz. Type 2 computational complexity of functions on Cantor’s space. Theoretical Computer Science, 82(1):1–18, May 22, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lando:1991:PUP

[225] B. Lando. Periodicity and ultimate periodicity of D0L systems. Theoretical Computer Science, 82(1):19– 33, May 22, 1991. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hebrard:1991:ADE

[226] J.-J. Hebrard. An algorithm for distinguishing efficiently bit-strings by their subsequences. Theoretical Computer Science, 82(1):35–49, May 22, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ko:1991:AVN

[227] K.-I. Ko. On adaptive versus nonadaptive bounded query machines. Theoretical Computer Science, 82(1):51– 69, May 22, 1991. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Mignosi:1991:NFS

[228] F. Mignosi. On the number of factors of Sturmian words. Theoretical Computer Science, 82(1):71–84, May 22, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Snir:1991:SDT

[229] M. Snir. Size-depth trade-offs for monotone arithmetic circuits. Theoretical Computer Science, 82(1):85– 93, May 22, 1991. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Engelfriet:1991:NSG

[230] J. Engelfriet, G. Leih, and G. Rozenberg. Nonterminal separation in graph grammars. Theoretical Computer Science, 82(1):95–111, May 22, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dietzfelbinger:1991:CMT

[231] M. Dietzfelbinger, W. Maass, and G. Schnitger. The complexity of matrix transposition on one-tape off-line Turing machines. Theoretical Computer Science, 82(1):113–129, May 22, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Salomaa:1991:DSE

[232] K. Salomaa and S. Yu. Decidability of structural equivalence of E0L grammars. Theoretical Computer Science, 82(1):131–139, May 22, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Robson:1991:LRR

[233] J. M. Robson. An O(T log T) reduction from RAM computations to satisfiability. Theoretical Computer Science, 82(1):141–149, May 22, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Calude:1991:DSS

[234] C. Calude and G. Istrate. Determining and stationary sets for some classes of partial recursive functions. Theoretical Computer Science, 82(1):151– 155, May 22, 1991. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Palko:1991:ACM

[235] V. Palko, O. Sykora, and I. Vrto. Area complexity of merging. Theoretical Computer Science, 82(1):157– 163, May 22, 1991. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Leo:1991:GCF

[236] J. M. I. M. Leo. A general contextfree parsing algorithm running in linear time on every LR(k) grammar without using lookahead. Theoretical Computer Science, 82(1):165–176, May 22, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bauderon:1991:IHB

[237] M. Bauderon. Infinite hypergraphs. I. basic properties. Theoretical Computer Science, 82(2):177–214, May 31, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Germano:1991:CFG

[238] G. M. Germano and S. Mazzanti. Closure functions and general iterates as reflectors. Theoretical Computer Science, 82(2):215–252, May 31, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Abadi:1991:ERM

[239] M. Abadi and L. Lamport. The existence of refinement mappings. Theoretical Computer Science, 82(2):253– 284, May 31, 1991. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Baeten:1991:RPD

[240] J. C. M. Baeten and J. A. Bergtstra. Recursive process definitions with the state operator. Theoretical Computer Science, 82(2):285–302, May 31, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lukassen:1991:FFI

[241] A. Lukassen and G. Vossen. A formal framework for independence with respect to transactions in the universal relation model. Theoretical Computer Science, 82(2):303–327, May 31, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Szwast:1991:HS

[242] W. Szwast. On Horn spectra. Theoretical Computer Science, 82(2):329– 339, May 31, 1991. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Howell:1991:TFT

[243] R. R. Howell, L. E. Rosier, and Hsu-Chun Yen. A taxonomy of fairness and temporal logic problems for Petri nets. Theoretical Computer Science, 82(2): 341–372, May 31, 1991. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Denis:1991:TAS

[244] F. Denis and J.-P. Delahaye. Is there an axiomatic semantics for standard pure prolog? Theoretical Computer Science, 82(2):373–388, May 31, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Curien:1991:AFE

[245] P.-L. Curien. An abstract framework for environment machines. Theoretical Computer Science, 82(2):389– 402, May 31, 1991. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Badouel:1991:GR

[246] E. Badouel and P. Darondeau. On guarded recursion. Theoretical Computer Science, 82(2):403–408, May 31, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hoofman:1991:WEM

[247] R. Hoofman. Weakly expressive models for Hoare logic. Theoretical Computer Science, 82(2):409–418, May 31, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Breazu-Tannen:1991:PRC

[248] V. Breazu-Tannen and J. Gallier. Polymorphic rewriting conserves algebraic strong normalization. Theoretical Computer Science, 83(1):3– 28, June 21, 1991. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cardone:1991:RTF

[249] F. Cardone. Recursive types for fun. Theoretical Computer Science, 83(1): 29–56, June 21, 1991. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Colson:1991:APR

[250] L. Colson. About primitive recursive algorithms. Theoretical Computer Science, 83(1):57–69, June 21, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dershowitz:1991:RRR

[251] N. Dershowitz, S. Kaplan, and D. A. Plaisted. Rewrite, rewrite, rewrite, rewrite, rewrite. Theoretical Computer Science, 83(1):71–96, June 21, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Manna:1991:CTP

[252] Z. Manna and A. Pnueli. Completing the temporal picture (logic). Theoretical Computer Science, 83(1):97– 130, June 21, 1991. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Parisi-Presicce:1991:FRB

[253] F. Parisi-Presicce. Foundations of rulebased design of modular systems. Theoretical Computer Science, 83(1):131– 155, June 21, 1991. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Winskel:1991:NMC

[254] G. Winskel. A note on model checking the modal ν-calculus. Theoretical Computer Science, 83(1):157–167, June 21, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Luchkardt:1991:NFU

[255] H. Luchkardt. New formally undecidable propositions: non-trivial lower bounds on proof complexity and related theorems. Theoretical Computer Science, 83(2):169–188, June 28, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Mauro:1991:DLS

[256] D. Mauro. Derived linear systems of context-free grammars. Theoretical Computer Science, 83(2):189–203, June 28, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

deBoer:1991:PDP

[257] M. J. M. de Boer, A. Lindenmayer, and Z. Tuza. A periodic division pattern that cannot be generated by D0L systems. Theoretical Computer Science, 83(2):205–218, June 28, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Labhalla:1991:CCC

[258] S. Labhalla. Computational complexity of continuous fraction real number development. Theoretical Computer Science, 83(2):219–235, June 28, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Seki:1991:SLC

[259] S. Seki and Y. Kobuchi. On standard locally catenative L schemes. Theoretical Computer Science, 83(2):237– 248, June 28, 1991. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fachini:1991:NCC

[260] E. Fachini, A. M. Schettini, G. Resta, and D. Sangiorgi. Nonacceptability criteria and closure properties for the class of languages accepted by binary systolic tree automata. Theoretical Computer Science, 83(2):249–260, June 28, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kuster:1991:HPF

[261] G. Kuster. On the Hurwitz product of formal power series and automata. Theoretical Computer Science, 83(2): 261–273, June 28, 1991. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Azar:1991:PCM

[262] Y. Azar. Parallel comparison merging of many-ordered lists. Theoretical Computer Science, 83(2):275– 285, June 28, 1991. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kundu:1991:MSR

[263] S. Kundu. Minimal strings in a regular language with respect to a partial order on the alphabet. Theoretical Computer Science, 83(2):287–300, June 28, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cohen-Chesnot:1991:EPT

[264] J. Cohen-Chesnot. On the expressive power of temporal logic for infinite words. Theoretical Computer Science, 83(2):301–312, June 28, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hemachandra:1991:KCC

[265] L. A. Hemachandra and G. Wechsung. Kolmogorov characterizations of complexity classes. Theoretical Computer Science, 83(2):313–322, June 28, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Mostowski:1991:HWA

[266] A. W. Mostowski. Hierarchies of weak automata and weak monadic formulas. Theoretical Computer Science, 83(2): 323–335, June 28, 1991. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Watanabe:1991:IC

[267] O. Watanabe. On the p-isomorphism conjecture. Theoretical Computer Science, 83(2):337–343, June 28, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Beauquier:1991:LS

[268] D. Beauquier and J.-E. Pin. Languages and scanners. Theoretical Computer Science, 84(1):3–21, July 22, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Brassard:1991:CRP

[269] G. Brassard, C. Crepeau, and M. Yung. Constant-round perfect zero-knowledge computationally convincing protocols. Theoretical Computer Science, 84(1): 23–52, July 22, 1991. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bruyere:1991:MCB

[270] V. Bruyere. Maximal codes with bounded deciphering delay. Theoretical Computer Science, 84(1):53– 76, July 22, 1991. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chazelle:1991:SES

[271] B. Chazelle, H. Edelsbrunner, L. J. Guibas, and M. Sharir. A singly exponential stratification scheme for real semi-algebraic varieties and its applications. Theoretical Computer Science, 84(1):77–105, July 22, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gambosi:1991:PFD

[272] G. Gambosi, E. Nardelli, and M. Talamo. A pointer-free data structure for merging heaps and min-max heaps. Theoretical Computer Science, 84(1): 107–126, July 22, 1991. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Papadimitriou:1991:SPM

[273] C. H. Papadimitriou and M. Yannakakis. Shortest paths without a map.

Theoretical Computer Science, 84(1): 127–150, July 22, 1991. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Clausen:1991:ZTI

[274] M. Clausen, A. Dress, J. Grabmeier, and M. Karpinski. On zero-testing and interpolation of k-sparse multivariate polynomials over finite fields. Theoretical Computer Science, 84(2):151– 164, July 29, 1991. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Saoudi:1991:GAI

[275] A. Saoudi. Generalized automata on infinite trees and Muller-McNaughton’s theorem. Theoretical Computer Science, 84(2):165–177, July 29, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Moran:1991:OCC

[276] S. Moran and Y. Wolfstahl. Optimal covering of cacti by vertex-disjoint paths. Theoretical Computer Science, 84(2):179–197, July 29, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Beigel:1991:BQS

[277] R. Beigel. Bounded queries to SAT and the Boolean hierarchy. Theoretical Computer Science, 84(2):199– 223, July 29, 1991. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Becker:1991:CFM

[278] B. Becker and U. Sparmann. Computations over finite monoids and their test complexity. Theoretical Computer Science, 84(2):225–250, July 29, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Rupert:1991:WKS

[279] C. P. Rupert. Which Kleene semigroups are finite? Theoretical Computer Science, 84(2):251–264, July 29, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Maurer:1991:BDC

[280] H. A. Maurer, A. Salomaa, and D. Wood. Bounded delay L codes. Theoretical Computer Science, 84(2): 265–279, July 29, 1991. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dassow:1991:DSA

[281] J. Dassow and H. Jurgensen. Deterministic soliton automata with a single exterior node. Theoretical Computer Science, 84(2):281–292, July 29, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Baldwin:1991:MT

[282] W. A. Baldwin and G. O. Strawn. Multidimensional trees. Theoretical Computer Science, 84(2):293–311, July 29, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Wang:1991:CSC

[283] Jie Wang. On p-creative sets and pcompletely creative sets. Theoretical Computer Science, 85(1):1–31, August 05, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Devolder:1991:FGB

[284] J. Devolder and I. Litovsky. Finitely generated bi omega-languages. Theoretical Computer Science, 85(1):33– 52, August 05, 1991. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ibarra:1991:PPO

[285] O. H. Ibarra, Tao Jiang, and Hui Wang. Parallel parsing on a one-way linear array of finite-state machines. Theoretical Computer Science, 85(1):53– 74, August 05, 1991. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lindell:1991:AFP

[286] S. Lindell. An analysis of fixedpoint queries on binary trees. Theoretical Computer Science, 85(1):75– 95, August 05, 1991. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Rodenburg:1991:ASD

[287] P. H. Rodenburg. Algebraic specifiability of data types with minimal computable parameters. Theoretical Computer Science, 85(1):97–116, August 05, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Szpankowski:1991:CDS

[288] W. Szpankowski. A characterization of digital search trees from the successful search viewpoint. Theoretical Computer Science, 85(1):117– 134, August 05, 1991. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kutylowski:1991:MOW

[289] M. Kutylowski. Multihead one-way finite automata. Theoretical Computer Science, 85(1):135–153, August 05, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Wegener:1991:CPF

[290] I. Wegener. The complexity of the parity function in unbounded fanin, unbounded depth circuits. Theoretical Computer Science, 85(1):155– 170, August 05, 1991. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cherubini:1991:QFA

[291] A. Cherubini, C. Citrini, S. C. Reghizzi, and D. Mandrioli. QRT FIFO automata, breadth-first grammars and their relations. Theoretical Computer Science, 85(1):171–203, August 05, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Michel:1991:NCL

[292] P. Michel. An NP-complete language accepted in linear time by a one-tape Turing machine. Theoretical Computer Science, 85(1):205–212, August 05, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cialdea:1991:RSF

[293] M. Cialdea. Resolution for some firstorder modal systems. Theoretical Computer Science, 85(2):213–229, August 12, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Doggaz:1991:CU

[294] N. Doggaz and C. Kirchner. Completion for unification. Theoretical Computer Science, 85(2):231–251, August 12, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Sopena:1991:HRC

[295] E. Sopena. Hypermap rewriting: a combinatorial approach. Theoretical Computer Science, 85(2):253–281, August 12, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Tennent:1991:CPW

[296] R. D. Tennent and J. K. Tobin. Continuations in possible-world semantics. Theoretical Computer Science, 85(2): 283–303, August 12, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kamper:1991:NPS

[297] J. Kamper. Nonuniform proof systems: a new framework to describe nonuniform and probabilistic complexity classes. Theoretical Computer Science, 85(2):305–331, August 12, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Goto:1991:PNN

[298] S. Goto. Proof normalization with nonstandard objects. Theoretical Computer Science, 85(2):333–351, August 12, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bethke:1991:CSU

[299] I. Bethke. Coherence spaces are untopological. Theoretical Computer Science, 85(2):353–357, August 12, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:1991:ICL

[300] Anonymous. 6th International Conference on Logic Programming (ICLP 89). Theoretical Computer Science, 86 (1):??, August 19, 1991. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

deBoer:1991:SMC

[301] F. S. de Boer, J. J. M. M. Rutten, J. N. Kok, and C. Palamidessi. Semantic models for concurrent logic languages. Theoretical Computer Science, 86(1):3– 33, August 19, 1991. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bol:1991:ALC

[302] R. N. Bol, K. R. Apt, and J. W. Klop. An analysis of loop checking mechanisms for logic programs. Theoretical Computer Science, 86(1):35– 79, August 19, 1991. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cavedon:1991:ALP

[303] L. Cavedon. Acyclic logic programs and the completeness of SLDNF-resolution. Theoretical Computer Science, 86(1): 81–92, August 19, 1991. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lobo:1991:SHD

[304] J. Lobo, A. Rajasekar, and J. Minker. Semantics of Horn and disjunctive logic programs. Theoretical Computer Science, 86(1):93–106, August 19, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Seki:1991:UFT

[305] H. Seki. Unfold/fold transformation of stratified programs. Theoretical Computer Science, 86(1):107– 139, August 19, 1991. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Averbuch:1991:CAM

[306] A. Averbuch, Z. Galil, and S. Winograd. Classification of all the minimal bilinear algorithms for computing the coefficients of the product of two polynomials modulo a polynomial. II. the algebra G(u)(u^{n}). Theoretical Computer Science, 86(2):143–203, September 02, 1991. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Spehner:1991:MMP

[307] J.-C. Spehner. Merging in maps and in pavings. Theoretical Computer Science, 86(2):205–232, September 02, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hashiguchi:1991:RCS

[308] K. Hashiguchi. Recognizable closures and submonoids of free partially commutative monoids. Theoretical Computer Science, 86(2):233–241, September 02, 1991. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chrobak:1991:POL

[309] M. Chrobak and D. Eppstein. Planar orientations with low out-degree and compaction of adjacency matrices. Theoretical Computer Science, 86(2): 243–266, September 02, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Krause:1991:SET

[310] M. Krause, C. Meinel, and S. Waack. Separating the eraser Turing machine classes L_{e}, NL_{e}, co-NL_{e} and P_{e}. Theoretical Computer Science, 86(2):267– 275, September 02, 1991. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gire:1991:ALB

[311] F. Gire and M. Nivat. Algebraic languages of biinfinite words. Theoretical Computer Science, 86(2):277–323, September 02, 1991. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bertoni:1991:CCN

[312] A. Bertoni, M. Goldwurm, and N. Sabadini. The complexity of computing the number of strings of given length in context-free languages. Theoretical Computer Science, 86(2):325– 342, September 02, 1991. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Seger:1991:ESI

[313] C.-J. Seger. On the existence of speed-independent circuits. Theoretical Computer Science, 86(2):343–364, September 02, 1991. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Capocelli:1991:DIS

[314] R. M. Capocelli, L. Gargano, and U. Vaccaro. Decoders with initial state invariance for multivalued encodings. Theoretical Computer Science, 86(2): 365–375, September 02, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Benedek:1991:LRF

[315] G. M. Benedek and A. Itai. Learnability with respect to fixed distributions. Theoretical Computer Science, 86(2): 377–389, September 02, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Stoltenberg-Hanssen:1991:AFP

[316] V. Stoltenberg-Hanssen and J. V. Tucker. Algebraic and fixed point equations over inverse limits of algebras. Theoretical Computer Science, 87(1):1– 24, September 16, 1991. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Farmer:1991:SSO

[317] W. M. Farmer. Simple second-order languages for which unification is undecidable. Theoretical Computer Science, 87(1):25–41, September 16, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Heibig:1991:CMN

[318] A. G. Heibig. Control machines: a new model of parallelism for compositional specifications and their effective compilation. Theoretical Computer Science, 87(1):43–80, September 16, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Balsters:1991:SCS

[319] H. Balsters and M. M. Fokkinga. Subtyping can have a simple semantics. Theoretical Computer Science, 87(1): 81–96, September 16, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fournier:1991:EDM

[320] R. Fournier and G. von Bochmann. The equivalence in the DCP model. Theoretical Computer Science, 87(1): 97–114, September 16, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hallnas:1991:PID

[321] L. Hallnas. Partial inductive definitions. Theoretical Computer Science, 87(1):115–142, September 16, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gardiner:1991:DRP

[322] P. Gardiner and C. Morgan. Data refinement of predicate transformers. Theoretical Computer Science, 87(1): 143–162, September 16, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Zierer:1991:RAD

[323] H. Zierer. Relation algebraic domain constructions. Theoretical Computer Science, 87(1):163–188, September 16, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Wilm:1991:DNP

[324] A. Wilm. Determinism and nondeterminism in PDL. Theoretical Computer Science, 87(1):189–202, September 16, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Mezghiche:1991:WCT

[325] M. Mezghiche. Weak completeness of type assignment in lambda-calculus models: a generalization of Hindley’s result. Theoretical Computer Science, 87(1):203–208, September 16, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Milner:1991:CRS

[326] R. Milner and M. Tofte. Conduction in relational semantics. Theoretical Computer Science, 87(1):209–220, September 16, 1991. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

McCune:1991:APF

[327] W. McCune and L. Wos. The absence and the presence of fixed point combinators. Theoretical Computer Science, 87(1):221–228, September 16, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Esik:1991:RHR

[328] Z. Esik. Results on homomorphic realization of automata by α_{0}-products. Theoretical Computer Science, 87(2): 229–249, September 23, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Diks:1991:OPC

[329] K. Diks and W. Rytter. On optimal parallel computations for sequences of brackets. Theoretical Computer Science, 87(2):251–262, September 23, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gyssens:1991:CBA

[330] M. Gyssens and D. Van Gucht. A comparison between algebraic query languages for flat and nested databases. Theoretical Computer Science, 87(2): 263–286, September 23, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Csuhaj-Varju:1991:BIG

[331] E. Csuhaj-Varju and J. Dassow. On bounded interpretations of grammar forms. Theoretical Computer Science, 87(2):287–313, September 23, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

deLuca:1991:FIC

[332] A. de Luca and S. Varricchio. Finiteness and iteration conditions for semigroups. Theoretical Computer Science, 87(2):315–327, September 23, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Nirmal:1991:MCE

[333] N. Nirmal and R. Rama. Machine characterization of (E0L-E0L) array languages. Theoretical Computer Science, 87(2):329–346, September 23, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Calude:1991:RTS

[334] C. Calude. Relativized topological size of sets of partial recursive functions. Theoretical Computer Science, 87(2): 347–352, September 23, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gavalda:1991:SRS

[335] R. Gavalda and J. L. Balcazar. Strong and robustly strong polynomial-time reducibilities to sparse sets. Theoretical Computer Science, 88(1):1–14, September 30, 1991. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Larsen:1991:PSC

[336] K. G. Larsen and B. Thomsen. Partial specifications and compositional verification. Theoretical Computer Science, 88(1):15–32, September 30, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Miyano:1991:CLF

[337] S. Miyano. Δ_{2}^{P} -complete lexicographically first maximal subgraph problems. Theoretical Computer Science, 88(1): 33–57, September 30, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Crochemore:1991:UKM

[338] M. Crochemore and W. Rytter. Usefulness of the Karp-Miller-Rosenberg algorithm in parallel computations on strings and arrays. Theoretical Computer Science, 88(1):59–82, September 30, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Garzon:1991:CGG

[339] M. Garzon and Y. Zalcstein. The complexity of Grigorchuk groups with application to cryptography. Theoretical Computer Science, 88(1):83–98, September 30, 1991. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cho:1991:FAA

[340] Sang Cho and Dung T. Huynh. Finiteautomaton aperiodicity is PSPACEcomplete. Theoretical Computer Science, 88(1):99–116, September 30, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Shankar:1991:GBR

[341] P. Shankar and B. S. Adiga. A graphbased regularity test for deterministic context-free languages. Theoretical Computer Science, 88(1):117–125, September 30, 1991. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Salomaa:1991:DAM

[342] A. Salomaa. A deterministic algorithm for modular knapsack problems. Theoretical Computer Science, 88(1):127– 138, September 30, 1991. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Engelfriet:1991:RCG

[343] J. Engelfriet. A regular characterization of graph languages definable in monadic second-order logic. Theoretical Computer Science, 88(1):139–150, September 30, 1991. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Aizawa:1991:GGP

[344] K. Aizawa and A. Nakamura. Graph grammars with path-controlled embedding. Theoretical Computer Science, 88(1):151–170, September 30, 1991.

CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Labhalla:1991:RRN

[345] S. Labhalla and H. Lombardi. Representation of real numbers by developments in whole bases and complexity. Theoretical Computer Science, 88(1): 171–182, September 30, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Khuller:1991:PGC

[346] S. Khuller and V. V. Vazirani. Planar graph coloring is not self-reducible, assuming P not=NP. Theoretical Computer Science, 88(1):183–189, September 30, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Seki:1991:MCF

[347] H. Seki, T. Matsumura, M. Fujii, and T. Kasami. On multiple context-free grammars. Theoretical Computer Science, 88(2):191–229, October 07, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Huckenbeck:1991:RAP

[348] U. Huckenbeck. A result about the power of geometric oracle machines. Theoretical Computer Science, 88(2): 231–251, October 07, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Peeva:1991:ERM

[349] K. Peeva. Equivalence, reduction and minimization of finite automata over semirings. Theoretical Computer Science, 88(2):269–285, October 07, 1991.

CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Inoue:1991:NRT

[350] K. Inoue, A. Ito, and I. Takanami. A note on real-time one-way alternating multicounter machines. Theoretical Computer Science, 88(2):287– 296, October 07, 1991. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bebjak:1991:SDN

[351] A. Bebjak and I. Stefanekova. Separation of deterministic, nondeterministic and alternating complexity classes. Theoretical Computer Science, 88(2): 297–311, October 07, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Baur:1991:ACR

[352] W. Baur. On the algebraic complexity of rational iteration procedures. Theoretical Computer Science, 88(2):313– 324, October 07, 1991. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Weber:1991:DAF

[353] A. Weber and H. Seidl. On the degree of ambiguity of finite automata. Theoretical Computer Science, 88(2):325– 349, October 07, 1991. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Yoo:1991:EAL

[354] H. Yoo and K. Hashiguchi. Extended automata-like regular expressions of star degree at most (2, 1). Theoretical Computer Science, 88(2):351– 363, October 07, 1991. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Seebold:1991:FMS

[355] P. Seebold. Fibonacci morphisms and Sturmian words. Theoretical Computer Science, 88(2):365–384, October 07, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ganzinger:1991:OSC

[356] H. Ganzinger. Order-sorted completion: the many-sorted way. Theoretical Computer Science, 89(1):3– 32, October 21, 1991. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Habel:1991:DBP

[357] A. Habel, H.-J. Kreowski, and W. Vogler. Decidable boundedness problems for sets of graphs generated by hyperedge-replacement. Theoretical Computer Science, 89(1):33– 62, October 21, 1991. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hanus:1991:HCP

[358] M. Hanus. Horn clause programs with polymorphic types: semantics and resolution. Theoretical Computer Science, 89(1):63–106, October 21, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Harper:1991:TCU

[359] R. Harper and R. Pollack. Type checking with universes. Theoretical Computer Science, 89(1):107–136, October 21, 1991. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Pfenning:1991:MPC

[360] F. Pfenning and P. Lee. Metacircularity in the polymorphic λ-calculus. Theoretical Computer Science, 89(1):137– 159, October 21, 1991. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Stirling:1991:LMC

[361] C. Stirling and D. Walker. Local model checking in the modal mu-calculus. Theoretical Computer Science, 89(1): 161–177, October 21, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Vissers:1991:SSD

[362] C. A. Vissers, G. Scollo, M. van Sinderen, and E. Brinksma. Specification styles in distributed systems design and verification. Theoretical Computer Science, 89(1):179–206, October 21, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Krob:1991:CSB

[363] D. Krob. Complete systems of Brational identities. Theoretical Computer Science, 89(2):207–343, October 28, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Knuth:1991:TPC

[364] D. E. Knuth. Theory and practice (computer science). Theoretical Computer Science, 90(1):1–15, November 11, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Pottosin:1991:APO

[365] I. V. Pottosin. Analysis of program optimization possibilities and further development. Theoretical Computer Science, 90(1):17–36, November 11, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kasyanov:1991:TAP

[366] V. Kasyanov. Transformational approach to program concretization. Theoretical Computer Science, 90(1):37– 46, November 11, 1991. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bulyonkov:1991:PEM

[367] M. A. Bulyonkov. From partial evaluation to mixed computation. Theoretical Computer Science, 90(1):47– 60, November 11, 1991. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Futamura:1991:EGP

[368] Y. Futamura, K. Nogi, and A. Takano. Essence of generalized partial computation. Theoretical Computer Science, 90(1):61–79, November 11, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Itkin:1991:AMC

[369] V. E. Itkin. An algebra of mixed computation. Theoretical Computer Science, 90(1):81–93, November 11, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Jones:1991:SST

[370] N. D. Jones. Static semantics, types, and binding time analysis. Theoretical Computer Science, 90(1):95–118, November 11, 1991. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Turski:1991:PB

[371] W. M. Turski. Prescribing behaviours. Theoretical Computer Science, 90(1): 119–125, November 11, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

deBakker:1991:FDC

[372] J. W. de Bakker and J. H. A. Warmerdam. Four domains for concurrency. Theoretical Computer Science, 90(1): 127–149, November 11, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cherkasova:1991:ACN

[373] L. A. Cherkasova and V. E. Kotov. An algebra of concurrent nondeterministic processes. Theoretical Computer Science, 90(1):151–170, November 11, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Mazurkiewicz:1991:CS

[374] A. Mazurkiewicz, A. Rabinovich, and B. A. Trakhtenbrot. Connectedness and synchronization. Theoretical Computer Science, 90(1):171–184, November 11, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Tyugu:1991:HOD

[375] E. Tyugu. Higher order dataflow schemas. Theoretical Computer Science, 90(1):185–198, November 11, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Barzdin:1991:RCA

[376] J. M. Barzdin and G. J. Barzdin. Rapid construction of algebraic axioms from samples. Theoretical Computer Science, 90(1):199–208, November 11, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Blikle:1991:CES

[377] A. Blikle, A. Tarlecki, and M. Thorup. On conservative extensions of syntax in system development. Theoretical Computer Science, 90(1):209– 233, November 11, 1991. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hoare:1991:TDC

[378] C. A. R. Hoare. A theory for the derivation of combinational C-MOS circuit designs. Theoretical Computer Science, 90(1):235–251, November 11, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Nepejvoda:1991:BBC

[379] N. N. Nepejvoda. A bridge between constructive logic and computer programming. Theoretical Computer Science, 90(1):253–270, November 11, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Rus:1991:ACC

[380] T. Rus. Algebraic construction of compilers. Theoretical Computer Science, 90(2):271–308, November 25, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Tatsuta:1991:PSU

[381] M. Tatsuta. Program synthesis using realizability. Theoretical Computer Science, 90(2):309–353, November 25, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Saito:1991:DER

[382] T. Saito. Dynamics of equivalence relations in automata networks. Theoretical Computer Science, 90(2):355– 367, November 25, 1991. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Toyama:1991:HPE

[383] Y. Toyama. How to prove equivalence of term rewriting systems without induction. Theoretical Computer Science, 90(2):369–390, November 25, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dong:1991:LCO

[384] Guozhu Dong and S. Ginsburg. Localizable constraints for object histories. Theoretical Computer Science, 90(2): 391–432, November 25, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Vakarelov:1991:MLK

[385] D. Vakarelov. Modal logics for knowledge representation systems. Theoretical Computer Science, 90(2):433–456, November 25, 1991. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Srinivasan:1991:EAM

[386] A. Srinivasan and C. P. Rangan. Efficient algorithms for the minimum weighted dominating clique problem on permutation graphs. Theoretical Computer Science, 91(1):1–21, December 09, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Buneman:1991:UPG

[387] P. Buneman, A. Jung, and A. Ohori. Using powerdomains to generalize relational databases. Theoretical Computer Science, 91(1):23–55, December 09, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Culik:1991:ESF

[388] K. Culik, II and S. Dube. An efficient solution of the firing mob problem. Theoretical Computer Science, 91 (1):57–69, December 09, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Arfi:1991:POH

[389] M. Arfi. Polynomial operations and hierarchies of concatenation. Theoretical Computer Science, 91(1):71–84, December 09, 1991. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hashiguchi:1991:ADR

[390] K. Hashiguchi. Algorithms for determining relative inclusion star height and inclusion star height. Theoretical Computer Science, 91(1):85–100, December 09, 1991. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kinber:1991:CSS

[391] E. B. Kinber. On complete sets of samples for generalized regular expressions.

Theoretical Computer Science, 91(1): 101–117, December 09, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Moews:1991:SGB

[392] D. Moews. Sums of games born on Days 2 and 3. Theoretical Computer Science, 91(1):119–128, December 09, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kearney:1991:EFP

[393] P. Kearney and J. Staples. An extensional fixed-point semantics for nondeterministic data flow. Theoretical Computer Science, 91(2):129–179, December 23, 1991. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Knijnenburg:1991:MPD

[394] P. M. W. Knijnenburg and J. van Leeuwen. On models for propositional dynamic logic. Theoretical Computer Science, 91(2):181–203, December 23, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Vogler:1991:ENP

[395] W. Vogler. Executions: a new partialorder semantics of Petri nets. Theoretical Computer Science, 91(2):205– 238, December 23, 1991. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Tarlecki:1991:SFA

[396] A. Tarlecki, R. M. Burstall, and J. A. Goguen. Some fundamental algebraic tools for the semantics of computation. 3. indexed categories. Theoretical Computer Science, 91(2):239–264, December 23, 1991. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Zhang:1991:CEA

[397] Wenhui Zhang. Cut elimination and automatic proof procedures. Theoretical Computer Science, 91(2):265–284, December 23, 1991. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Rozoy:1991:EST

[398] B. Rozoy and P. S. Thiagarajan. Event structures and trace monoids. Theoretical Computer Science, 91(2):285– 313, December 23, 1991. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:1992:CPM

[399] Anonymous. Combinatorial Pattern Matching School. Theoretical Computer Science, 92(1):??, January 06, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Crochemore:1992:FSI

[400] M. Crochemore. Foreword to the Special Issue on Selected Papers of the Combinatorial Pattern Matching School, Paris, July 1990. Theoretical Computer Science, 92(1):1–??, January 06, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Apostolico:1992:FLS

[401] A. Apostolico, S. Browne, and C. Guerra. Fast linear-space computations of longest common subsequences. Theoretical Computer Science, 92(1):

3–17, January 06, 1992. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Baeza-Yates:1992:ART

[402] R. A. Baeza-Yates and M. Regnier. Average running time of the Boyer– Moore–Horspool algorithm. Theoretical Computer Science, 92(1):19–31, January 06, 1992. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Crochemore:1992:SMO

[403] M. Crochemore. String-matching on ordered alphabets. Theoretical Computer Science, 92(1):33–47, January 06, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Galil:1992:DPC

[404] Z. Galil and K. Park. Dynamic programming with convexity, concavity, and sparsity. Theoretical Computer Science, 92(1):49–76, January 06, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hashiguchi:1992:TRS

[405] K. Hashiguchi and K. Yamada. Two recognizable string-matching problems over free partially commutative monoids. Theoretical Computer Science, 92(1):77–86, January 06, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Iliopoulos:1992:OAC

[406] C. S. Iliopoulos and W. F. Smyth. Optimal algorithms for computing the canonical form of a circular string. Theoretical Computer Science, 92(1):87– 105, January 06, 1992. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kim:1992:ASM

[407] J. Y. Kim and J. Shawe-Taylor. An approximate string-matching algorithm. Theoretical Computer Science, 92(1): 107–117, January 06, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lecroq:1992:VBM

[408] T. Lecroq. A variation on the Boyer– Moore algorithm. Theoretical Computer Science, 92(1):119–144, January 06, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Neraud:1992:SMI

[409] J. Neraud and M. Crochemore. A string matching interpretation of the equation x^{m}y^{n} = z^{p}. Theoretical Computer Science, 92(1):145–164, January 06, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Quong:1992:FAC

[410] R. W. Quong. Fast average-case pattern matching by multiplexing sparse tables. Theoretical Computer Science, 92(1):165–179, January 06, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Revuz:1992:MAD

[411] D. Revuz. Minimisation of acyclic deterministic automata in linear time. Theoretical Computer Science, 92(1): 181–189, January 06, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ukkonen:1992:ASM

[412] E. Ukkonen. Approximate stringmatching with q-grams and maximal matches. Theoretical Computer Science, 92(1):191–211, January 06, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Zipstein:1992:DCF

[413] M. Zipstein. Data compression with factor automata. Theoretical Computer Science, 92(1):213–221, January 06, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ehrenfeucht:1992:AS

[414] A. Ehrenfeucht and G. Rozenberg. Angular 2-structures. Theoretical Computer Science, 92(2):227–248, January 20, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Goldwurm:1992:PEN

[415] M. Goldwurm. Probabilistic estimation of the number of prefixes of a trace. Theoretical Computer Science, 92(2): 249–268, January 20, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hebisch:1992:GSS

[416] U. Hebisch and H. J. Weinert. Generalized semigroup semirings which are zero-divisor-free or multiplicatively left-cancellative. Theoretical Computer Science, 92(2):269–289, January 20, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bitar:1992:PCF

[417] J. Bitar and E. Goles. Parallel chip firing games on graphs. Theoretical Computer Science, 92(2):291–300, January 20, 1992. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lutz:1992:IRO

[418] J. H. Lutz. On independent random oracles. Theoretical Computer Science, 92(2):301–307, January 20, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hemachandra:1992:SCC

[419] L. A. Hemachandra and R. S. Rubinstein. Separating complexity classes with tally oracles. Theoretical Computer Science, 92(2):309–318, January 20, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Edelsbrunner:1992:ACP

[420] H. Edelsbrunner, L. Guibas, J. Pach, R. Pollack, Seidel, R., and M. Sharir. Arrangements of curves in the planetopology, combinatorics and algorithms. Theoretical Computer Science, 92(2):319–336, January 20, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kfoury:1992:EPF

[421] A. J. Kfoury, J. Tiuryn, and P. Urzyczyn. On the expressive power of finitely typed and universally polymorphic recursive procedures. Theoretical Computer Science, 93(1):1–41, February 03, 1992. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Vagvolgyi:1992:TTT

[422] S. Vagvolgyi. Top-down tree transducers with two-way tree walking lookahead. Theoretical Computer Science, 93(1):43–74, February 03, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Revesz:1992:LOE

[423] G. E. Revesz. A list-oriented extension of the lambda-calculus satisfying the Church–Rosser theorem. Theoretical Computer Science, 93(1):75– 89, February 03, 1992. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Harrison:1992:NAR

[424] P. G. Harrison and H. Khoshnevisan. A new approach to recursion removal. Theoretical Computer Science, 93(1): 91–113, February 03, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Subrahmanian:1992:PDD

[425] V. S. Subrahmanian. Paraconsistent disjunctive deductive databases. Theoretical Computer Science, 93(1):115– 141, February 03, 1992. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Zhang:1992:SN

[426] Guo-Qiang Zhang. Stable neighbourhoods. Theoretical Computer Science, 93(1):143–157, February 03, 1992. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Goerdt:1992:URV

[427] A. Goerdt. Unrestricted resolution versus N-resolution. Theoretical Computer Science, 93(1):159–167, February 03, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Peladeau:1992:LDS

[428] P. Peladeau. Logically defined subsets of N^{k}. Theoretical Computer Science, 93(2):169–183, February 17, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hiraishi:1992:SCW

[429] K. Hiraishi and A. Ichikawa. On structural conditions for weak persistency and semilinearity of Petri nets. Theoretical Computer Science, 93(2):185– 199, February 17, 1992. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Louchard:1992:DAD

[430] G. Louchard, B. Randrianarimanana, and R. Schott. Dynamic algorithms in D. E. Knuth’s model: a probabilistic analysis. Theoretical Computer Science, 93(2):201–225, February 17, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bonizzoni:1992:AIT

[431] P. Bonizzoni and G. Mauri. On automata on infinite trees. Theoretical Computer Science, 93(2):227–244, February 17, 1992. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kedem:1992:OPA

[432] Z. M. Kedem and K. V. Palem. Optimal parallel algorithms for forest and term matching. Theoretical Computer Science, 93(2):245–264, February 17, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Toda:1992:RRP

[433] S. Toda. Restricted relativizations of probabilistic polynomial time. Theoretical Computer Science, 93(2):265– 277, February 17, 1992. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ramanan:1992:TOA

[434] P. Ramanan. Testing the optimality of alphabetic trees. Theoretical Computer Science, 93(2):279–301, February 17, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Boyar:1992:AMC

[435] J. Boyar, G. Frandsen, and C. Sturtivant. An arithmetic model of computation equivalent to threshold circuits. Theoretical Computer Science, 93(2): 303–319, February 17, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dehornoy:1992:CPN

[436] P. Dehornoy. A criterion for proving noetherianity of a relation. Theoretical Computer Science, 93(2):321– 325, February 17, 1992. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Herbst:1992:KFP

[437] T. Herbst. On a kind of Fatou property of context-free groups. Theoretical Computer Science, 93(2):327– 331, February 17, 1992. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Waldmann:1992:SOS

[438] U. Waldmann. Semantics of ordersorted specifications. Theoretical Computer Science, 94(1):1–35, March 02, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lamarche:1992:QDI

[439] F. Lamarche. Quantitative domains and infinitary algebras. Theoretical Computer Science, 94(1):37–62, March 02, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Filippenko:1992:DLP

[440] I. Filippenko and F. L. Morris. Domains for logic programming. Theoretical Computer Science, 94(1):63– 69, March 02, 1992. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Manca:1992:SCB

[441] V. Manca and A. Salibra. Soundness and completeness of the Birkhoff equational calculus for many-sorted algebras with possibly empty carrier sets. Theoretical Computer Science, 94(1): 101–124, March 02, 1992. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Caspi:1992:CDL

[442] P. Caspi. Clocks in dataflow languages. Theoretical Computer Science, 94(1): 125–140, March 02, 1992. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Koutney:1992:APT

[443] M. Koutney. Adequacy-preserving transformations of COSY path programs. Theoretical Computer Science, 94(1):141–158, March 02, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Mauduit:1992:I

[444] C. Mauduit. Introduction. Theoretical Computer Science, 94(2):159– ??, March 09, 1992. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Allouche:1992:PSS

[445] J.-P. Allouche, P. Morton, and J. Shallit. Pattern spectra, substring enumeration, and automatic sequences. Theoretical Computer Science, 94(2):161– 174, March 09, 1992. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Baratchart:1992:RAP

[446] L. Baratchart, M. Olivi, and F. Wielonsky. On a rational approximation problem in the real Hardy space H_{2}. Theoretical Computer Science, 94(2):175– 197, March 09, 1992. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dehornoy:1992:PMG

[447] P. Dehornoy. Problème de mots dans les gerbes libres. (French) [Word problems in free groups]. Theoretical Computer Science, 94(2):199–213, March 09, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ferenczi:1992:TMS

[448] S. Ferenczi. Tiling the Morse sequence. Theoretical Computer Science, 94(2): 215–221, March 09, 1992. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Frougny:1992:SNL

[449] Ch. Frougny. Systèmes de numération linéaires et theta-représentations. (French) [Linear numeration systems and theta-representation]. Theoretical Computer Science, 94(2):223–236, March 09, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Galmiche:1992:PDC

[450] D. Galmiche. Program development in constructive type theory. Theoretical Computer Science, 94(2):237– 259, March 09, 1992. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gardy:1992:MCL

[451] D. Gardy. Méthode de col et lois limités en analyse combinatoire. (French) [Bottleneck method and limit laws in combinational analysis]. Theoretical Computer Science, 94(2):261–280, March 09, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gil:1992:CEG

[452] B. Gil. Complete extension of general logic programs. Theoretical Computer Science, 94(2):281–294, March 09, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lachaud:1992:ASC

[453] G. Lachaud. Artin–Schreier curves, exponential sums, and coding theory. Theoretical Computer Science, 94(2): 295–310, March 09, 1992. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Mery:1992:NSD

[454] D. Mery. The NU system as a development system for concurrent programs: delta NU. Theoretical Computer Science, 94(2):311–334, March 09, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Perrin:1992:RPP

[455] D. Perrin and M. Parigot. Recursive programming with proofs. Theoretical Computer Science, 94(2):335– 356, March 09, 1992. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Perrin:1992:PM

[456] D. Perrin. On positive matrices. Theoretical Computer Science, 94(2):357– ??, March 09, 1992. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Restivo:1992:NRS

[457] A. Restivo. A note on renewal systems. Theoretical Computer Science, 94(2):

367–371, March 09, 1992. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Wen:1992:SST

[458] Zhi-Xiong Wen and Zhi-Ying Wen. Some studies on the (p, q)-type sequences. Theoretical Computer Science, 94(2):373–393, March 09, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Vianu:1992:CLC

[459] V. Vianu and G. Vossen. Conceptual level concurrency control of relational update transactions. Theoretical Computer Science, 95(1):1–42, March 23, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Giordano:1992:EHC

[460] L. Giordano, A. Martelli, and G. Rossi. Extending Horn clause logic with implication goals. Theoretical Computer Science, 95(1):43–74, March 23, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Sain:1992:TLN

[461] I. Sain. Temporal logics need their clocks. Theoretical Computer Science, 95(1):75–95, March 23, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Arikawa:1992:LEF

[462] S. Arikawa, T. Shinohara, and A. Yamamoto. Learning elementary formal systems. Theoretical Computer Science, 95(1):97–113, March 23, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bellin:1992:DPR

[463] G. Bellin and J. Ketonen. A decision procedure revisited: Notes on direct logic, linear logic and its implementation. Theoretical Computer Science, 95(1):115–142, March 23, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Jacobs:1992:FMP

[464] B. Jacobs, I. Margaria, and M. Zacchi. Filter models with polymorphic types. Theoretical Computer Science, 95(1): 143–158, March 23, 1992. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Schwer:1992:FCV

[465] S. R. Schwer. Fine covers of a VAS language. Theoretical Computer Science, 95(1):159–168, March 23, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Beauquier:1992:TDP

[466] J. Beauquier. Two distributed problems involving Byzantine processes. Theoretical Computer Science, 95(1): 169–185, March 23, 1992. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ollagnier:1992:PDC

[467] J. Moulin Ollagnier. Proof of Dejean’s conjecture for alphabets with 5, 6, 7, 8, 9, 10 and 11 letters. Theoretical Computer Science, 95(2):187– 205, March 30, 1992. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Quinton:1992:SCA

[468] P. Quinton and Y. Robert. Systolic convolution of arithmetic functions. Theoretical Computer Science, 95(2): 207–229, March 30, 1992. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Rabani:1992:SCS

[469] Y. Rabani and Z. Galil. On the space complexity of some algorithms for sequence comparison. Theoretical Computer Science, 95(2):231–244, March 30, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ausiello:1992:LCM

[470] G. Ausiello, G. F. Italiano, A. Marchetti Spaccamela, and U. Nanni. Online computation of minimal and maximal length paths. Theoretical Computer Science, 95(2):245–261, March 30, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Zhang:1992:PTA

[471] S. Zhang. Polynomial-time algorithms for testing strong isomorphism and computing the automorphism group of R-strongly connected automata. Theoretical Computer Science, 95(2):263– 277, March 30, 1992. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Pierre:1992:RIG

[472] L. Pierre. Rational indexes of generators of the cone of context-free languages. Theoretical Computer Science, 95(2):279–305, March 30, 1992.

CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Spencer:1992:USG

[473] J. Spencer. Ulam’s searching game with a fixed number of lies. Theoretical Computer Science, 95(2):307– 321, March 30, 1992. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Akl:1992:DMI

[474] S. G. Akl, M. Cosnard, and A. G. Ferreira. Data-movement-intensive problems: two folk theorems in parallel computation revisited. Theoretical Computer Science, 95(2):323– 337, March 30, 1992. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Shankar:1992:GBR

[475] P. Shankar and B. S. Adiga. A graph-based regularity test for deterministic context-free languages. Theoretical Computer Science, 95(2):339– 340, March 30, 1992. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:1992:WCC

[476] Anonymous. 2nd Workshop on Concurrency and Compositionality. Theoretical Computer Science, 96(1):??, April 06, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

DeNicola:1992:P

[477] R. De Nicola and U. Montanari. Preface. Theoretical Computer Science, 96 (1):1–??, April 06, 1992. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Nielsen:1992:ETS

[478] M. Nielsen, G. Rozenberg, and P. S. Thiagarajan. Elementary transition systems. Theoretical Computer Science, 96(1):3–33, April 06, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Mukund:1992:LCW

[479] M. Mukund and P. S. Thiagarajan. A logical characterization of well branching event structures. Theoretical Computer Science, 96(1):35–72, April 06, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Meseguer:1992:CRL

[480] J. Meseguer. Conditional rewriting logic as a unified model of concurrency. Theoretical Computer Science, 96(1): 73–155, April 06, 1992. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bradfield:1992:LMC

[481] J. Bradfield and C. Stirling. Local model checking for infinite state spaces. Theoretical Computer Science, 96(1): 157–174, April 06, 1992. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Best:1992:PNS

[482] E. Best and M. Koutny. Petri net semantics of priority systems. Theoretical Computer Science, 96(1):175– 215, April 06, 1992. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Berry:1992:CAM

[483] G. Berry and G. Boudol. The chemical abstract machine. Theoretical Computer Science, 96(1):217–248, April 06, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Astesiano:1992:OST

[484] E. Astesiano, A. Giovini, and G. Reggio. Observational structures and their logics. Theoretical Computer Science, 96(1):249–283, April 06, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Garg:1992:CRE

[485] V. K. Garg and M. T. Ragunath. Concurrent regular expressions and their relationship to Petri nets. Theoretical Computer Science, 96(2):285– 304, April 13, 1992. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Huynh:1992:NCR

[486] D. T. Huynh. Nonuniform complexity and the randomness of certain complete languages. Theoretical Computer Science, 96(2):305–324, April 13, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ito:1992:LWE

[487] M. Ito, H. Jurgensen, H. J. Shyr, and G. Thierrin. Languages whose n-element subsets are codes. Theoretical Computer Science, 96(2):325– 344, April 13, 1992. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Barua:1992:HKH

[488] R. Barua. The Hausdorff-Kuratowski hierarchy of omega-regular languages and a hierarchy of Muller automata. Theoretical Computer Science, 96(2): 345–360, April 13, 1992. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Plambeck:1992:DKS

[489] T. E. Plambeck. Daisies, Kayles, and the Sibert-Conway decomposition in misere octal games. Theoretical Computer Science, 96(2):361–388, April 13, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ferguson:1992:MBK

[490] T. S. Ferguson. Mate with bishop and knight in Kriegspiel (mathematical games). Theoretical Computer Science, 96(2):389–403, April 13, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Duchamp:1992:PCS

[491] G. Duchamp and D. Krob. On the partially commutative shuffle product. Theoretical Computer Science, 96(2): 405–410, April 13, 1992. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Slobodova:1992:SPS

[492] A. Slobodova. Some properties of space-bounded synchronized alternating Turing machines with universal states only. Theoretical Computer Science, 96(2):411–419, April 13, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Girard:1992:BLL

[493] J. Y. Girard, A. Scedrov, and P. J. Scott. Bounded linear logic: a modular approach to polynomial-time computability. Theoretical Computer Science, 97(1):1–66, April 20, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

German:1992:SRF

[494] S. M. German. Semantics and reasoning with free procedures. Theoretical Computer Science, 97(1):67– 81, April 20, 1992. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Perraud:1992:OSK

[495] J. Perraud, O. Roux, and M. Huou. Operational semantics of a kernel of the language ELECTRE. Theoretical Computer Science, 97(1):83–103, April 20, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Monfroglio:1992:IPL

[496] A. Monfroglio. Integer programs for logic constraint satisfaction. Theoretical Computer Science, 97(1):105– 130, April 20, 1992. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Darondeau:1992:FDD

[497] P. Darondeau, D. Nolte, L. Priese, and S. Yoccoz. Fairness, distances and degrees. Theoretical Computer Science, 97(1):131–142, April 20, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Baratella:1992:CRA

[498] S. Baratella. A completeness result for allowed semi-strict programs with respect to well-behaved and allowed query clauses. Theoretical Computer Science, 97(1):143–156, April 20, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Weibel:1992:ECL

[499] T. Weibel. Extension of combinatory logic to a theory of combinatory representation. Theoretical Computer Science, 97(1):157–173, April 20, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Doets:1992:SST

[500] K. Doets. A slight strengthening of a theorem of Blair and Kunen. Theoretical Computer Science, 97(1):175– 181, April 20, 1992. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Krentel:1992:GOP

[501] M. W. Krentel. Generalizations of opt P to the polynomial hierarchy. Theoretical Computer Science, 97(2):183– 198, April 27, 1992. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Watanabe:1992:PTT

[502] O. Watanabe and Shouwen Tang. On polynomial-time Turing and many-one completeness in PSPACE. Theoretical Computer Science, 97(2):199– 215, April 27, 1992. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Yoo:1992:ERE

[503] H. Yoo and K. Hashiguchi. Extended regular expressions of arbitrary star degrees. Theoretical Computer Science, 97(2):217–231, April 27, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Muller:1992:AAW

[504] D. E. Muller, A. Saoudi, and P. E. Schupp. Alternating automata, the weak monadic theory of trees and its complexity. Theoretical Computer Science, 97(2):233–244, April 27, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Karner:1992:NTP

[505] G. Karner. Nivat’s theorem for pushdown transducers. Theoretical Computer Science, 97(2):245–262, April 27, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Shore:1992:DRS

[506] R. A. Shore and T. A. Slaman. The p-T -degrees of the recursive sets: lattice embeddings, extensions of embeddings and the two-quantifier theory. Theoretical Computer Science, 97(2):263– 284, April 27, 1992. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hromkovic:1992:LBA

[507] J. Hromkovic, S. A. Lozkin, A. I. Rybko, A. A. Sapozenko, and N. A. Skalikova. Lower bounds on the area complexity of Boolean circuits. Theoretical Computer Science, 97(2):285– 300, April 27, 1992. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Guaiana:1992:SFT

[508] G. Guaiana, A. Restivo, and S. Salemi. Star-free trace languages. Theoretical Computer Science, 97(2):301–311, April 27, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Duchamp:1992:SWA

[509] G. Duchamp, G. Jacob, and D. Krob. Second Workshop on Algebraic and Computer-Theoretic Aspects of Formal Power Series. Theoretical Computer Science, 98(1):??, May 11, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Choffrut:1992:RRR

[510] C. Choffrut. Rational relations and rational series. Theoretical Computer Science, 98(1):5–13, May 11, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Karhumaki:1992:MDV

[511] J. Karhumaki. Multiplicities: A deterministic view of nondeterminism. Theoretical Computer Science, 98(1):15– 25, May 11, 1992. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Karner:1992:TFP

[512] G. Karner. On transductions of formal power series over complete semirings. Theoretical Computer Science, 98 (1):27–39, May 11, 1992. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Varricchio:1992:RSC

[513] S. Varricchio. Rational series with coefficients in a commutative ring. Theoretical Computer Science, 98(1):41– 50, May 11, 1992. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bjorner:1992:RMF

[514] A. Bjorner and C. Reutenauer. Rationality of the Möbius function of subword order. Theoretical Computer Science, 98(1):53–63, May 11, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Delest:1992:AGU

[515] M. P. Delest and J. M. Fedou. Attribute grammars are useful for combinatorics. Theoretical Computer Science, 98(1):65–76, May 11, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cauchon:1992:SMN

[516] G. Cauchon. Séries de Malcev-Neumann sur le groupe libre et questions de rationalité. (French) [Malcev-Neumann series on the free group and questions on nationality]. Theoretical Computer Science, 98(1):79– 97, May 11, 1992. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dumas:1992:SPS

[517] F. Dumas. Skew power series rings with general commutation formula. Theoretical Computer Science, 98(1):99– 114, May 11, 1992. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Mourrain:1992:CIA

[518] B. Mourrain. Computable identities in the algebra of formal matrices. Theoretical Computer Science, 98(1):115– 133, May 11, 1992. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Diop:1992:DAD

[519] S. Diop. Differential-algebraic decision methods and some applications to system theory. Theoretical Computer Science, 98(1):137–161, May 11, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Allouche:1992:RRS

[520] J.-P. Allouche and J. Shallit. The ring of k-regular sequences. Theoretical Computer Science, 98(2):163– 197, May 18, 1992. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Schwer:1992:CFL

[521] S. R. Schwer. The context-freeness of the languages associated with vector addition systems is decidable. Theoretical Computer Science, 98(2):199– 247, May 18, 1992. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Santean:1992:INC

[522] L. Santean and J. Kari. The impact of the number of cooperating grammars on the generative power. Theoretical Computer Science, 98(2):249– 262, May 18, 1992. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Yen:1992:MAD

[523] Hsu-Chun Yen. A multiparameter analysis of domino tiling with an application to concurrent systems. Theoretical Computer Science, 98(2):263– 287, May 18, 1992. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Palis:1992:UBR

[524] M. A. Palis and S. M. Shende. Upper bounds on recognition of a hierarchy of non-context-free languages. Theoretical Computer Science, 98(2):289– 319, May 18, 1992. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Van:1992:VFM

[525] Do Long Van and Phan Trung Huy. Varieties of finite monoids and Buchi-McNaughton theorem. Theoretical Computer Science, 98(2):321–337, May 18, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chrobak:1992:HCT

[526] M. Chrobak and L. L. Larmore. HARMONIC is 3-competitive for two servers. Theoretical Computer Science, 98(2):339–346, May 18, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Amtoft:1992:PMO

[527] T. Amtoft and J. Larsson Traff. Partial memoization for obtaining linear time behavior of a 2DPDA. Theoretical Computer Science, 98(2):347– 356, May 18, 1992. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cazanescu:1992:GRA

[528] V.-E. Cazanescu and Gh. Stefanescu. A general result on abstract flowchart schemes with applications to the study of accessibility, reduction and minimization. Theoretical Computer Science, 99(1):1–63, June 01, 1992. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hein:1992:CPL

[529] J. Hein. Completions of perpetual logic programs. Theoretical Computer Science, 99(1):65–78, June 01, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lambert:1992:SDR

[530] J. L. Lambert. A structure to decide reachability in Petri nets. Theoretical Computer Science, 99(1):79– 104, June 01, 1992. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hesselink:1992:PFU

[531] W. H. Hesselink. Processes and formalisms for unbounded choice. Theoretical Computer Science, 99(1):105– 119, June 01, 1992. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Priese:1992:SFU

[532] L. Priese and D. Nolte. Strong fairness and ultra metrics. Theoretical Computer Science, 99(1):121–140, June 01, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Mulry:1992:MAS

[533] P. S. Mulry. Monads and algebras in the semantics of partial data types. Theoretical Computer Science, 99(1): 141–155, June 01, 1992. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Rajasekaran:1992:NAP

[534] S. Rajasekaran and J. H. Reif. Nested annealing: a provable improvement to simulated annealing. Theoretical Computer Science, 99(1):157–176, June 01, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bozapalidis:1992:ATR

[535] S. Bozapalidis. Alphabetic tree relations. Theoretical Computer Science, 99(2):177–211, June 08, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Nummenmaa:1992:CCR

[536] J. Nummenmaa. Constructing compact rectilinear planar layouts using canonical representation of planar graphs. Theoretical Computer Science, 99(2): 213–230, June 08, 1992. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Neraud:1992:RSF

[537] J. Neraud. On the rank of the subsets of a free monoid. Theoretical Computer Science, 99(2):231–241, June 08, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ibarra:1992:SBS

[538] O. H. Ibarra and N. Q. Tran. On space-bounded synchronized alternating Turing machines. Theoretical Computer Science, 99(2):243–264, June 08, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Zhang:1992:EST

[539] S. Zhang. Efficient simplicity testing of automata. Theoretical Computer Science, 99(2):265–??, June 08, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Balcazar:1992:LAC

[540] J. L. Balcazar and U. Schoning. Logarithmic advice classes. Theoretical Computer Science, 99(2):279–290, June 08, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Varricchio:1992:DEP

[541] S. Varricchio. On the decidability of the equivalence problem for partially commutative rational power series. Theoretical Computer Science, 99(2):291– 299, June 08, 1992. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ibarra:1992:CET

[542] O. H. Ibarra, Tao Jiang, and Hui Wang. A characterization of exponential-time languages by alternating context-free grammars. Theoretical Computer Science, 99(2):301–315, June 08, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fouks:1992:TFR

[543] J. D. Fouks. Tseitin’s formulas revisited. Theoretical Computer Science, 99 (2):315–??, June 08, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Mosse:1992:PMR

[544] B. Mosse. Puissances de mots et reconnaissabilité des points fixes d’une substitution. (French) [Word power and recognizability of substitutional fixed points]. Theoretical Computer Science, 99(2):327–??, June 08, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hoang:1992:PAM

[545] C. T. Hoang. A parallel algorithm for minimum weighted colouring of triangulated graphs. Theoretical Computer Science, 99(2):335–344, June 08, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Trahan:1992:MDS

[546] J. L. Trahan, M. C. Loui, and V. Ramachandran. Multiplication, division, and shift instructions in parallel random access machines. Theoretical Computer Science, 100(1):1–44, June 22, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Goerdt:1992:CCC

[547] A. Goerdt. Characterizing complexity classes by higher type primitive recursive definitions. Theoretical Computer Science, 100(1):45–66, June 22, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

DeLuca:1992:NRC

[548] A. De Luca and S. Varricchio. On noncounting regular classes. Theoretical Computer Science, 100(1):67– 104, June 22, 1992. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Holte:1992:PST

[549] R. Holte, L. Rosier, I. Tulchinsky, and D. Varvel. Pinwheel scheduling with two distinct numbers. Theoretical Computer Science, 100(1):105– 135, June 22, 1992. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Benninger:1992:ESD

[550] D. Benninger and J. Schmid. Effective subdirect decomposition: a case study. Theoretical Computer Science, 100(1): 137–156, June 22, 1992. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Board:1992:NOA

[551] R. Board and L. Pitt. On the necessity of Occam algorithms. Theoretical Computer Science, 100(1):157–184, June 22, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Santoro:1992:DSA

[552] N. Santoro, J. B. Sidney, and S. J. Sidney. A distributed selection algorithm and its expected communication complexity. Theoretical Computer Science, 100(1):185–204, June 22, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Toda:1992:PTT

[553] S. Toda and O. Watanabe. Polynomial time 1-Turing reductions from #ph to #p. Theoretical Computer Science, 100(1):205–221, June 22, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Erdos:1992:DLN

[554] P. Erdos and D. F. Hsu. Distributed loop network with minimum transmission delay. Theoretical Computer Science, 100(1):223–241, June 22, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Prodinger:1992:HAA

[555] H. Prodinger. Hypothetical analyses: approximate counting in the style of Knuth, path length in the style of Flajolet. Theoretical Computer Science, 100(1):243–251, June 22, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Beame:1992:CCS

[556] P. Beame, E. Brisson, and R. Ladner. The complexity of computing symmetric functions using threshold circuits. Theoretical Computer Science, 100(1): 253–265, June 22, 1992. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Moss:1992:FAC

[557] L. S. Moss, J. Meseguer, and J. A. Goguen. Final algebras, cosemicomputable algebras and degrees of unsolvability. Theoretical Computer Science, 100(2):267–302, June 29, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dezani-Ciancaglini:1992:ITC

[558] M. Dezani-Ciancaglini and J. R. Hindley. Intersection types for combinatory logic. Theoretical Computer Science, 100(2):303–324, June 29, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bartha:1992:FTS

[559] M. Bartha. Foundations of a theory of synchronous systems. Theoretical Computer Science, 100(2):325–346, June 29, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Wang:1992:CBD

[560] Ke Wang. On characterizing boundedness of database schemes with bounded dependencies. Theoretical Computer Science, 100(2):347–364, June 29, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Back:1992:CAD

[561] R. J. R. Back and J. Von Wright. Combining angels, demons and miracles in program specifications. Theoretical Computer Science, 100(2):365– 383, June 29, 1992. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Meinke:1992:UAH

[562] K. Meinke. Universal algebra in higher types. Theoretical Computer Science, 100(2):385–417, June 29, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Grandjean:1992:E

[563] E. Grandjean. Editorial. Theoretical Computer Science, 101(1):1– ??, July 13, 1992. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Courcelle:1992:MSO

[564] B. Courcelle. The monadic secondorder logic of graphs VII: Graphs as relational structures. Theoretical Computer Science, 101(1):3–33, July 13, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gradel:1992:CCC

[565] E. Gradel. Capturing complexity classes by fragments of second-order logic. Theoretical Computer Science, 101(1):35–57, July 13, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Mazoyer:1992:LST

[566] J. Mazoyer and N. Reimen. A linear speed-up theorem for cellular automata. Theoretical Computer Science, 101(1):59–98, July 13, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Michel:1992:SSC

[567] P. Michel. A survey of space complexity. Theoretical Computer Science, 101 (1):99–132, July 13, 1992. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Peladeau:1992:FRL

[568] P. Peladeau. Formulas, regular languages and Boolean circuits. Theoretical Computer Science, 101(1):133– 141, July 13, 1992. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

deRougemont:1992:FDI

[569] M. de Rougemont. The functional dimension of inductive definitions. Theoretical Computer Science, 101(1):143– 158, July 13, 1992. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kwiatkowska:1992:E

[570] M. Z. Kwiatkowska. Editorial. Theoretical Computer Science, 101(2):159– ??, July 20, 1992. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Best:1992:TCH

[571] Eike Best, Jörg Desel, and Javier Esparza. Traps characterize home states in free choice systems. Theoretical Computer Science, 101(2):161–176, July 20, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1992&volume=101&issue=2&aid=1130.

Brookes:1992:TTP

[572] Stephen Brookes and Shai Geva. Towards a theory of parallel algorithms on concrete data structures. Theoretical Computer Science, 101(2):177–221, July 20, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1992&volume=101&issue=2&aid=1131.

Bloom:1992:EPE

[573] Bard Bloom and Albert R. Meyer. Experimenting with process equivalence. Theoretical Computer Science, 101(2):223–237, July 20, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.
cgi?year=1992&volume=101&issue=
2&aid=1132.

deBoer:1992:FSC

[574] F. S. de Boer, J. N. Kok, C. Palamidessi, and J. J. M. M. Rutten. From failure to success: comparing a denotational and a declarative semantics for Horn clause logic. Theoretical Computer Science, 101(2):239–263, July 20, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1992&volume=101&issue=2&aid=1133.

Gunawardena:1992:CA

[575] Jeremy Gunawardena. Causal automata. Theoretical Computer Science, 101(2):265–288, July 20, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1992&volume=101&issue=2&aid=1134.

Hooman:1992:CAS

[576] J. J. M. Hooman, S. Ramesh, and W. P. de Roever. A compositional axiomatization of Statecharts. Theoretical Computer Science, 101(2):289–335, July 20, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1992&volume=101&issue=2&aid=1135.

Katz:1992:DCI

[577] Shmuel Katz and Doron Peled. Defining conditional independence using collapses. Theoretical Computer Science, 101(2):337–359, July 20, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1992&volume=101&issue=2&aid=1136.

Grue:1992:MT

[578] Klaus Grue. Map theory. Theoretical Computer Science, 102(1):1–134, August 03, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1992&volume=102&issue=1&aid=1169.

VanBakel:1992:CRI

[579] S. Van Bakel. Complete restrictions of the intersection type discipline. Theoretical Computer Science, 102(1):135– 163, August 03, 1992. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Devillers:1992:MPB

[580] R. Devillers. Maximality preserving bisimulation. Theoretical Computer Science, 102(1):165–183, August 03, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Esparza:1992:PTA

[581] J. Esparza and M. Silva. A polynomialtime algorithm to decide liveness of bounded free choice nets. Theoretical Computer Science, 102(1):185– 205, August 03, 1992. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dosen:1992:NCL

[582] K. Dosen. Nonmodal classical linear predicate logic is a fragment of intuitionistic linear logic. Theoretical Computer Science, 102(1):207– 214, August 03, 1992. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bruschi:1992:SSP

[583] D. Bruschi. Strong separations of the polynomial hierarchy with oracles: constructive separations by immune and simple sets. Theoretical Computer Science, 102(2):215–252, August 10, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Skubiszewski:1992:BPS

[584] M. Skubiszewski. Binary periodic synchronizing sequences. Theoretical Computer Science, 102(2):253–281, August 10, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dagum:1992:APG

[585] P. Dagum and M. Luby. Approximating the permanent of graphs with large factors. Theoretical Computer Science, 102(2):283–305, August 10, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Mirwald:1992:MCQ

[586] R. Mirwald and C. P. Schnorr. The multiplicative complexity of quadratic Boolean forms. Theoretical Computer Science, 102(2):307–328, August 10, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Halpenny:1992:CMS

[587] Liam Halpenny and Christopher J. Smyth. A classification of minimal standard-path 2 × 2 switching networks. Theoretical Computer Science, 102(2):329–354, August 10, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1992&volume=102&issue=2&aid=1227.

Prasad:1992:AEN

[588] L. Prasad and S. S. Iyengar. An asymptotic equality for the number of necklaces in a shuffle-exchange network. Theoretical Computer Science, 102(2): 355–365, August 10, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Choffrut:1992:E

[589] C. Choffrut and Th Lengaur. Editorial. Theoretical Computer Science, 103(1): 1–??, August 24, 1992. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Beigel:1992:CCT

[590] R. Beigel and J. Gill. Counting classes: thresholds, parity, mods, and fewness. Theoretical Computer Science, 103(1): 3–23, August 24, 1992. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cai:1992:GII

[591] Jin-Yi Cai, A. Condon, and R. J. Lipton. On games of incomplete information. Theoretical Computer Science, 103(1):25–38, August 24, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Clerbout:1992:SAL

[592] M. Clerbout and Y. Roos. Semicommutations and algebraic languages. Theoretical Computer Science, 103(1):39– 49, August 24, 1992. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Corradini:1992:ASS

[593] A. Corradini and U. Montanari. An algebraic semantics for structured transition systems and its applications to logic programs. Theoretical Computer Science, 103(1):51–106, August 24, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Das:1992:MVH

[594] G. Das and D. Joseph. Minimum vertex hulls for polyhedral domains. Theoretical Computer Science, 103(1):107– 135, August 24, 1992. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lambert:1992:SSC

[595] J. L. Lambert. Sorting the sums (x_{i} + y_{j}) in O(n^{2}) comparisons. Theoretical Computer Science, 103(1):137– 141, August 24, 1992. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Thomas:1992:ITA

[596] W. Thomas. Infinite trees and automaton-definable relations over omega-words. Theoretical Computer Science, 103(1):143–159, August 24, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bauderon:1992:IHI

[597] Michel Bauderon. Infinite hypergraphs II. Systems of recursive equations. Theoretical Computer Science, 103(2):165–190, September 14, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1992&volume=103&issue=2&aid=1055.

Hamaguchi:1992:RTL

[598] Kiyoharu Hamaguchi, Hiromi Hiraishi, and Shuzo Yajima. ∞-regular temporal logic and its model checking problem. Theoretical Computer Science, 103(2):191–204, September 14, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1992&volume=103&issue=2&aid=1148.

Palopoli:1992:TLP

[599] L. Palopoli. Testing logic programs for local stratification. Theoretical Computer Science, 103(2):205–234, September 14, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Felleisen:1992:RRS

[600] M. Felleisen and R. Hieb. The revised report on the syntactic theories of sequential control and state. Theoretical Computer Science, 103(2):235– 271, September 14, 1992. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kurihara:1992:MST

[601] M. Kurihara and A. Ohuchi. Modularity of simple termination of term rewriting systems with shared constructors. Theoretical Computer Science, 103(2):273–282, September 14, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kim:1992:NSM

[602] Byeong Man Kim and Jung Wan Cho. A new subsumption method in the connection graph proof procedure. Theoretical Computer Science, 103(2):283– 309, September 14, 1992. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gunter:1992:MP

[603] C. A. Gunter. The mixed powerdomain. Theoretical Computer Science, 103(2):311–334, September 14, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Maggiolo-Schettini:1992:TAT

[604] A. Maggiolo-Schettini and J. Winkowski. Towards an algebra for timed behaviours. Theoretical Computer Science, 103(2):335–363, September 14, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Marek:1992:RBS

[605] W. Marek and V. S. Subrahmanian. The relationship between stable, supported, default and autoepistemic semantics for general logic programs. Theoretical Computer Science, 103(2): 365–386, September 14, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Mairson:1992:SPT

[606] Harry G. Mairson. A simple proof of a theorem of Statman. Theoretical Computer Science, 103(2):387–394, September 14, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1992&volume=103&issue=2&aid=1244.

Streicher:1992:IIP

[607] Thomas Streicher. Independence of the induction principle and the axiom of choice in the pure calculus of constructions. Theoretical Computer Science, 103(2):395–408, September 14, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1992&volume=103&issue=2&aid=1216.

Dauchet:1992:STM

[608] Max Dauchet. Simulation of Turing machines by a regular rewrite rule. Theoretical Computer Science, 103(2):409–420, September 14, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1992&volume=103&issue=2&aid=1243.

Miola:1992:F

[609] A. Miola. Foreword. Theoretical Computer Science, 104(1):1–??, October 05, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bol:1992:GCR

[610] Roland N. Bol. Generalizing completeness results for loop checks in logic programming. Theoretical Computer Science, 104(1):3–28, October 05, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1992&volume=104&issue=1&aid=1164.

Jouannaud:1992:TCM

[611] Jean-Pierre Jouannaud and Claude Marché. Termination and completion modulo associativity, commutativity and identity. Theoretical Computer Science, 104(1):29–51, October 05, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1992&volume=104&issue=1&aid=1166.

Hennicker:1992:SAA

[612] Rolf Hennicker. A semi-algorithm for algebraic implementation proofs. Theoretical Computer Science, 104(1): 53–87, October 05, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.
cgi?year=1992&volume=104&issue=
1&aid=1168.

Limongelli:1992:ASS

[613] C. Limongelli and M. Temperini. Abstract specification of structures and methods in symbolic mathematical computation. Theoretical Computer Science, 104(1):89–107, October 05, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1992&volume=104&issue=1&aid=1165.

Stickel:1992:PTT

[614] Mark E. Stickel. A Prolog technology theorem prover: a new exposition and implementation in Prolog. Theoretical Computer Science, 104(1):109–128, October 05, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1992&volume=104&issue=1&aid=1167.

Talcott:1992:TPD

[615] Carolyn Talcott. A theory for program and data type specification. Theoretical Computer Science, 104(1):129–159, October 05, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1992&volume=104&issue=1&aid=1163.

Straubing:1992:CCD

[616] H. Straubing and P. Weil. On a conjecture concerning dot-depth two languages. Theoretical Computer Science, 104(2):161–183, October 12, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kim:1992:RBP

[617] Changwook Kim and I. H. Sudborough. On reversal-bounded picture languages. Theoretical Computer Science, 104(2):185–206, October 12, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Takada:1992:LSS

[618] Y. Takada. Learning semilinear sets from examples and via queries. Theoretical Computer Science, 104(2):207– 233, October 12, 1992. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Weir:1992:GHB

[619] D. J. Weir. A geometric hierarchy beyond context-free languages. Theoretical Computer Science, 104(2):235– 261, October 12, 1992. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bovet:1992:UAD

[620] D. P. Bovet, P. Crescenzi, and R. Silvestri. A uniform approach to define complexity classes. Theoretical Computer Science, 104(2):263–283, October 12, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Jansen:1992:POF

[621] K. Jansen. Processor optimization for flow graphs. Theoretical Computer Science, 104(2):285–298, October 12, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Leermakers:1992:RAP

[622] René Leermakers. Recursive ascent parsing: from Earley to Marcus. Theoretical Computer Science, 104(2):299–312, October 12, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1992&volume=104&issue=2&aid=1194.

Leermakers:1992:FLP

[623] R. Leermakers, L. Augusteijn, and F. E. J. Kruseman Aretz. A functional LR parser. Theoretical Computer Science, 104(2):313–323, October 12, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dung:1992:RBS

[624] Phan Minh Dung. On the relations between stable and well-founded semantics of logic programs. Theoretical Computer Science, 105(1):7–25, October 26, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1992&volume=105&issue=1&aid=1172.

Kanchanasut:1992:TNL

[625] Kanchana Kanchanasut and Peter J. Stuckey. Transforming normal logic programs to constraint logic programs. Theoretical Computer Science, 105(1):27–56, October 26, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1992&volume=105&issue=1&aid=1173.

Sato:1992:EPF

[626] Taisuke Sato. Equivalence-preserving first-order unfold/fold transformation systems. Theoretical Computer Science, 105(1):57–84, October 26, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1992&volume=105&issue=1&aid=1174.

Gabbrielli:1992:UFS

[627] Maurizio Gabbrielli and Giorgio Levi. Unfolding and fixpoint semantics of concurrent constraint logic programs. Theoretical Computer Science, 105(1):85–128, October 26, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1992&volume=105&issue=1&aid=1175.

Hofbauer:1992:TPM

[628] Dieter Hofbauer. Termination proofs by multiset path orderings imply primitive recursive derivation lengths. Theoretical Computer Science, 105(1):129–140, October 26, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1992&volume=105&issue=1&aid=1176.

Debart:1992:MLP

[629] Françoise Debart, Patrice Enjalbert, and Madeleine Lescot. Multimodal logic programming using equational and order-sorted logic. Theoretical Computer Science, 105(1):141– 166, October 26, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.

cgi?year=1992&volume=105&issue=
1&aid=1177.

Mason:1992:IEF

[630] Ian Mason and Carolyn L. Talcott. Inferring the equivalence of functional programs that mutate data. Theoretical Computer Science, 105(2):167–215, November 09, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1992&volume=105&issue=2&aid=1095.

Goguen:1992:OSA

[631] Joseph A. Goguen and José Meseguer. Order sorted algebra I. equational deduction for multiple inheritance, overloading, exceptions and partial operations. Theoretical Computer Science, 105(2):217–273, November 09, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1992&volume=105&issue=2&aid=1268.

Zhang:1992:DTE

[632] Ming-Hua Zhang. Data types with errors and exceptions. Theoretical Computer Science, 105(2):275–299, November 09, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1992&volume=105&issue=2&aid=1267.

Arnold:1992:P

[633] A. Arnold. Preface. Theoretical Computer Science, 106(1):1–??, November 30, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Boudol:1992:GVL

[634] G. Boudol and K. G. Larsen. Graphical versus logical specifications. Theoretical Computer Science, 106(1):3– 20, November 30, 1992. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cai:1992:MEB

[635] J. Cai, R. Paige, and R. Tarjan. More efficient bottom-up multi-pattern matching in trees. Theoretical Computer Science, 106(1):21–60, November 30, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Caucal:1992:RSP

[636] D. Caucal. On the regular structure of prefix rewriting. Theoretical Computer Science, 106(1):61–86, November 30, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kounalis:1992:TGC

[637] E. Kounalis. Testing for the ground (co-)reducibility property in termrewriting systems. Theoretical Computer Science, 106(1):87–117, November 30, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Schwartzbach:1992:IRD

[638] M. I. Schwartzbach. Interpretations of recursively defined types. Theoretical Computer Science, 106(1):119– 134, November 30, 1992. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Seidl:1992:SVT

[639] H. Seidl. Single-valuedness of tree transducers is decidable in polynomial time. Theoretical Computer Science, 106(1):135–181, November 30, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Frougny:1992:CLN

[640] Ch. Frougny. Confluent linear numeration systems. Theoretical Computer Science, 106(2):183–219, December 14, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Michel:1992:CLT

[641] P. Michel. Complexity of logical theories involving coprimality. Theoretical Computer Science, 106(2):221–241, December 14, 1992. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

El-Taha:1992:DAQ

[642] M. El-Taha and S. Stidham, Jr. Deterministic analysis of queueing systems with heterogeneous servers. Theoretical Computer Science, 106(2):243– 264, December 14, 1992. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bern:1992:FGA

[643] M. W. Bern, H. J. Karloff, P. Raghavan, and B. Schieber. Fast geometric approximation techniques and geometric embedding problems. Theoretical Computer Science, 106(2):265– 281, December 14, 1992. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hausser:1992:CLA

[644] R. Hausser. Complexity in leftassociative grammar. Theoretical Computer Science, 106(2):283–308, December 14, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bodlaender:1992:CCG

[645] Hans L. Bodlaender and Dieter Kratsch. The complexity of coloring games on perfect graphs. Theoretical Computer Science, 106(2):309– 326, December 14, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.
cgi?year=1992&volume=106&issue=
2&aid=1332.

Robson:1992:MLG

[646] J. M. Robson. More languages of generalised star height 1. Theoretical Computer Science, 106(2):327–335, December 14, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1992&volume=106&issue=2&aid=1233.

Villemaire:1992:TU

[647] Roger Villemaire. The theory of ⌊N + V_{k}, V_{l}⌋ is undecidable. Theoretical Computer Science, 106(2):337– 349, December 14, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/

tree/store/tcs/cas_sub/browse/browse.
cgi?year=1992&volume=106&issue=
2&aid=1234.

Damm:1992:SCC

[648] C. Damm and Ch. Meinel. Separating complexity classes related to Omegadecision trees. Theoretical Computer Science, 106(2):351–360, December 14, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Huang:1992:SPA

[649] Shou-Hsuan Stephen Huang, Hongfei Liu, and Venkatraman Viswanathan. A sublinear parallel algorithm for some dynamic programming problems. Theoretical Computer Science, 106(2):361–371, December 14, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1992&volume=106&issue=2&aid=1324.

Subramanian:1992:LT

[650] K. G. Subramanian, R. Siromoney, and L. Mathew. Lyndon trees. Theoretical Computer Science, 106(2):373– 383, December 14, 1992. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Razborov:1992:DCD

[651] A. A. Razborov. On the distributional complexity of disjointness. Theoretical Computer Science, 106(2):385–390, December 14, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/

                                        cas_sub/browse/browse.cgi?year=
                                        1992&volume=106&issue=2&aid=1223.

Andersson:1992:CBP

[652] Arne Andersson. Comments on “On the balance property of Patricia tries: External path length viewpoint”. Theoretical Computer Science, 106(2):391–393, December 14, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1992&volume=106&issue=2&aid=1356.

See reply [653].

Kirschenhofer:1992:PMD

[653] Peter Kirschenhofer, Helmut Prodinger, and Wojciech Szpankowski. Probabilistic modeling of data structures on words: a reply to Professor Andersson’s letter. Theoretical Computer Science, 106(2):395–400, December 14, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1992&volume=106&issue=2&aid=1366.

See comments [652].

Selman:1993:F

[654] A. L. Selman. Foreword. Theoretical Computer Science, 107(1):1– ??, January 04, 1993. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Alvarez:1993:VHL

[655] C. Alvarez and B. Jenner. A very hard log-space counting class. Theoretical Computer Science, 107(1):3– 30, January 04, 1993. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bedard:1993:EBM

[656] F. Bedard, F. Lemieux, and P. McKenzie. Extensions to Barrington’s Mprogram model. Theoretical Computer Science, 107(1):31–61, January 04, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Heiman:1993:ROT

[657] R. Heiman, I. Newman, and A. Wigderson. On read-once threshold formulae and their randomized decision tree complexity. Theoretical Computer Science, 107(1):63–76, January 04, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lange:1993:UC

[658] K.-J. Lange. Unambiguity of circuits. Theoretical Computer Science, 107(1): 77–94, January 04, 1993. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lutz:1993:CSR

[659] J. H. Lutz and W. J. Schmidt. Circuit size relative to pseudorandom oracles. Theoretical Computer Science, 107(1): 95–120, January 04, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Mansour:1993:CCU

[660] Y. Mansour, N. Nisan, and P. Tiwari. The computational complexity of universal hashing. Theoretical Computer Science, 107(1):121–133, January 04, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Nisan:1993:ROV

[661] N. Nisan. On read-once vs. multiple access to randomness in logspace. Theoretical Computer Science, 107(1):135– 144, January 04, 1993. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Panconesi:1993:QA

[662] A. Panconesi and D. Ranjan. Quantifiers and approximation. Theoretical Computer Science, 107(1):145–163, January 04, 1993. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Jacobs:1993:CCS

[663] Bart Jacobs. Comprehension categories and the semantics of type dependency. Theoretical Computer Science, 107(2):169–207, January 18, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=107&issue=2&aid=1128.

Bellia:1993:EVF

[664] Marco Bellia and M. Eugenia Occhiuto. C-expressions: a variablefree calculus for equational logic programming. Theoretical Computer Science, 107(2):209–252, January 18, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=107&issue=2&aid=1147.

Hirokawa:1993:PTB

[665] Sachio Hirokawa. Principal types of BCK-lambda-terms. Theoretical Computer Science, 107(2):253–276, January 18, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=107&issue=2&aid=1149.

Cortesi:1993:GPN

[666] Agostino Cortesi and Gilberto Filé. Graph properties for normal logic programs. Theoretical Computer Science, 107(2):277–303, January 18, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=107&issue=2&aid=1150.

Main:1993:SMT

[667] Michael G. Main and David L. Black. Semantic models for total correctness and fairness. Theoretical Computer Science, 107(2):305–332, January 18, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=107&issue=2&aid=1151.

Sreedhar:1993:CSR

[668] Vugranam Sreedhar and Kazem Taghva. Capturing strong reduction in director string calculus. Theoretical Computer Science, 107(2):333–347, January 18, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1993&volume=107&issue=2&aid=1146.

Dowek:1993:UPM

[669] Gilles Dowek. The undecidability of pattern matching in calculi where primitive recursive functions are representable. Theoretical Computer Science, 107(2):349–356, January 18, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1993&volume=107&issue=2&aid=1295.

Milner:1993:UDP

[670] Robin Milner and Faron Moller. Unique decomposition of processes. Theoretical Computer Science, 107(2): 357–363, January 18, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/tcs/cas_sub/browse/browse.
                                        cgi?year=1993&volume=107&issue=
                                        2&aid=1249.

Anonymous:1993:ICW

[671] Anonymous. International Colloquium on Words, Languages and Combinatorics. Theoretical Computer Science, 108(1):??, February 01, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ito:1993:F

[672] M. Ito. Foreword. Theoretical Computer Science, 108(1):1–??, February 01, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Almeida:1993:LCP

[673] Jorge Almeida. Locally commutative power semigroups and counting factors of words. Theoretical Computer Science, 108(1):3–16, February 01, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=108&issue=1&aid=1228.

Crvenkovic:1993:KA

[674] Siniša Crvenković and Rozália Sz. Madarász. On Kleene algebras. Theoretical Computer Science, 108(1):17– 24, February 01, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.
cgi?year=1993&volume=108&issue=
1&aid=1214.

Diekert:1993:MFC

[675] Volker Diekert. Möbius functions and confluent semi-commutations. Theoretical Computer Science, 108(1):25– 43, February 01, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.
cgi?year=1993&volume=108&issue=
1&aid=1213.

Frougny:1993:SRR

[676] Christiane Frougny and Jacques Sakarovitch. Synchronized rational relations of finite and infinite words. Theoretical Computer Science, 108(1):45–82, February 01, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1993&volume=108&issue=1&aid=1251.

Garzon:1993:CA

[677] Max Garzon. Cayley automata. Theoretical Computer Science, 108(1):83– 102, February 01, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/tcs/cas_sub/browse/browse.
                                        cgi?year=1993&volume=108&issue=
                                        1&aid=1212.

Karhumaki:1993:EFS

[678] J. Karhumäki. Equations over finite sets of words and equivalence problems in automata theory. Theoretical Computer Science, 108(1):103–118, February 01, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1993&volume=108&issue=1&aid=1211.

Katsura:1993:GFE

[679] Masashi Katsura and Genjiro Tanaka. Groups of finite elementary codes. Theoretical Computer Science, 108(1):119– 149, February 01, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/tcs/cas_sub/browse/browse.
                                        cgi?year=1993&volume=108&issue=
                                        1&aid=1210.

Kunze:1993:SAS

[680] Michael Kunze. Standard automata and semidirect products of transformation semigroups. Theoretical Computer Science, 108(1):151–171, February 01, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=108&issue=1&aid=1209.

Zhang:1993:DRS

[681] Liang Zhang and Weide D. Qiu. Decompositions of recognizable strong maximal codes. Theoretical Computer Science, 108(1):173–183, February 01, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=108&issue=1&aid=1208.

Fulop:1993:TTE

[682] Z. Fulop, F. Herrmann, S. Vagvolgyi, and H. Vogler. Tree transducers with external functions. Theoretical Computer Science, 108(2):185– 236, February 15, 1993. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Van:1993:SZS

[683] Do Long Van, B. Le Saec, and I. Litovsky. Stability for the zigzag submonoids. Theoretical Computer Science, 108(2):237–249, February 15, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Madonia:1993:GSP

[684] M. Madonia, S. Salemi, and T. Sportelli. A generalization of Sardinas and Patterson’s algorithm to z-codes. Theoretical Computer Science, 108(2):251– 270, February 15, 1993. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dietzfelbinger:1993:CMT

[685] M. Dietzfelbinger and W. Maass. The complexity of matrix transposition on one-tape off-line Turing machines with output tape. Theoretical Computer Science, 108(2):271–290, February 15, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Schmid:1993:ACL

[686] U. Schmid. The average CRI-length of a controlled ALOHA collision resolution algorithm. Theoretical Computer Science, 108(2):291–310, February 15, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

BarYehuda:1993:RTG

[687] Reuven Bar Yehuda, Tuvi Etzion, and Shlomo Moran. Rotatingtable games and derivatives of words. Theoretical Computer Science, 108(2):311–329, February 15, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1993&volume=108&issue=2&aid=1318.

Apostolico:1993:ECP

[688] Alberto Apostolico. Efficient CRCW-PRAM algorithms for universal substring searching. Theoretical Computer Science, 108(2):331–344, February 15, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/

cas_sub/browse/browse.cgi?year=
1993&volume=108&issue=2&aid=1277.

Grossi:1993:FCS

[689] Roberto Grossi. On finding common subtrees. Theoretical Computer Science, 108(2):345–356, February 15, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=108&issue=2&aid=1230.

Slowinski:1993:PWI

[690] Karine Slowinski. Picture words with invisible lines. Theoretical Computer Science, 108(2):357–363, February 15, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=108&issue=2&aid=1271.

Middendorf:1993:SCN

[691] M. Middendorf. The shortest common nonsubsequence problem is NPcomplete. Theoretical Computer Science, 108(2):365–369, February 15, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=108&issue=2&aid=1232.

dAmore:1993:WLU

[692] Fabrizio d’Amore, Alberto Marchetti-Spaccamela, and Umberto Nanni. The weighted list update problem and the lazy adversary. Theoretical Computer Science, 108(2):371–384, February 15, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1993&volume=108&issue=2&aid=1231.

Lehr:1993:SRM

[693] S. Lehr. Sums and rational multiples of q-automatic sequences are qautomatic. Theoretical Computer Science, 108(2):385–391, February 15, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1993&volume=108&issue=2&aid=1281.

Hromkovic:1993:NRO

[694] Juraj Hromkovič and Katsushi Inoue. A note on realtime one-way synchronized alternating one-counter automata. Theoretical Computer Science, 108(2):393–400, February 15, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1993&volume=108&issue=2&aid=1265.

Anonymous:1993:IWC

[695] Anonymous. International Workshop on Computing by Graph Transformation. Theoretical Computer Science, 109(1-2):??, March 01, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Courcelle:1993:P

[696] B. Courcelle and G. Rozenberg. Preface. Theoretical Computer Science, 109 (1-2):1–??, March 01, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ehrig:1993:EBR

[697] H. Ehrig and M. Löwe. The ESPRIT Basic Research Working Group COM-PUGRAPH: “Computing by Graph Transformation”: a survey. Theoretical Computer Science, 109(1– 2):3–6, March 01, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.
cgi?year=1993&volume=109&issue=
1-2&aid=1255.

Corradini:1993:HRJ

[698] Andrea Corradini and Francesca Rossi. Hyperedge replacement jungle rewriting for term-rewriting systems and logic programming. Theoretical Computer Science, 109(1–2):7–48, March 01, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=109&issue=1-2&aid=1256.

Courcelle:1993:MSO

[699] B. Courcelle and M. Mosbah. Monadic second-order evaluations on treedecomposable graphs. Theoretical Computer Science, 109(1–2):49–82, March 01, 1993. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.
cgi?year=1993&volume=109&issue=
1-2&aid=1257.

Drewes:1993:RCH

[700] Frank Drewes. Recognising kconnected hypergraphs in cubic time. Theoretical Computer Science, 109(1–2):83–122, March 01, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1993&volume=109&issue=1-2&aid=1258.

Ehrig:1993:PDD

[701] H. Ehrig and M. Löwe. Parallel and distributed derivations in the single-pushout approach. Theoretical Computer Science, 109(1–2):123– 143, March 01, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/tcs/cas_sub/browse/browse.
                                        cgi?year=1993&volume=109&issue=
                                        1-2&aid=1259.

Janssens:1993:ECA

[702] D. Janssens. Equivalence of computations in actor grammars. Theoretical Computer Science, 109(1–2): 145–180, March 01, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/tcs/cas_sub/browse/browse.
                                        cgi?year=1993&volume=109&issue=
                                        1-2&aid=1260.

Lowe:1993:AAS

[703] Michael Löwe. Algebraic approach to single-pushout graph transformation. Theoretical Computer Science, 109(1–2):181–224, March 01, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=109&issue=1-2&aid=1261.

Montanari:1993:GRP

[704] Ugo Montanari and Francesca Rossi. Graph rewriting for a partial ordering semantics of concurrent constraint programming. Theoretical Computer Science, 109(1–2):225–256, March 01, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=109&issue=1-2&aid=1262.

Schneider:1993:CGG

[705] H. J. Schneider. On categorical graph grammars integrating structural transformations and operations on labels. Theoretical Computer Science, 109(1–2):257–274, March 01, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=109&issue=1-2&aid=1263.

Engelfriet:1993:AW

[706] Joost Engelfriet and Hendrik Jan Hoogeboom. X-automata on ω-words. Theoretical Computer Science, 110 (1):1–51, March 15, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.
cgi?year=1993&volume=110&issue=
1&aid=1182.

Goles:1993:LSS

[707] Eric Goles, Alejandro Maass, and Servet Martinez. On the limit set of some universal cellular automata. Theoretical Computer Science, 110(1):53–78, March 15, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1993&volume=110&issue=1&aid=1145.

Blizard:1993:DMF

[708] Wayne D. Blizard. Dedekind multisets and function shells. Theoretical Computer Science, 110(1):79–98, March 15, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1993&volume=110&issue=1&aid=1144.

Patt-Shamir:1993:TST

[709] Boaz Patt-Shamir and David Peleg. Time-space tradeoffs for set operations. Theoretical Computer Science, 110(1):99–129, March 15, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1993&volume=110&issue=1&aid=1143.

Freivalds:1993:PII

[710] R. Freivalds, E. B. Kinber, and R. Wiehagen. On the power of inductive inference from good examples. Theoretical Computer Science, 110(1):131–144, March 15, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=110&issue=1&aid=1142.

Habel:1993:CCF

[711] Annegret Habel, Hans-Jörg Kreowski, and Clemens Lautemann. A comparison of compatible, finite, and inductive graph properties. Theoretical Computer Science, 110(1):145–168, March 15, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=110&issue=1&aid=1141.

Zwick:1993:MG

[712] Uri Zwick and Michael S. Paterson. The memory game. Theoretical Computer Science, 110(1):169–196, March 15, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=110&issue=1&aid=1333.

Fraenkel:1993:GMG

[713] A. S. Fraenkel and S. Simonson. Geography (mathematical games). Theoretical Computer Science, 110(1):197–214, March 15, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=110&issue=1&aid=1376.

Bodlaender:1993:CPF

[714] Hans L. Bodlaender. Complexity of path-forming games. Theoretical Computer Science, 110(1):215–245, March 15, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1993&volume=110&issue=1&aid=1375.

Nivat:1993:EIT

[715] M. Nivat. Editorial introducing the Tutorial Section. Theoretical Computer Science, 110(2):247–??, March 29, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gallier:1993:CLP

[716] Jean Gallier. Constructive logics. Part I: A tutorial on proof systems and typed λ-calculi. Theoretical Computer Science, 110(2):249–339, March 29, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1993&volume=110&issue=2&aid=1381.

Charron-Bost:1993:CCD

[717] Bernadette Charron-Bost. Coupling coefficients of a distributed execution. Theoretical Computer Science, 110(2):341–376, March 29, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1993&volume=110&issue=2&aid=1248.

Maher:1993:TSD

[718] Michael J. Maher. A transformation system for deductive database modules with perfect model semantics. Theoretical Computer Science, 110(2):377–403, March 29, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=110&issue=2&aid=1160.

Forster:1993:SCW

[719] Thomas Forster. A semantic characterization of the well-typed formulae of λ-calculus. Theoretical Computer Science, 110(2):405–418, March 29, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=110&issue=2&aid=1215.

Sprenger:1993:HDL

[720] M. Sprenger and M. Wymann-Böni. How to decide the lark (combinatorial logic). Theoretical Computer Science, 110(2):419–432, March 29, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=110&issue=2&aid=1331.

Anonymous:1993:SWM

[721] Anonymous. Sixth Workshop on the Mathematical Foundations of Programming Semantics. Theoretical Computer Science, 111(1–2):??, April 12, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Mislove:1993:F

[722] M. Mislove and R. D. Tennent. Foreword. Theoretical Computer Science, 111(1–2):1–??, April 12, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Abramsky:1993:CIL

[723] Samson Abramsky. Computational interpretations of linear logic. Theoretical Computer Science, 111(1– 2):3–57, April 12, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/tcs/cas_sub/browse/browse.
                                        cgi?year=1993&volume=111&issue=
                                        1-2&aid=1286.

Heckmann:1993:PDS

[724] Reinhold Heckmann. Power domains and second-order predicates. Theoretical Computer Science, 111(1–2): 59–88, April 12, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/tcs/cas_sub/browse/browse.
                                        cgi?year=1993&volume=111&issue=
                                        1-2&aid=1287.

Droste:1993:SD

[725] Manfred Droste. On stable domains. Theoretical Computer Science, 111(1–2):89–101, April 12, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1993&volume=111&issue=1-2&aid=1288.

Lamarche:1993:SDG

[726] François Lamarche. Stable domains are generalized topological spaces. Theoretical Computer Science, 111(1–2):103–123, April 12, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=111&issue=1-2&aid=1289.

Main:1993:CPR

[727] Michael G. Main. Complete proof rules for strong fairness and strong extreme fairness. Theoretical Computer Science, 111(1–2):125–143, April 12, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=111&issue=1-2&aid=1290.

Dybkjaer:1993:CHC

[728] Hans Dybkjær and Austin Melton. Comparing Hagino’s categorical programming language and typed lambdacalculi. Theoretical Computer Science, 111(1–2):145–189, April 12, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=111&issue=1-2&aid=1291.

Moss:1993:MLA

[729] Lawrence S. Moss and Satish R. Thatte. Modal logic and algebraic specifications. Theoretical Computer Science, 111(1–2):191–210, April 12, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1993&volume=111&issue=1-2&aid=1292.

Gilmore:1993:LCT

[730] Paul C. Gilmore and George K. Tsiknis. A logic for category theory. Theoretical Computer Science, 111(1–2):211–252, April 12, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1993&volume=111&issue=1-2&aid=1293.

Gilmore:1993:LFP

[731] Paul C. Gilmore and George K. Tsiknis. Logical foundations for programming semantics. Theoretical Computer Science, 111(1–2):253–290, April 12, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1993&volume=111&issue=1-2&aid=1294.

Rus:1993:E

[732] T. Rus. Editorial. Theoretical Computer Science, 112(1):1–??, April 26, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Janicki:1993:SC

[733] Ryszard Janicki and Maciej Koutny. Structure of concurrency. Theoretical Computer Science, 112(1):5–52, April 26, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/

cas_sub/browse/browse.cgi?year=
1993&volume=112&issue=1&aid=1300.

Srinivas:1993:STA

[734] Yellamraju V. Srinivas. A sheaftheoretic approach to pattern matching and related problems. Theoretical Computer Science, 112(1):53–97, April 26, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=112&issue=1&aid=1301.

Talcott:1993:TBS

[735] Carolyn Talcott. A theory of binding structures and applications to rewriting. Theoretical Computer Science, 112(1):99–143, April 26, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=112&issue=1&aid=1302.

Thomas:1993:SDK

[736] Muffy Thomas and Phil Watson. Solving divergence in Knuth–Bendix completion by enriching signatures. Theoretical Computer Science, 112(1): 145–185, April 26, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.
cgi?year=1993&volume=112&issue=
1&aid=1303.

Herbst:1993:GPF

[737] Thomas Herbst and Richard M. Thomas. Group presentations, formal languages and characterizations of one-counter groups. Theoretical Computer Science, 112(2):187–214, May 10, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=112&issue=2&aid=1314.

Bourgin:1993:SCP

[738] Richard D. Bourgin and Sally E. Howe. Shortest curves in planar regions with curved boundary. Theoretical Computer Science, 112(2):215–253, May 10, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=112&issue=2&aid=1242.

Ogiwara:1993:SHS

[739] Mitsunori Ogiwara and Antoni Lozano. On sparse hard sets for counting classes. Theoretical Computer Science, 112(2):255–275, May 10, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=112&issue=2&aid=1222.

Csuhaj-Varju:1993:DCC

[740] Erzsébet Csuhaj-Varjú and Alica Kelemenová. Descriptional complexity of context-free grammar forms. Theoretical Computer Science, 112(2):277–289, May 10, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=112&issue=2&aid=1224.

Sturtivant:1993:CEF

[741] Carl Sturtivant and Gudmund Skovbjerg Frandsen. The computational efficacy of finite-field arithmetic. Theoretical Computer Science, 112(2):291–309, May 10, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=112&issue=2&aid=1334.

Neraud:1993:DWF

[742] Jean Néraud. Deciding whether a finite set of words has rank at most two. Theoretical Computer Science, 112(2):311–337, May 10, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=112&issue=2&aid=1178.

Boissonnat:1993:RCD

[743] Jean-Daniel Boissonnat and Monique Teillaud. On the randomized construction of the Delaunay tree. Theoretical Computer Science, 112(2):339–354, May 10, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=112&issue=2&aid=1206.

Aslam:1993:LAC

[744] Javed A. Aslam and Aditi Dhagat. On-line algorithms for 2-coloring hypergraphs via chip games. Theoretical Computer Science, 112(2):355–369, May 10, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1993&volume=112&issue=2&aid=1391.

Fraenkel:1993:UEG

[745] Aviezri S. Fraenkel, Edward R. Scheinerman, and Daniel Ullman. Undirected edge geography. Theoretical Computer Science, 112(2):371–382, May 10, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1993&volume=112&issue=2&aid=1377.

Calude:1993:NTS

[746] Cristian Calude and Cezar Câmpeanu. Note on the topological structure of random strings. Theoretical Computer Science, 112(2):383–390, May 10, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1993&volume=112&issue=2&aid=1320.

Ibarra:1993:NSP

[747] Oscar H. Ibarra and Nicholas Q. Trân. A note on simple programs with two variables. Theoretical Computer Science, 112(2):391–397, May 10, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1993&volume=112&issue=2&aid=1316.

Korec:1993:ISC

[748] Ivan Korec. Irrational speeds of configurations growth in generalized Pascal triangles. Theoretical Computer Science, 112(2):399–412, May 10, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=112&issue=2&aid=1279.

Skurczynski:1993:BHI

[749] Jerzy Skurczyński. The Borel hierarchy is infinite in the class of regular sets of trees. Theoretical Computer Science, 112(2):413–418, May 10, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=112&issue=2&aid=1282.

Sykora:1993:ESG

[750] Ondrej Sýkora and Imrich Vrťo. Edge separators for graphs of bounded genus with applications. Theoretical Computer Science, 112(2):419–420, May 10, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=112&issue=2&aid=1373.

Choffrut:1993:E

[751] C. Choffrut and M. Jantzen. Editorial. Theoretical Computer Science, 113 (1):1–??, May 24, 1993. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Albert:1993:ACA

[752] Luc Albert, Rafael Casas, and François Fages. Average-case analysis of unification algorithms. Theoretical Computer Science, 113(1):3–34, May 24, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1993&volume=113&issue=1&aid=1297.

Diekert:1993:CIT

[753] Volker Diekert. On the concatenation of infinite traces. Theoretical Computer Science, 113(1):35–54, May 24, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1993&volume=113&issue=1&aid=1276.

Fortnow:1993:IPS

[754] Lance Fortnow and Carsten Lund. Interactive proof systems and alternating time-space complexity. Theoretical Computer Science, 113(1):55–73, May 24, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1993&volume=113&issue=1&aid=1275.

Huhne:1993:PSQ

[755] Martin Hühne. On the power of several queues (Turing machines). Theoretical Computer Science, 113(1):75–91, May 24, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1993&volume=113&issue=1&aid=1272.

Jeron:1993:TUF

[756] Thierry Jéron and Claude Jard. Testing for unboundedness of fifo channels. Theoretical Computer Science, 113(1):93–117, May 24, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=113&issue=1&aid=1274.

Madlener:1993:WCM

[757] K. Madlener, P. Narendran, F. Otto, and L. Zhang. On weakly confluent monadic string-rewriting systems. Theoretical Computer Science, 113(1): 119–165, May 24, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.
cgi?year=1993&volume=113&issue=
1&aid=1298.

Tarui:1993:PPF

[758] Jun Tarui. Probabilistic polynomials, AC^{0} functions, and the polynomialtime hierarchly. Theoretical Computer Science, 113(1):167–183, May 24, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=113&issue=1&aid=1394.

Weihrauch:1993:CCM

[759] Klaus Weihrauch. Computability on computable metric spaces. Theoretical Computer Science, 113(2):191–210, June 07, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1993&volume=113&issue=2&aid=1250.

Yi:1993:AC

[760] Bo Yi and Jiafu Xu. Analogy calculus. Theoretical Computer Science, 113(2):211–230, June 07, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1993&volume=113&issue=2&aid=1217.

Ruqian:1993:TCM

[761] Lu Ruqian. A true concurrency model of CCS semantics. Theoretical Computer Science, 113(2):231–258, June 07, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1993&volume=113&issue=2&aid=1218.

Lu:1993:APD

[762] Jianguo Lu and Jiafu Xu. Analogical program derivation based on type theory. Theoretical Computer Science, 113(2):259–272, June 07, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1993&volume=113&issue=2&aid=1219.

Bucciarelli:1993:TS

[763] Antonio Bucciarelli and Thomas Ehrhard. A theory of sequentiality. Theoretical Computer Science, 113(2): 273–291, June 07, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.
cgi?year=1993&volume=113&issue=
2&aid=1221.

vanGlabbeek:1993:MSP

[764] Rob J. van Glabbeek and Frits W. Vaandrager. Modular specification of process algebras. Theoretical Computer Science, 113(2):293–348, June 07, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=113&issue=2&aid=1220.

Masseron:1993:GPLa

[765] M. Masseron, C. Tollu, and J. Vauzeilles. Generating plans in linear logic I. Actions as proofs. Theoretical Computer Science, 113(2):349–370, June 07, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=113&issue=2&aid=1266.

Masseron:1993:GPLb

[766] M. Masseron. Generating plans in linear logic. II. A geometry of conjunctive actions. Theoretical Computer Science, 113(2):371–375, June 07, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=113&issue=2&aid=1296.

Anonymous:1993:WCC

[767] Anonymous. 3rd Workshop on Concurrency and Compositionality. Theoretical Computer Science, 114(1):??, June 14, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Best:1993:P

[768] E. Best and G. Rozenberg. Preface. Theoretical Computer Science, 114(1): 1–??, June 14, 1993. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Abadi:1993:LVC

[769] Martín Abadi and Gordon D. Plotkin. A logical view of composition. Theoretical Computer Science, 114(1):3–30, June 14, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1993&volume=114&issue=1&aid=1305.

Boudol:1993:OL

[770] G. Boudol, I. Castellani, M. Hennessy, and A. Kiehn. Observing localities. Theoretical Computer Science, 114(1):31–61, June 14, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1993&volume=114&issue=1&aid=1306.

Degano:1993:UAB

[771] Pierpaolo Degano, Rocco De Nicola, and Ugo Montanari. Universal axioms for bisimulations. Theoretical Computer Science, 114(1):63–91, June 14, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=114&issue=1&aid=1307.

Desel:1993:RCE

[772] Jörg Desel and Javier Esparza. Reachability in cyclic extended freechoice systems. Theoretical Computer Science, 114(1):93–118, June 14, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=114&issue=1&aid=1308.

Larsen:1993:EPI

[773] Kim Guldstrand Larsen. The expressive power of implicit specifications. Theoretical Computer Science, 114(1):119–147, June 14, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=114&issue=1&aid=1309.

Milner:1993:MLM

[774] Robin Milner, Joachim Parrow, and David Walker. Modal logics for mobile processes. Theoretical Computer Science, 114(1):149–171, June 14, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=114&issue=1&aid=1310.

Vogler:1993:BAR

[775] Walter Vogler. Bisimulation and action refinement. Theoretical Computer Science, 114(1):173–200, June 14, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1993&volume=114&issue=1&aid=1311.

Vickers:1993:ISC

[776] Steven Vickers. Information systems for continuous posets. Theoretical Computer Science, 114(2):201–229, June 21, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1993&volume=114&issue=2&aid=1283.

Eiter:1993:PCE

[777] Thomas Eiter and Georg Gottlob. Propositional circumscription and extended closed-world reasoning are Π_{2}^{P} -complete. Theoretical Computer Science, 114(2):231–245, June 21, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1993&volume=114&issue=2&aid=1313.

See erratum [857].

Desharnais:1993:RDO

[778] Jules Desharnais, Ali Jaoua, Fatma Mili, Noureddine Boudriga, and Ali Mili. A relational division operator: the conjugate kernel. Theoretical Computer Science, 114(2):247–272, June 21, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=114&issue=2&aid=1246.

Dougherty:1993:HOU

[779] Daniel J. Dougherty. Higher-order unification via combinators. Theoretical Computer Science, 114(2):273–298, June 21, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=114&issue=2&aid=1225.

Barr:1993:TCW

[780] Michael Barr. Terminal coalgebras in well-founded set theory. Theoretical Computer Science, 114(2):299–315, June 21, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=114&issue=2&aid=1245.

Elvang-Goransson:1993:SSC

[781] Morten Elvang-Gøransson and Olaf Owe. A simple sequent calculus for partial functions. Theoretical Computer Science, 114(2):317–330, June 21, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=114&issue=2&aid=1387.

Abramsky:1993:P

[782] S. Abramsky and P.-L Curien. Preface. Theoretical Computer Science, 115 (1):1–??, July 05, 1993. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Blute:1993:LLC

[783] Richard Blute. Linear logic, coherence and dinaturality. Theoretical Computer Science, 115(1):3–41, July 05, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=115&issue=1&aid=1358.

Burroni:1993:HDW

[784] Albert Burroni. Higher-dimensional word problems with applications to equational logic. Theoretical Computer Science, 115(1):43–62, July 05, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=115&issue=1&aid=1359.

Coquand:1993:API

[785] Thierry Coquand. Another proof of the intuitionistic Ramsey theorem. Theoretical Computer Science, 115 (1):63–75, July 05, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.
cgi?year=1993&volume=115&issue=
1&aid=1360.

Edalat:1993:CFS

[786] Abbas Edalat and Michael B. Smyth. I-categories as a framework for solving domain equations. Theoretical Computer Science, 115(1):77–106, July 05, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=115&issue=1&aid=1362.

Freyd:1993:SP

[787] P. Freyd. Structural polymorphism. Theoretical Computer Science, 115(1): 107–129, July 05, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.
cgi?year=1993&volume=115&issue=
1&aid=1361.

Jarzembski:1993:PPA

[788] Grzegorz Jarzembski. Programs in partial algebras. Theoretical Computer Science, 115(1):131–149, July 05, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=115&issue=1&aid=1363.

Jay:1993:TRT

[789] C. Barry Jay. Tail recursion through universal invariants. Theoretical Computer Science, 115(1):151–189, July 05, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=115&issue=1&aid=1364.

Litovsky:1993:CGR

[790] Igor Litovsky and Yves Métivier. Computing with graph rewriting systems with priorities. Theoretical Computer Science, 115(2):191–224, July 19, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=115&issue=2&aid=1278.

Allender:1993:AEC

[791] Eric Allender, Richard Beigel, Ulrich Hertrampf, and Steven Homer. Almost-everywhere complexity hierarchies for nondeterministic time. Theoretical Computer Science, 115(2): 225–241, July 19, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.
cgi?year=1993&volume=115&issue=
2&aid=1380.

Carpi:1993:OFW

[792] Arturo Carpi. Overlap-free words and finite automata. Theoretical Computer Science, 115(2):243–260, July 19, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=115&issue=2&aid=1317.

Ibarra:1993:SFA

[793] Oscar H. Ibarra and Nicholas Q. Trân. Synchronized finite automata and 2DFA reductions. Theoretical Computer Science, 115(2):261–275, July 19, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/

cas_sub/browse/browse.cgi?year=
1993&volume=115&issue=2&aid=1237.

Longpre:1993:HPP

[794] Luc Longpré and Alan L. Selman. Hard promise problems and nonuniform complexity. Theoretical Computer Science, 115(2):277–290, July 19, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=115&issue=2&aid=1236.

Troesch:1993:IGA

[795] A. Troesch. Interpretation geométrique de l’algorithme d’Euclide et reconnaissance de segments. (French) [Geometric interpretation of Euclid’s algorithm with recognition segments]. Theoretical Computer Science, 115(2):291–319, July 19, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=115&issue=2&aid=1235.

Goles:1993:GLG

[796] Eric Goles and Marcos A. Kiwi. Games on line graphs and sand piles. Theoretical Computer Science, 115(2):321–349, July 19, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=115&issue=2&aid=1395.

Srikant:1993:OPC

[797] R. Srikant, Ravi Sundaram, Karan Sher Singh, and C. Pandu Rangan. Optimal path cover problem on block graphs and bipartite permutation graphs. Theoretical Computer Science, 115(2):351–357, July 19, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=115&issue=2&aid=1372.

Katsura:1993:SAD

[798] Masashi Katsura and Yuji Kobayashi. The shuffle algebra and its derivations. Theoretical Computer Science, 115(2):359–369, July 19, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=115&issue=2&aid=1315.

Tang:1993:ETS

[799] Shouwen Tang, Bin Fu, and Tian Liu. Exponential-time and subexponentialtime sets. Theoretical Computer Science, 115(2):371–381, July 19, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=115&issue=2&aid=1322.

Homer:1993:TTH

[800] Steven Homer, Stuart Kurtz, and James Royer. On 1-truth-table-hard languages. Theoretical Computer Science, 115(2):383–389, July 19, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/

cas_sub/browse/browse.cgi?year=
1993&volume=115&issue=2&aid=1319.

Vagvolgyi:1993:FAC

[801] Sándor Vágvölgyi. A fast algorithm for constructing a tree automaton recognizing a congruential tree language. Theoretical Computer Science, 115(2):391–399, July 19, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=115&issue=2&aid=1325.

Abiteboul:1993:F

[802] S. Abiteboul and P. Kanellakis. Foreword. Theoretical Computer Science, 116(1):1–??, August 02, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fagin:1993:FMT

[803] Ronald Fagin. Finite-model theory — a personal perspective. Theoretical Computer Science, 116(1):3–31, August 02, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=116&issue=1&aid=1346.

Kuper:1993:CQL

[804] Gabriel M. Kuper and Moshe Y. Vardi. On the complexity of queries in the logical data model. Theoretical Computer Science, 116(1):33–57, August 02, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/

cas_sub/browse/browse.cgi?year=
1993&volume=116&issue=1&aid=1347.

Beeri:1993:AOO

[805] Catriel Beeri and Yoram Kornatzky. Algebraic optimization of objectoriented query languages. Theoretical Computer Science, 116(1):59–94, August 02, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=116&issue=1&aid=1348.

Consens:1993:LCA

[806] Mariano P. Consens and Alberto O. Mendelzon. Low-complexity aggregation in GraphLog and Datalog. Theoretical Computer Science, 116(1):95–116, August 02, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=116&issue=1&aid=1349.

Revesz:1993:CFE

[807] Peter Z. Revesz. A closed-form evaluation for Datalog queries with integer (gap)-order constraints. Theoretical Computer Science, 116(1): 117–149, August 02, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.
cgi?year=1993&volume=116&issue=
1&aid=1343.

vanderMeyden:1993:RID

[808] Ron van der Meyden. Recursively indefinite databases. Theoretical Computer Science, 116(1):151–194, August 02, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=116&issue=1&aid=1344.

Lipton:1993:ESS

[809] Richard J. Lipton, Jeffrey F. Naughton, Donovan A. Schneider, and S. Seshadri. Efficient sampling strategies for relational database operations. Theoretical Computer Science, 116(1):195–226, August 02, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=116&issue=1&aid=1345.

Ehrenfeucht:1993:SFT

[810] A. Ehrenfeucht and G. Rozenberg. T -structures, T -functions and texts. Theoretical Computer Science, 116(2):227–290, August 16, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=116&issue=2&aid=1338.

Bruderlin:1993:UGR

[811] Beat Brüderlin. Using geometric rewrite rules for solving geometric problems symbolically. Theoretical Computer Science, 116(2):291– 303, August 16, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.

                                        cgi?year=1993&volume=116&issue=
                                        2&aid=1280.

Karhumaki:1993:ECT

[812] Juhani Karhumäki, Wojciech Rytter, and Stefan Jarominek. Efficient constructions of test sets for regular and context-free languages. Theoretical Computer Science, 116(2):305–316, August 16, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1993&volume=116&issue=2&aid=1239.

Saouter:1993:CRE

[813] Yannick Saouter and Patrice Quinton. Computability of recurrence equations. Theoretical Computer Science, 116(2):317–337, August 16, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1993&volume=116&issue=2&aid=1240.

Andrasiu:1993:LTP

[814] Mircea Andraşiu, Gheorghe Păun, Jürgen Dassow, and Arto Salomaa. Language-theoretic problems arising from Richelieu cryptosystems. Theoretical Computer Science, 116(2):339–357, August 16, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1993&volume=116&issue=2&aid=1238.

Feuerstein:1993:DAS

[815] Esteban Feuerstein and Alberto Marchetti-Spaccamela. Dynamic algorithms for shortest paths in planar graphs. Theoretical Computer Science, 116(2):359–371, August 16, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=116&issue=2&aid=1312.

Culik:1993:AAR

[816] Karel Culik II and Simant Dube. Affine automata and related techniques for generation of complex images. Theoretical Computer Science, 116(2):373–398, August 16, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=116&issue=2&aid=1241.

Koutsoupias:1993:IKT

[817] Elias Koutsoupias. Improvements on Khrapchenko’s theorem. Theoretical Computer Science, 116(2):399– 403, August 16, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.
cgi?year=1993&volume=116&issue=
2&aid=1378.

Buning:1993:GHF

[818] Hans Kleine Büning. On generalized Horn formulas and k-resolution. Theoretical Computer Science, 116(2): 405–413, August 16, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/tcs/cas_sub/browse/browse.
                                        cgi?year=1993&volume=116&issue=
                                        2&aid=1339.

Pudlak:1993:SNB

[819] Pavel Pudlák and Petr Savický. On shifting networks (Boolean circuit). Theoretical Computer Science, 116(2):415–419, August 16, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1993&volume=116&issue=2&aid=1340.

Monien:1993:FRD

[820] Burkhard Monien, Wojciech Rytter, and Leopold Schäpers. Fast recognition of deterministic CFL’s with a smaller number of processors. Theoretical Computer Science, 116(2):421–429, August 16, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1993&volume=116&issue=2&aid=1431.

See corrigendum [954].

Anonymous:1993:CFP

[821] Anonymous. Conference on Formal Power Series and Algebraic Combinatorics. Theoretical Computer Science, 117(1–2):??, August 30, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Delest:1993:F

[822] M. Delest, G. Jacob, and P. Leroux. Forward. Theoretical Computer Science, 117(1–2):1–??, August 30, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Labelle:1993:SAS

[823] Gilbert Labelle. Sur la symétrie et l’asymétrie des structures combinatoires. (French) [On the symmetry and asymmetry of combinatorial structures]. Theoretical Computer Science, 117(1–2):3–22, August 30, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=117&issue=1-2&aid=1408.

Zeilberger:1993:ISI

[824] Doron Zeilberger. Identities in search of identity. Theoretical Computer Science, 117(1–2):23–38, August 30, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=117&issue=1-2&aid=1409.

Anselmo:1993:OFP

[825] Marcella Anselmo. The operation ↑ on formal power series. Theoretical Computer Science, 117(1–2):39–43, August 30, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=117&issue=1-2&aid=1410.

Arques:1993:CCP

[826] Didier Arques and Isabelle Jacques. Classification des cartes pointées de genre 1 et relation fonctionnelle associée. (French) [Classification of genus-1 pointed maps and associated functional relation]. Theoretical Computer Science, 117(1–2):45–65, August 30, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=117&issue=1-2&aid=1411.

Betrema:1993:AAG

[827] J. Bétréma and J. G. Penaud. Animaux et arbres guingois. (French) [Animals and lop-sided trees]. Theoretical Computer Science, 117(1–2):67–89, August 30, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=117&issue=1-2&aid=1412.

Bjorner:1993:MFF

[828] Anders Björner. The Möbius function of factor order. Theoretical Computer Science, 117(1–2):91–98, August 30, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=117&issue=1-2&aid=1413.

Casas:1993:ACA

[829] R. Casas, J. Díaz, and C. Martínez. Average-case analysis on simple families of trees using a balanced probability model. Theoretical Computer Science, 117(1–2):99–112, August 30, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=117&issue=1-2&aid=1414.

Chen:1993:CFG

[830] William Y. C. Chen. Context-free grammars, differential operators and formal power series. Theoretical Computer Science, 117(1–2):113–130, August 30, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=117&issue=1-2&aid=1415.

Chiricota:1993:RSE

[831] Yves Chiricota. Représentation symbolique d’espèces moléculaires. (French) [Symbolic representation of molecular species]. Theoretical Computer Science, 117(1–2):131–136, August 30, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=117&issue=1-2&aid=1416.

Conference on Formal Power Series and Algebraic Combinatorics (Bordeaux, 1991).

Choi:1993:EGY

[832] Seul Hee Choi and Dominique Gouyou-Beauchamps. Enumeration of generalized Young tableaux with bounded height. Theoretical Computer Science, 117(1–2):137–151, August 30, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=117&issue=1-2&aid=1417.

Constantineau:1993:ASC

[833] I. Constantineau. Auto-similarité dans la combinatoire des polynômes orthogonaux. (French) [Auto-similarity in the combination of orthogonal polynomials]. Theoretical Computer Science, 117(1–2):153–167, August 30, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Decoste:1993:SIS

[834] Hélène Décoste. Séries indicatrices et q-séries. (French) [Indicator series and q-series]. Theoretical Computer Science, 117(1–2):169–186, August 30, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=117&issue=1-2&aid=1419.

Conference on Formal Power Series and Algebraic Combinatorics (Bordeaux, 1991).

Dulucq:1993:CAO

[835] S. Dulucq and S. Gire. Complexité d’algorithmes et opérations sur les arbres. (French) [Algorithmic complexity and operations on trees]. Theoretical Computer Science, 117(1–2):187–198, August 30, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=117&issue=1-2&aid=1420.

Ekhad:1993:SEE

[836] Shalosh B. Ekhad. A short, elementary, and easy, WZ proof of the Askey-Gasper inequality that was used by de Branges in his proof of the Bieberbach conjecture. Theoretical Computer Science, 117(1–2):199–202, August 30, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=117&issue=1-2&aid=1421.

Lalanne:1993:ICD

[837] J. C. Lalanne. Sur une involution sur les chemins de Dyck. (French) [On an involution on Dyck paths]. Theoretical Computer Science, 117(1–2):203–215, August 30, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=117&issue=1-2&aid=1422.

Lalonde:1993:BLA

[838] Pierre Lalonde. Bases de Lyndon des algèbres de Lie libres partiellement commutatives. (French) [Lyndon bases for partially commutative free Lie algebras]. Theoretical Computer Science, 117(1–2):217–226, August 30, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=117&issue=1-2&aid=1423.

Lando:1993:PPM

[839] S. K. Lando and A. K. Zvonkin.

Plane and projective meanders. Theoretical Computer Science, 117(1–2): 227–241, August 30, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.
cgi?year=1993&volume=117&issue=
1-2&aid=1424.

Mantaci:1993:DAE

[840] Roberto Mantaci. Sur la distribution des anti-excédances dans le groupe symétrique et dans ses sousgroupes. (French) [On the anti-exceedance distribution in a symmetrical group in its sub-groups]. Theoretical Computer Science, 117(1–2):243–253, August 30, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=117&issue=1-2&aid=1425.

Melancon:1993:CBS

[841] Guy Melançon. Constructions des bases standard des K < A >-modules à droite. (French) [Construction of standard bases for right K < A >- modules]. Theoretical Computer Science, 117(1–2):255–272, August 30, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=117&issue=1-2&aid=1426.

Sagan:1993:CPH

[842] Bruce E. Sagan. Combinatorial proofs of hook generating functions for skew plane partitions. Theoretical Computer Science, 117(1–2):273–287, August 30, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=117&issue=1-2&aid=1427.

Svrtan:1993:NPO

[843] Dragutin Svrtan. New plethysm operation, Chern characters of exterior and symmetric powers with applications to Stiefel-Whitney classes of Grassmannians. Theoretical Computer Science, 117(1–2):289–301, August 30, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=117&issue=1-2&aid=1428.

West:1993:STT

[844] Julian West. Sorting twice through a stack. Theoretical Computer Science, 117(1–2):303–313, August 30, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=117&issue=1-2&aid=1429.

Rovan:1993:I

[845] B. Rovan. Introduction. Theoretical Computer Science, 118(1):1–??, September 13, 1993. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Baruah:1993:FPR

[846] Sanjoy K. Baruah, Rodney R. Howell, and Louis E. Rosier. Feasibility problems for recurring tasks on one processor. Theoretical Computer Science, 118(1):3–20, September 13, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=118&issue=1&aid=1353.

Darondeau:1993:RAE

[847] Philippe Darondeau and Pierpaolo Degano. Refinement of actions in event structures and causal trees. Theoretical Computer Science, 118(1):21– 48, September 13, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.
cgi?year=1993&volume=118&issue=
1&aid=1354.

Geffert:1993:STT

[848] Viliam Geffert. A speed-up theorem without tape compression. Theoretical Computer Science, 118(1):49– 65, September 13, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.
cgi?year=1993&volume=118&issue=
1&aid=1351.

Walukiewicz:1993:GTA

[849] Igor Walukiewicz. Gentzen-type axiomatization for PAL. Theoretical Computer Science, 118(1):67– 79, September 13, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.
cgi?year=1993&volume=118&issue=
1&aid=1355.

Wegener:1993:BHS

[850] Ingo Wegener. BOTTOM-UP-HEAP-SORT, a new variant of HEAP-SORT, beating, on an average, QUICKSORT (if n is not very small). Theoretical Computer Science, 118(1):81–98, September 13, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=118&issue=1&aid=1352.

Deransart:1993:PMD

[851] Pierre Deransart. Proof methods of declarative properties of definite programs. Theoretical Computer Science, 118(2):99–166, September 27, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=118&issue=2&aid=1247.

Comon:1993:CAS

[852] Hubert Comon. Complete axiomatizations of some quotient term algebras. Theoretical Computer Science, 118(2):167–191, September 27, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=118&issue=2&aid=1269.

Stewart:1993:MPC

[853] Iain A. Stewart. Methods for proving completeness via logical reductions. Theoretical Computer Science, 118(2):193–229, September 27, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=118&issue=2&aid=1329.

Draghicescu:1993:UTO

[854] M. Draghicescu and S. Purushothaman. A uniform treatment of order of evaluation and aggregate update. Theoretical Computer Science, 118(2):231–262, September 27, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=118&issue=2&aid=1284.

Groote:1993:TSS

[855] Jan Friso Groote. Transition system specifications with negative premises. Theoretical Computer Science, 118(2):263–299, September 27, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=118&issue=2&aid=1270.

Simpson:1993:CLF

[856] Alex K. Simpson. A characterisation of the least-fixed-point operator by dinaturality. Theoretical Computer Science, 118(2):301–314, September 27, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=118&issue=2&aid=1402.

Eiter:1993:EPC

[857] Thomas Eiter and Georg Gottlob. Erratum: Propositional circumscription and extended closed-world reasoning are Π_{2}^{P} -complete. Theoretical Computer Science, 118(2):315, September 27, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=118&issue=2&aid=1537.

See [777].

Anonymous:1993:SFS

[858] Anonymous. 5th Soviet — French Symposium on Theoretical Computer Science, Methods and Tools for Compilation, and Program Development (Informatika ’91). Theoretical Computer Science, 119(1):??, October 11, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Barbar:1993:ATG

[859] Kablan Barbar. Attributed tree grammars. Theoretical Computer Science, 119(1):3–22, October 11, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=119&issue=1&aid=1441.

Corsini:1993:TIP

[860] M.-M. Corsini and K. Musumbu. Type inference in Prolog: a new approach. Theoretical Computer Science, 119(1):23–38, October 11, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.

                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1993&volume=119&issue=1&aid=1442.

Devienne:1993:WSE

[861] Philippe Devienne, Patrick Lebègue, and Max Dauchet. Weighted systems of equations. Theoretical Computer Science, 119(1):39–62, October 11, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1993&volume=119&issue=1&aid=1443.

Dikovsky:1993:CCP

[862] A. Ja. Dikovsky. On computational complexity of Prolog programs. Theoretical Computer Science, 119(1):63–102, October 11, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1993&volume=119&issue=1&aid=1444.

Dosch:1993:GPD

[863] Walter Dosch. On a generalized product for domains. Theoretical Computer Science, 119(1):103–125, October 11, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1993&volume=119&issue=1&aid=1445.

Habbas:1993:CMP

[864] Zineb Habbas. A complete modal proof system for HAL: the Herbrand agent language. Theoretical Computer Science, 119(1):127–143, October 11, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=119&issue=1&aid=1446.

Letichevsky:1993:CAA

[865] A. A. Letichevsky, J. V. Kapitonova, and S. V. Konozenko. Computations in APS (algebraic programming system). Theoretical Computer Science, 119(1):145–171, October 11, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=119&issue=1&aid=1447.

Nepomniaschy:1993:POV

[866] V. A. Nepomniaschy and A. A. Sulimov. Problem-oriented verification system and its application to linear algebra programs. Theoretical Computer Science, 119(1):173–185, October 11, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=119&issue=1&aid=1448.

Sazonov:1993:HFS

[867] Vladimir Yu Sazonov. Hereditarilyfinite sets, data bases and polynomialtime computability. Theoretical Computer Science, 119(1):187–214, October 11, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=119&issue=1&aid=1449.

Slissenko:1993:FTS

[868] A. O. Slissenko. On fault tolerance of syntax. Theoretical Computer Science, 119(1):215–222, October 11, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1993&volume=119&issue=1&aid=1450.

Valiev:1993:USP

[869] M. K. Valiev. P_{1}^{1}-universality of some propositional logics of concurrent programs. Theoretical Computer Science, 119(1):223–232, October 11, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1993&volume=119&issue=1&aid=1451.

Kutylowski:1993:SVS

[870] Miroslaw Kutylowski. Stack versus sensitivity for one-way automata. Theoretical Computer Science, 119(2):233–245, October 25, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1993&volume=119&issue=2&aid=1253.

Apostolico:1993:EDQ

[871] Alberto Apostolico and Andrzej Ehrenfeucht. Efficient detection of quasiperiodicities in strings. Theoretical Computer Science, 119(2):247–265, October 25, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.

com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=119&issue=2&aid=1252.

Birget:1993:POW

[872] Jean-Camille Birget. Partial orders on words, minimal elements of regular languages, and state complexity. Theoretical Computer Science, 119(2):267–291, October 25, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=119&issue=2&aid=1432.

Zimand:1993:IEN

[873] Marius Zimand. If not empty, NP-P is topologically large. Theoretical Computer Science, 119(2):293–310, October 25, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=119&issue=2&aid=1264.

Stromquist:1993:SCC

[874] Walter Stromquist and Daniel Ullman. Sequential compounds of combinatorial games. Theoretical Computer Science, 119(2):311–321, October 25, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=119&issue=2&aid=1487.

Wolfe:1993:SDG

[875] David Wolfe. Snakes in Domineering games. Theoretical Computer Science, 119(2):323–329, October 25, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=119&issue=2&aid=1491.

Tamassia:1993:DRP

[876] Roberto Tamassia and Ioannis G. Tollis. Dynamic reachability in planar digraphs with one source and one sink. Theoretical Computer Science, 119(2):331–343, October 25, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=119&issue=2&aid=1404.

Litow:1993:ACA

[877] B. Litow and Ph. Dumas. Additive cellular automata and algebraic series. Theoretical Computer Science, 119(2):345–354, October 25, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=119&issue=2&aid=1367.

Oommen:1993:OAL

[878] B. John Oommen and David T. H. Ng. An optimal absorbing list organization strategy with constant memory requirements. Theoretical Computer Science, 119(2):355–361, October 25, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=119&issue=2&aid=1369.

Jiang:1993:CLS

[879] Tao Jiang and Ming Li. On the complexity of learning strings and sequences. Theoretical Computer Science, 119(2):363–371, October 25, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=119&issue=2&aid=1390.

Knill:1993:EAR

[880] E. Knill, P. T. Cox, and T. Pietrzykowski. Equality and abductive residua for Horn clauses. Theoretical Computer Science, 120(1):1–44, November 08, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=120&issue=1&aid=1384.

Nota:1993:LCT

[881] G. Nota, S. Orefice, G. Pacini, F. Ruggiero, and G. Tortora. Legality concepts for three-valued logic programs. Theoretical Computer Science, 120(1):45–68, November 08, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=120&issue=1&aid=1382.

Grabowski:1993:SPP

[882] Michal Grabowski. On the status of proving program properties in effective interpretations. Theoretical Computer Science, 120(1):69–81, November 08, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=120&issue=1&aid=1365.

Baratella:1993:CPW

[883] Stefano Baratella. A class of programs for which SLDNF resolution and NAF rule are complete. Theoretical Computer Science, 120(1):83–99, November 08, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=120&issue=1&aid=1330.

Gastin:1993:PIT

[884] Paul Gastin and Brigitte Rozoy. The poset of infinitary traces. Theoretical Computer Science, 120(1):101– 121, November 08, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.
cgi?year=1993&volume=120&issue=
1&aid=1511.

Cousot:1993:BIA

[885] P. Cousot and R. Cousot. “A la Burstall” intermittent assertions induction principles for proving inevitability properties of programs. Theoretical Computer Science, 120(1):123– 155, November 08, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.
cgi?year=1993&volume=120&issue=
1&aid=1500.

Zhang:1993:CFP

[886] Wenhui Zhang. Cut formulas in propositional logic. Theoretical Computer Science, 120(1):157–168, November 08, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=120&issue=1&aid=1399.

Marcoux:1993:CAQ

[887] Yves Marcoux. Composition is almost (but not quite) as good as s − 1 − 1. Theoretical Computer Science, 120(2):169–195, November 22, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=120&issue=2&aid=1397.

Bruggemann-Klein:1993:REF

[888] Anne Brüggemann-Klein. Regular expressions into finite automata. Theoretical Computer Science, 120(2):197–213, November 22, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=120&issue=2&aid=1327.

Chen:1993:DAR

[889] Gen-Huey H. Chen, Biing-Feng F. Wang, and Hungwen Li. Deriving algorithms on reconfigurable networks based on function decomposition. Theoretical Computer Science, 120(2):215–227, November 22, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=120&issue=2&aid=1328.

Honkala:1993:DSI

[890] Juha Honkala. On D0L systems with immigration. Theoretical Computer Science, 120(2):229–245, November 22, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=120&issue=2&aid=1323.

Kim:1993:SSE

[891] Changwook Kim and Dong Hoon Lee. Separating k-separated eNCE graph languages. Theoretical Computer Science, 120(2):247–259, November 22, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=120&issue=2&aid=1326.

Subbiah:1993:ODA

[892] Rajanarayanan Subbiah, Sitharama S. Iyengar, Sridhar Radhakrishnan, and R. L. Kashyap. An optimal distributed algorithm for recognizing meshconnected networks. Theoretical Computer Science, 120(2):261–278, November 22, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=120&issue=2&aid=1321.

Fu:1993:SDN

[893] Bin Fu and Hong zhou Li. On symmetric differences of NP-hard sets with weakly P-selective sets. Theoretical Computer Science, 120(2):279– 291, November 22, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.
cgi?year=1993&volume=120&issue=
2&aid=1438.

Paun:1993:CPS

[894] Gheorghe Păun and Arto Salomaa. Closure properties of slender languages. Theoretical Computer Science, 120(2):293–301, November 22, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=120&issue=2&aid=1406.

Sikkel:1993:PLP

[895] Klaas Sikkel. Parallel on-line parsing in constant time per word. Theoretical Computer Science, 120(2):303– 310, November 22, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.
cgi?year=1993&volume=120&issue=
2&aid=1490.

Ferreira:1993:SEA

[896] A. Ferreira. On space-efficient algorithms for certain NP-complete problems. Theoretical Computer Science, 120(2):311–315, November 22, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1993&volume=120&issue=2&aid=1474.

Abadi:1993:SSB

[897] M. Abadi, L. Cardelli, and P.-L Curien. A short scientific biography of Corrado Bohm. Theoretical Computer Science, 121(1–2):1–??, December 06, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Abadi:1993:FPP

[898] M. Abadi, L. Cardelli, and P.-L. Curien. Formal parametric polymorphism. Theoretical Computer Science, 121(1–2):9–58, December 06, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1993&volume=121&issue=1-2&aid=1533.

Barendregt:1993:CPR

[899] Henk Barendregt. Constructive proofs of the range property in lambda calculus. Theoretical Computer Science, 121(1–2):59–69, December 06, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1993&volume=121&issue=1-2&aid=1518.

Berarducci:1993:SNR

[900] Alessandro Berarducci and Benedetto Intrigila. Some new results on easy lambda-terms. Theoretical Computer Science, 121(1–2):71–88, December 06, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=121&issue=1-2&aid=1530.

Constable:1993:CFB

[901] Robert L. Constable and Scott F. Smith. Computational foundations of basic recursive function theory. Theoretical Computer Science, 121(1–2):89–112, December 06, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=121&issue=1-2&aid=1521.

Coppo:1993:TIA

[902] Mario Coppo and Alberto Ferrari. Type inference, abstract interpretation and strictness analysis. Theoretical Computer Science, 121(1–2):113–143, December 06, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=121&issue=1-2&aid=1529.

Huet:1993:ABT

[903] Gérard Huet. An analysis of Böhm’s theorem. Theoretical Computer Science, 121(1–2):145–167, December 06, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=121&issue=1-2&aid=1523.

Jacopini:1993:GRF

[904] G. Jacopini and G. Sontacchi. General recursive functions in a very simply interpretable typed λ-calculus. Theoretical Computer Science, 121(1–2): 169–178, December 06, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/tcs/cas_sub/browse/browse.
                                        cgi?year=1993&volume=121&issue=
                                        1-2&aid=1528.

Brookes:1993:HIC

[905] Stephen Brookes. Historical introduction to “Concrete domains” by G. Kahn and G. D. Plotkin. Theoretical Computer Science, 121(1–2):179– 186, December 06, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/tcs/cas_sub/browse/browse.
                                        cgi?year=1993&volume=121&issue=
                                        1-2&aid=1532.

Kahn:1993:CDP

[906] G. Kahn and G. D. Plotkin. Concrete domains (programming languages semantics). Theoretical Computer Science, 121(1–2):187–277, December 06, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1993&volume=121&issue=1-2&aid=1527.

Klop:1993:CRS

[907] Jan Willem Klop, Vincent van Oostrom, and Femke van Raamsdonk. Combinatory reduction systems: introduction and survey. Theoretical Computer Science, 121(1–2):279–308, December 06, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=121&issue=1-2&aid=1520.

Leivant:1993:FFA

[908] Daniel Leivant. Functions over free algebras definable in the simply typed lambda calculus. Theoretical Computer Science, 121(1–2):309–321, December 06, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1993&volume=121&issue=1-2&aid=1525.

Longo:1993:GTP

[909] Giuseppe Longo, Kathleen Milsted, and Sergei Soloviev. The genericity theorem and parametricity in the polymorphic λ-calculus. Theoretical Computer Science, 121(1–2):323– 349, December 06, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.
cgi?year=1993&volume=121&issue=
1-2&aid=1524.

Plotkin:1993:STO

[910] Gordon D. Plotkin. Set-theoretical and other elementary models of the λ-calculus. Theoretical Computer Science, 121(1–2):351–409, December 06, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/

                                        cas_sub/browse/browse.cgi?year=
                                        1993&volume=121&issue=1-2&aid=1526.

Scott:1993:TTA

[911] Dana S. Scott. A type-theoretical alternative to ISWIM, CUCH, OWHY. Theoretical Computer Science, 121 (1–2):411–440, December 06, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL

                                        http://www.elsevier.com/cgi-bin/
                                        cas/tree/store/tcs/cas_sub/browse/
                                        browse.cgi?year=1993&volume=121&
                                        issue=1-2&aid=1531.

Statman:1993:SEN

[912] Rick Statman. Some examples of non-existent combinators. Theoretical Computer Science, 121(1–2):441– 448, December 06, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/tcs/cas_sub/browse/browse.
                                        cgi?year=1993&volume=121&issue=
                                        1-2&aid=1519.

Goto:1994:PSI

[913] S. Goto and K. Satoh. Preface to the Special Issue of Selected Papers of the Fifth Generation Computer Systems Conference, Tokyo, June 1–5, 1992. Theoretical Computer Science, 122(1– 2):1–??, January 03, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bossi:1994:CSL

[914] A. Bossi, M. Gabbrielli, G. Levi, and M. C. Meo. A compositional semantics for logic programs. Theoretical Computer Science, 122(1–2):3–47, January 03, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=122&issue=1-2&aid=1453.

Pereira:1994:ACW

[915] Luís Moniz Pereira, José J. Alferes, and Joaquim N. Aparício. Adding closed world assumptions to wellfounded semantics. Theoretical Computer Science, 122(1–2):49–68, January 03, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=122&issue=1-2&aid=1454.

Kawamura:1994:LPS

[916] Tadashi Kawamura. Logic program synthesis from first-order logic specifications. Theoretical Computer Science, 122(1–2):69–96, January 03, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=122&issue=1-2&aid=1455.

Martens:1994:SCP

[917] Bern Martens, Danny De Schreye, and Tamás Horváth. Sound and complete partial deduction with unfolding based on well-founded measures. Theoretical Computer Science, 122(1–2):97–117, January 03, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=122&issue=1-2&aid=1456.

Tatsuta:1994:RIC

[918] Makoto Tatsuta. Realizability interpretation of coinductive definitions and program synthesis with streams. Theoretical Computer Science, 122(1–2):119–136, January 03, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1994&volume=122&issue=1-2&aid=1457.

Takayama:1994:DCP

[919] Yukihide Takayama. Defining concurrent processes constructively. Theoretical Computer Science, 122(1–2): 137–164, January 03, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/tcs/cas_sub/browse/browse.
                                        cgi?year=1994&volume=122&issue=
                                        1-2&aid=1458.

Corradini:1994:AMC

[920] Andrea Corradini, Ugo Montanari, and Francesca Rossi. An abstract machine for concurrent modular systems: CHARM. Theoretical Computer Science, 122(1–2):165–200, January 03, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1994&volume=122&issue=1-2&aid=1459.

Poirriez:1994:MST

[921] V. Poirriez. MLOG: a strongly typed confluent functional language with logical variables. Theoretical Computer Science, 122(1–2):201–223, January 03, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=122&issue=1-2&aid=1460.

Denecker:1994:DAM

[922] Marc Denecker and Danny De Schreye. On the duality of abduction and model generation in a framework for model generation with equality. Theoretical Computer Science, 122(1–2):225–262, January 03, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=122&issue=1-2&aid=1461.

Ait-Kaci:1994:FCS

[923] Hassan Aït-Kaci, Andreas Podelski, and Gert Smolka. A feature constraint system for logic programming with entailment. Theoretical Computer Science, 122(1–2):263–283, January 03, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=122&issue=1-2&aid=1462.

Anonymous:1994:JMI

[924] Anonymous. 10th ‘Journées Mathématiques-Informatique’ (mathematics — computer science). Theoretical Computer Science, 123(1):??, January 17, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Allouche:1994:NCT

[925] Jean-Paul Allouche. Note on the cyclic Towers of Hanoi. Theoretical Computer Science, 123(1):3–7, January 17, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1994&volume=123&issue=1&aid=1463.

Assous:1994:FLP

[926] Marc Roland Assous, Vincent Bouchitté, Christine Charretton, and Brigitte Rozoy. Finite labelling problem in event structures. Theoretical Computer Science, 123(1):9–19, January 17, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1994&volume=123&issue=1&aid=1464.

Bertrand:1994:CDM

[927] A. Bertrand. Sur une conjecture d’Yves Métivier. (French) [On a conjecture of Yves Métivier (rewriting system)]. Theoretical Computer Science, 123(1):21–30, January 17, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1994&volume=123&issue=1&aid=1514.

Bertrand:1994:NSF

[928] Claude Bertrand. A natural semantics of first-order type dependency. Theoretical Computer Science, 123(1):31–53, January 17, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=123&issue=1&aid=1515.

Blanchard:1994:QPE

[929] F. Blanchard and S. Fabre. Quelques procédés engendrant des suites infinies. (French) [Some production procedures of infinite sequences]. Theoretical Computer Science, 123(1):55–60, January 17, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=123&issue=1&aid=1465.

Borel:1994:SRP

[930] Jean-Pierre Borel. Symbolic representation of piecewise linear functions on the unit interval and application to discrepancy. Theoretical Computer Science, 123(1):61–87, January 17, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=123&issue=1&aid=1466.

Chaunier:1994:NPI

[931] Claude Chaunier and Nik Lygero¯ s. Le nombre de posets à isomorphie près ayant 12 éléments. (French) [The number of nearly isomorphic posets having 12 elements]. Theoretical Computer Science, 123(1):89–94, January 17, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/

                                        cas_sub/browse/browse.cgi?year=
                                        1994&volume=123&issue=1&aid=1467.

Daude:1994:UBA

[932] Hervé Daudé and Brigitte Vallée. An upper bound on the average number of iterations of the LLL algorithm. Theoretical Computer Science, 123(1):95–115, January 17, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1994&volume=123&issue=1&aid=1468.

Duval:1994:SP

[933] Dominique Duval and Pascale Sénéchaud. Sketches and parametrization. Theoretical Computer Science, 123(1):117– 130, January 17, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/tcs/cas_sub/browse/browse.
                                        cgi?year=1994&volume=123&issue=
                                        1&aid=1469.

Faure:1994:MQM

[934] Henri Faure. Multidimensional quasi Monte-Carlo methods. Theoretical Computer Science, 123(1):131– 137, January 17, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/tcs/cas_sub/browse/browse.
                                        cgi?year=1994&volume=123&issue=
                                        1&aid=1470.

Kowalski:1994:MRU

[935] Zbigniew S. Kowalski. Multiple returns under a bounded number of iterations. Theoretical Computer Science, 123(1):139–144, January 17, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=123&issue=1&aid=1471.

Mignotte:1994:ECI

[936] M. Mignotte. Sur l’équation de Catalan. II. (French) [On Catalan’s equation. II (number theory)]. Theoretical Computer Science, 123(1):145–149, January 17, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=123&issue=1&aid=1472.

Remila:1994:LAT

[937] Eric Rémila. A linear algorithm to tile the trapezes with h_{m} and v_{n}. Theoretical Computer Science, 123(1):151–165, January 17, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=123&issue=1&aid=1473.

Ye:1994:CUE

[938] Xiangdong Ye. Coexistence of uniquely ergodic subsystems of interval mapping. Theoretical Computer Science, 123(1):167–181, January 17, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=123&issue=1&aid=1516.

Huynh:1994:DBN

[939] Dung T. Huynh and Lu Tian. Deciding bisimilarity of normed context-free processes is in Σ_{2}^{p}. Theoretical Computer Science, 123(2):183–197, January 31, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1994&volume=123&issue=2&aid=1374.

Martin:1994:UCA

[940] Bruno Martin. A universal cellular automaton in quasi-linear time and its S–m–n form. Theoretical Computer Science, 123(2):199–237, January 31, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1994&volume=123&issue=2&aid=1370.

Blanchet-Sadri:1994:EMV

[941] F. Blanchet-Sadri. Equations and monoid varieties of dot-depth one and two. Theoretical Computer Science, 123(2):239–258, January 31, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1994&volume=123&issue=2&aid=1336.

Clerbout:1994:AS

[942] M. Clerbout and D. Gonzalez. Atomic semicommutations. Theoretical Computer Science, 123(2):259–272, January 31, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.

com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=123&issue=2&aid=1341.

Birget:1994:WPI

[943] Jean-Camille Birget, Stuart W. Margolis, and John C. Meakin. The word problem for inverse monoids presented by one idempotent relator. Theoretical Computer Science, 123(2):273–289, January 31, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=123&issue=2&aid=1335.

Flajolet:1994:MTA

[944] Philippe Flajolet, Peter Grabner, Peter Kirschenhofer, Helmut Prodinger, and Robert F. Tichy. Mellin transforms and asymptotics: digital sums. Theoretical Computer Science, 123(2):291–314, January 31, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=123&issue=2&aid=1337.

Aoki:1994:THD

[945] Kunimasa Aoki, Juichi Shinoda, and Teruko Tsuda. On Π_{2} theories of hp-T degrees of low sets. Theoretical Computer Science, 123(2):315–327, January 31, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=123&issue=2&aid=1371.

Iwata:1994:OGB

[946] Shigeki Iwata and Takumi Kasai. The Othello game on an n × n board is PSPACE-complete. Theoretical Computer Science, 123(2):329–340, January 31, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1994&volume=123&issue=2&aid=1557.

Mey:1994:FGP

[947] Daniel Mey. Finite games for a predicate logic without contractions. Theoretical Computer Science, 123(2):341–349, January 31, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1994&volume=123&issue=2&aid=1498.

Bennett:1994:DDC

[948] Grahame Bennett. Double dipping: the case of the missing binomial coefficient identities. Theoretical Computer Science, 123(2):351–375, January 31, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1994&volume=123&issue=2&aid=1488.

Kozen:1994:OBC

[949] Dexter Kozen and Shmuel Zaks. Optimal bounds for the change-making problem. Theoretical Computer Science, 123(2):377–388, January 31, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=123&issue=2&aid=1485.

Klawe:1994:SG

[950] Maria M. Klawe. Shallow grates. Theoretical Computer Science, 123(2): 389–395, January 31, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.
cgi?year=1994&volume=123&issue=
2&aid=1484.

Burckel:1994:FEA

[951] S. Burckel. Functional equations associated with congruential functions. Theoretical Computer Science, 123(2):397–406, January 31, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=123&issue=2&aid=1478.

Hemaspaandra:1994:QIR

[952] Edith Hemaspaandra and Lane A. Hemaspaandra. Quasi-injective reductions. Theoretical Computer Science, 123(2):407–413, January 31, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=123&issue=2&aid=1482.

Nishimura:1994:RCP

[953] Naomi Nishimura. Restricted CRCW PRAMs. Theoretical Computer Science, 123(2):415–426, January 31, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=123&issue=2&aid=1483.

Monien:1994:FRD

[954] Burkhard Monien, Wojciech Rytter, and Helmut Schäpers. Fast recognition of deterministic CFL’s with a smaller number of processors (corrigendum). Theoretical Computer Science, 123(2):427–428, January 31, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=123&issue=2&aid=1606.

See [820].

Majster-Cederbaum:1994:CCB

[955] M. E. Majster-Cederbaum and F. Zetzsche. The comparison of a cpobased semantics with a cms-based semantics for CSP. Theoretical Computer Science, 124(1):1–40, February 14, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=124&issue=1&aid=1385.

Berard:1994:GSC

[956] Béatrice Bérard. Global serializability of concurrent programs. Theoretical Computer Science, 124(1):41– 70, February 14, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.

cgi?year=1994&volume=124&issue=
1&aid=1386.

Yamasaki:1994:DSD

[957] Susumu Yamasaki. A denotational semantics and dataflow construction for logic programs. Theoretical Computer Science, 124(1):71–91, February 14, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=124&issue=1&aid=1383.

Codish:1994:BAI

[958] Michael Codish, Dennis Dams, and Eyal Yardeni. Bottom-up abstract interpretation of logic programs. Theoretical Computer Science, 124(1):93–125, February 14, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=124&issue=1&aid=1388.

Thatte:1994:TIP

[959] Satish R. Thatte. Type inference with partial types. Theoretical Computer Science, 124(1):127–148, February 14, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=124&issue=1&aid=1501.

Bergstra:1994:WDT

[960] J. A. Bergstra and J. Heering. Which data types have omegacomplete initial algebra specifications? Theoretical Computer Science, 124(1):149–168, February 14, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=124&issue=1&aid=1504.

Goltz:1994:FPN

[961] Ursula Goltz and Arend Rensink. Finite Petri nets as models for recursive causal behaviour. Theoretical Computer Science, 124(1):169–179, February 14, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=124&issue=1&aid=1509.

Doets:1994:LTT

[962] Kees Doets. Left termination turned into termination. Theoretical Computer Science, 124(1):180–187, February 14, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=124&issue=1&aid=1574.

Barr:1994:ACT

[963] Michael Barr. Additions and corrections to “Terminal coalgebras in wellfounded set theory”. Theoretical Computer Science, 124(1):189–192, February 14, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=124&issue=1&aid=1572.

Pitts:1994:CIP

[964] Andrew M. Pitts. A co-induction principle for recursively defined domains. Theoretical Computer Science, 124(2):195–219, February 28, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=124&issue=2&aid=1503.

More:1994:IBS

[965] Malika More. Investigation of binary spectra by explicit polynomial transformations of graphs. Theoretical Computer Science, 124(2):221–272, February 28, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=124&issue=2&aid=1400.

Hesselink:1994:NRS

[966] Wim H. Hesselink. Nondeterminacy and recursion via stacks and games. Theoretical Computer Science, 124(2):273–295, February 28, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=124&issue=2&aid=1398.

Bossi:1994:NTT

[967] A. Bossi, N. Cocco, and M. Fabris. Norms on terms and their use in proving universal termination of a logic program. Theoretical Computer Science, 124(2):297–328, February 28, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1994&volume=124&issue=2&aid=1401.

Park:1994:CRM

[968] Jung-Heum Park and Kyung-Yong Chwa. On the construction of regular minimal broadcast digraphs. Theoretical Computer Science, 124(2):329– 342, February 28, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/tcs/cas_sub/browse/browse.
                                        cgi?year=1994&volume=124&issue=
                                        2&aid=1389.

Hebrard:1994:LAR

[969] Jean-Jacques Hébrard. A linear algorithm for renaming a set of clauses as a Horn set. Theoretical Computer Science, 124(2):343–350, February 28, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1994&volume=124&issue=2&aid=1560.

Mitrani:1994:P

[970] I. Mitrani and E. Gelenbe. Preface. Theoretical Computer Science, 125(1): 1–??, March 14, 1994. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Coffman:1994:PMP

[971] E. G. Coffman, Jr., Leopold Flatto, Bjorn Poonen, and Paul E. Wright. The processor minimization problem with independent waiting-time constraints. Theoretical Computer Science, 125(1):3–16, March 14, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=125&issue=1&aid=1539.

Combe:1994:OST

[972] M. B. Combé and O. J. Boxma. Optimization of static traffic allocation policies. Theoretical Computer Science, 125(1):17–43, March 14, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=125&issue=1&aid=1540.

Louth:1994:CCL

[973] Graham Louth, Michael Mitzenmacher, and Frank Kelly. Computational complexity of loss networks. Theoretical Computer Science, 125(1): 45–59, March 14, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.
cgi?year=1994&volume=125&issue=
1&aid=1541.

Hofri:1994:AAP

[974] Micha Hofri and Yaakov Kogan. Asymptotic analysis of product-form distributions related to large interconnection networks. Theoretical Computer Science, 125(1):61–90, March 14, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/

                                        cas_sub/browse/browse.cgi?year=
                                        1994&volume=125&issue=1&aid=1542.

Chakka:1994:HMS

[975] Ram Chakka and Isi Mitrani. Heterogeneous multiprocessor systems with breakdowns: Performance and optimal repair strategies. Theoretical Computer Science, 125(1):91–109, March 14, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1994&volume=125&issue=1&aid=1543.

Akyildiz:1994:ESN

[976] Ian F. Akyildiz and Horst von Brand. Exact solutions for networks of queues with blocking-after-service. Theoretical Computer Science, 125(1):111–130, March 14, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1994&volume=125&issue=1&aid=1544.

Gelenbe:1994:VTM

[977] Erol Gelenbe and Marisela Hernández. Virus tests to maximize availability of software systems. Theoretical Computer Science, 125(1):131–147, March 14, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1994&volume=125&issue=1&aid=1545.

Baccelli:1994:DET

[978] François Baccelli, William A. Massey, and Paul E. Wright. Determining the exit time distribution for a closed cyclic network. Theoretical Computer Science, 125(1):149–165, March 14, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=125&issue=1&aid=1546.

Gastin:1994:EKO

[979] Paul Gastin, Antoine Petit, and Wieslaw Zielonka. An extension of Kleene’s and Ochmanski’s theorems to infinite traces. Theoretical Computer Science, 125(2):167–204, March 28, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=125&issue=2&aid=1430.

Middendorf:1994:MCC

[980] Martin Middendorf. More on the complexity of common superstring and supersequence problems. Theoretical Computer Science, 125(2):205–228, March 28, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=125&issue=2&aid=1368.

Lempel:1994:AFS

[981] Mody Lempel and Azaria Paz. An algorithm for finding a shortest vector in a two-dimensional modular lattice. Theoretical Computer Science, 125(2):229–241, March 28, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1994&volume=125&issue=2&aid=1403.

Jiang:1994:SRC

[982] Tao Jiang, Oscar H. Ibarra, and Hui Wang. Some results concerning 2-D on-line tessellation acceptors and 2-D alternating finite automata. Theoretical Computer Science, 125(2):243–257, March 28, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1994&volume=125&issue=2&aid=1392.

Ehrenfeucht:1994:PGC

[983] A. Ehrenfeucht, P. Ten Pas, and G. Rozenberg. Properties of grammatical codes of trees. Theoretical Computer Science, 125(2):259–293, March 28, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1994&volume=125&issue=2&aid=1379.

Wolf:1994:NCS

[984] Marty J. Wolf. Nondeterministic circuits, space complexity and quasigroups. Theoretical Computer Science, 125(2):295–313, March 28, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1994&volume=125&issue=2&aid=1396.

Yu:1994:SCS

[985] Sheng Yu, Qingyu Zhuang, and Kai Salomaa. The state complexities of some basic operations on regular languages. Theoretical Computer Science, 125(2):315–328, March 28, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=125&issue=2&aid=1393.

Goles:1994:LOS

[986] Eric Goles. Lyapunov operators to study the convergence of extremal automata. Theoretical Computer Science, 125(2):329–337, March 28, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=125&issue=2&aid=1563.

Litow:1994:CFL

[987] B. Litow. A context-free language decision problem. Theoretical Computer Science, 125(2):339–343, March 28, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=125&issue=2&aid=1566.

Bafna:1994:AIM

[988] Vineet Bafna, Bala Kalyanasundaram, and Kirk Pruhs. Not all insertion methods yield constant approximate tours in the Euclidean plane. Theoretical Computer Science, 125(2):345–353, March 28, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=125&issue=2&aid=1496.

Matthews:1994:IIA

[989] Robert E. Matthews. An inherently iterative algorithm for the Grzegorczyk hierarchy. Theoretical Computer Science, 125(2):355–360, March 28, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=125&issue=2&aid=1495.

Mateescu:1994:SDC

[990] Alexandru Mateescu. Scattered deletion and commutativity. Theoretical Computer Science, 125(2):361– 371, March 28, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.
cgi?year=1994&volume=125&issue=
2&aid=1494.

Chen:1994:CCC

[991] Pengyuan Chen. The communication complexity of computing differentiable functions in a multicomputer network. Theoretical Computer Science, 125(2):373–383, March 28, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=125&issue=2&aid=1578.

Raoult:1994:P

[992] J.-C Raoult. Preface. Theoretical Computer Science, 126(1):1–??, April 11, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Andersen:1994:MCB

[993] Henrik Reif Andersen. Model checking and Boolean graphs. Theoretical Computer Science, 126(1):3–30, April 11, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=126&issue=1&aid=1547.

Caron:1994:DRD

[994] Anne-Cécile Caron and Jean-Luc Coquide. Decidability of reachability for disjoint union of term rewriting systems. Theoretical Computer Science, 126(1):31–52, April 11, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=126&issue=1&aid=1548.

Courcelle:1994:MSO

[995] Bruno Courcelle. Monadic secondorder definable graph transductions: a survey. Theoretical Computer Science, 126(1):53–75, April 11, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=126&issue=1&aid=1549.

Dam:1994:CEF

[996] Mads Dam. CTL∗ and ECTL∗ as fragments of the modal μcalculus. Theoretical Computer Science, 126(1):77–96, April 11, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=126&issue=1&aid=1550.

Potthoff:1994:MCQ

[997] Andreas Potthoff. Modulo-counting quantifiers over finite trees. Theoretical Computer Science, 126(1): 97–112, April 11, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.
cgi?year=1994&volume=126&issue=
1&aid=1551.

Seidl:1994:FTA

[998] Helmut Seidl. Finite tree automata with cost functions. Theoretical Computer Science, 126(1):113–142, April 11, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=126&issue=1&aid=1552.

Peled:1994:PPO

[999] Doron Peled and Amir Pnueli. Proving partial order properties. Theoretical Computer Science, 126(2): 143–182, April 25, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/

tree/store/tcs/cas_sub/browse/browse.
cgi?year=1994&volume=126&issue=
2&aid=1502.

Alur:1994:TTA

[1000] Rajeev Alur and David L. Dill. A theory of timed automata. Theoretical Computer Science, 126(2): 183–235, April 25, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.
cgi?year=1994&volume=126&issue=
2&aid=1507.

Gerhard:1994:UFD

[1001] J. A. Gerhard and Mario Petrich. Unification in free distributive lattices. Theoretical Computer Science, 126(2):237–257, April 25, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=126&issue=2&aid=1506.

vanOostrom:1994:CDD

[1002] Vincent van Oostrom. Confluence by decreasing diagrams. Theoretical Computer Science, 126(2):259–280, April 25, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=126&issue=2&aid=1405.

Regnier:1994:ELT

[1003] Laurent Regnier. Une equivalence sur les lambda-termes. (French) [Equivalence between lambda-terms]. Theoretical Computer Science, 126(2):

281–292, April 25, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.
cgi?year=1994&volume=126&issue=
2&aid=1591.

Hosono:1994:FDD

[1004] C. Hosono and Y. Ikeda. A formal derivation of the decidability of the theory SA. Theoretical Computer Science, 127(1):1–23, May 09, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=127&issue=1&aid=1439.

Salomaa:1994:STA

[1005] Kai Salomaa. Synchronized tree automata. Theoretical Computer Science, 127(1):25–51, May 09, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=127&issue=1&aid=1437.

Daykin:1994:PRA

[1006] J. W. Daykin, C. S. Iliopoulos, and W. F. Smyth. Parallel RAM algorithms for factorizing words. Theoretical Computer Science, 127(1):53–67, May 09, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=127&issue=1&aid=1475.

Coquide:1994:BTP

[1007] Jean-Luc Coquidé, Max Dauchet, Rémi Gilleron, and Sándor Vágvölgyi. Bottom-up tree pushdown automata: classification and connection with rewrite systems. Theoretical Computer Science, 127(1):69–98, May 09, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=127&issue=1&aid=1436.

Codenotti:1994:OCP

[1008] B. Codenotti, M. Leoncini, and G. Resta. Oracle computations in parallel numerical linear algebra. Theoretical Computer Science, 127(1):99–121, May 09, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=127&issue=1&aid=1435.

Hromkovic:1994:SHC

[1009] Juraj Hromkovič, Jarkko Kari, and Lila Kari. Some hierarchies for the communication complexity measures of cooperating grammar systems. Theoretical Computer Science, 127(1):123–147, May 09, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=127&issue=1&aid=1434.

Jantzen:1994:CCF

[1010] M. Jantzen and H. Petersen. Cancellation in context-free languages: enrichment by reduction. Theoretical Computer Science, 127(1):149–170, May 09, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=127&issue=1&aid=1433.

Inoue:1994:IAT

[1011] Katsushi Inoue, Akira Ito, and Itsuo Takanami. On 1-inkdot alternating Turing machines with small space. Theoretical Computer Science, 127(1):171–179, May 09, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=127&issue=1&aid=1567.

DeAgostino:1994:PCP

[1012] Sergio De Agostino. P-complete problems in data compression. Theoretical Computer Science, 127(1):181–186, May 09, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=127&issue=1&aid=1568.

See erratum [1203].

DelLungo:1994:PDT

[1013] A. Del Lungo. Polyominoes defined by two vectors. Theoretical Computer Science, 127(1):187–198, May 09, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=127&issue=1&aid=1570.

Cameron:1994:BAT

[1014] Helen Cameron and Derick Wood. Balance in AVL trees and space cost of brother trees. Theoretical Computer Science, 127(2):199–228, May 23, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=127&issue=2&aid=1481.

Kari:1994:RTL

[1015] Jarkko Kari. Rice’s theorem for the limit sets of cellular automata. Theoretical Computer Science, 127(2):229–254, May 23, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=127&issue=2&aid=1486.

Khuller:1994:LAW

[1016] Samir Khuller, Stephen G. Mitchell, and Vijay V. Vazirani. On-line algorithms for weighted bipartite matching and stable marriages. Theoretical Computer Science, 127(2):255–267, May 23, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=127&issue=2&aid=1588.

Honkala:1994:GDS

[1017] Juha Honkala. On generalized DT0L systems and their fixed points. Theoretical Computer Science, 127(2):269–286, May 23, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.

                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1994&volume=127&issue=2&aid=1489.

Matos:1994:PSI

[1018] Armando B. Matos. Periodic sets of integers. Theoretical Computer Science, 127(2):287–312, May 23, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1994&volume=127&issue=2&aid=1479.

See remark [1200].

Yunes:1994:SSS

[1019] J. B. Yunès. Seven-state solutions to the Firing Squad Synchronization Problem. Theoretical Computer Science, 127(2):313–332, May 23, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1994&volume=127&issue=2&aid=1477.

Barcucci:1994:RGD

[1020] E. Barcucci, R. Pinzani, and R. Sprugnoli. The random generation of directed animals. Theoretical Computer Science, 127(2):333–350, May 23, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1994&volume=127&issue=2&aid=1476.

Jain:1994:PSR

[1021] Sanjay Jain and Arun Sharma. Program size restrictions in computational learning. Theoretical Computer Science, 127(2):351–386, May 23, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=127&issue=2&aid=1480.

Deutz:1994:HCA

[1022] A. H. Deutz, A. Ehrenfeucht, and G. Rozenberg. Hyperedge channels are abelian. Theoretical Computer Science, 127(2):387–393, May 23, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=127&issue=2&aid=1583.

Hromkovic:1994:NOG

[1023] Juraj Hromkovič, Claus-Dieter Jeschke, and Burkhard Monien. Note on optimal gossiping in some weakconnected graphs. Theoretical Computer Science, 127(2):395–402, May 23, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=127&issue=2&aid=1589.

Aumann:1994:CCF

[1024] Yonatan Aumann and Michael O.

Rabin. Clock construction in fully asynchronous parallel systems and PRAM simulation. Theoretical Computer Science, 128(1–2):3–30, June 06, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/

                                        cas_sub/browse/browse.cgi?year=
                                        1994&volume=128&issue=1-2&aid=1596.

Leighton:1994:MMR

[1025] Tom Leighton. Methods for message routing in parallel machines. Theoretical Computer Science, 128(1–2):31–62, June 06, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1994&volume=128&issue=1-2&aid=1599.

Garofalakis:1994:TDD

[1026] J. Garofalakis, P. Spirakis, B. Tampakas, and S. Rajsbaum. Tentative and definite distributed computations: an optimistic approach to network synchronization. Theoretical Computer Science, 128(1–2):63–74, June 06, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1994&volume=128&issue=1-2&aid=1603.

Koren:1994:MML

[1027] Gilad Koren and Dennis Shasha. MOCA: a multiprocessor on-line competitive algorithm for real-time system scheduling. Theoretical Computer Science, 128(1–2):75–97, June 06, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1994&volume=128&issue=1-2&aid=1602.

Peled:1994:CFF

[1028] Doron Peled and Mathai Joseph. A compositional framework for fault tolerance by specification transformation. Theoretical Computer Science, 128(1–2):99–125, June 06, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=128&issue=1-2&aid=1600.

Schepers:1994:TBC

[1029] Henk Schepers and Jozef Hooman. A trace-based compositional proof theory for fault tolerant distributed systems. Theoretical Computer Science, 128(1–2):127–157, June 06, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=128&issue=1-2&aid=1601.

Krishnan:1994:SCF

[1030] Padmanabhan Krishnan. A semantic characterisation for faults in replicated systems. Theoretical Computer Science, 128(1–2):159–177, June 06, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=128&issue=1-2&aid=1598.

deMeer:1994:GRD

[1031] Hermann de Meer, Kishor S. Trivedi, and Mario Dal Cin. Guarded repair of dependable systems. Theoretical Computer Science, 128(1–2):179–210, June 06, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/

                                        cas_sub/browse/browse.cgi?year=
                                        1994&volume=128&issue=1-2&aid=1604.

Pai:1994:MAM

[1032] Vinay Sadananda Pai, Alejandro A. Schäffer, and Peter J. Varman. Markov analysis of multiple-disk prefetching strategies for external merging. Theoretical Computer Science, 128(1–2):211–239, June 06, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1994&volume=128&issue=1-2&aid=1605.

Bruck:1994:TFM

[1033] Jehoshua Bruck, Robert Cypher, and Ching-Tien Ho. Tolerating faults in a mesh with a row of spare nodes. Theoretical Computer Science, 128(1–2):241–252, June 06, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1994&volume=128&issue=1-2&aid=1597.

Abraham:1994:MEP

[1034] Uri Abraham and Menachem Magidor. On the mutual-exclusion problem — a quest for minimal solutions. Theoretical Computer Science, 129(1):1–38, June 20, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1994&volume=129&issue=1&aid=1512.

Yokouchi:1994:STA

[1035] Hirofumi Yokouchi. F -semantics for type assignment systems. Theoretical Computer Science, 129(1):39–77, June 20, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=129&issue=1&aid=1513.

Krivine:1994:GST

[1036] Jean-Louis L. Krivine. A general storage theorem for integers in call-by-name λ-calculus. Theoretical Computer Science, 129(1):79–94, June 20, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=129&issue=1&aid=1638.

Chen:1994:CCS

[1037] Cheng-Chia Chen and I-Peng Lin. The computational complexity of the satisfiability of modal Horn clauses for modal propositional logics. Theoretical Computer Science, 129(1):95–121, June 20, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=129&issue=1&aid=1508.

Wei:1994:CFT

[1038] Jiawang Wei. Correctness of fixpoint transformations. Theoretical Computer Science, 129(1):123–142, June 20, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.

                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1994&volume=129&issue=1&aid=1505.

Shepherdson:1994:RSA

[1039] J. C. Shepherdson. The role of standardising apart in logic programming. Theoretical Computer Science, 129(1):143–166, June 20, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1994&volume=129&issue=1&aid=1510.

Ognjanovic:1994:TLP

[1040] Zoran Ognjanović. A tableaulike proof procedure for normal modal logics. Theoretical Computer Science, 129(1):167–186, June 20, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1994&volume=129&issue=1&aid=1440.

Banach:1994:RRB

[1041] R. Banach. Regular relations and bicartesian squares. Theoretical Computer Science, 129(1):187–192, June 20, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1994&volume=129&issue=1&aid=1667.

Zhang:1994:DPD

[1042] Wenhui Zhang. Depth of proofs, depth of cut-formulas and complexity of cut formulas. Theoretical Computer Science, 129(1):193–206, June 20, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=129&issue=1&aid=1616.

Deutz:1994:CRS

[1043] A. H. Deutz, A. Ehrenfeucht, and G. Rozenberg. Clans and regions in 2-structures. Theoretical Computer Science, 129(2):207–262, July 04, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=129&issue=2&aid=1565.

Allouche:1994:CPF

[1044] Jean-Paul Allouche and Mireille Bousquet-Mélou. Canonical positions for the factors in paperfolding sequences. Theoretical Computer Science, 129(2):263–278, July 04, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=129&issue=2&aid=1562.

Delyon:1994:ENS

[1045] Bernard Delyon and Oded Maler.

On the effects of noise and speed on computations. Theoretical Computer Science, 129(2):279–291, July 04, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=129&issue=2&aid=1517.

Jaja:1994:EPA

[1046] Joseph F. Jájá and Kwan Woo Ryu. An efficient parallel algorithm for the single function coarsest partition problem. Theoretical Computer Science, 129(2):293–307, July 04, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1994&volume=129&issue=2&aid=1561.

Ganesan:1994:OWF

[1047] K. Ganesan. One-way functions and the isomorphism conjecture. Theoretical Computer Science, 129(2):309–321, July 04, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1994&volume=129&issue=2&aid=1564.

Rich:1994:COF

[1048] Craig A. Rich and Giora Slutzki. The complexity of optimizing finitestate transducers. Theoretical Computer Science, 129(2):323–336, July 04, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1994&volume=129&issue=2&aid=1493.

Knuutila:1994:ITL

[1049] Timo Knuutila and Magnus Steinby. The inference of tree languages from finite samples: an algebraic approach. Theoretical Computer Science, 129(2):337–367, July 04, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=129&issue=2&aid=1492.

Ferenczi:1994:TLR

[1050] S. Ferenczi. Tiling and local rank properties of the Morse sequence. Theoretical Computer Science, 129(2):369–383, July 04, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=129&issue=2&aid=1607.

Scheepers:1994:VGG

[1051] Marion Scheepers. Variations on a game of Gale (II): Markov strategies. Theoretical Computer Science, 129(2): 385–396, July 04, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.
cgi?year=1994&volume=129&issue=
2&aid=1675.

Yellin:1994:ADS

[1052] D. M. Yellin. An algorithm for dynamic subset and intersection testing. Theoretical Computer Science, 129(2):397–406, July 04, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=129&issue=2&aid=1634.

Egecioglu:1994:EUC

[1053] Ömer Eǧecioǧlu and Çetin Kaya Koç. Exponentiation using canonical recoding. Theoretical Computer Science, 129(2):407–417, July 04, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=129&issue=2&aid=1626.

Margenstern:1994:NTM

[1054] M. Margenstern. Nonerasing Turing machines: a frontier between a decidable halting problem and universality. Theoretical Computer Science, 129(2):419–424, July 04, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=129&issue=2&aid=1611.

Woeginger:1994:LSJ

[1055] Gerhard J. Woeginger. On-line scheduling of jobs with fixed start and end times. Theoretical Computer Science, 130(1):5–16, August 01, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=130&issue=1&aid=1614.

Motwani:1994:NS

[1056] Rajeev Motwani, Steven Phillips, and Eric Torng. Nonclairvoyant scheduling. Theoretical Computer Science, 130(1):17–47, August 01, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/

cas_sub/browse/browse.cgi?year=
1994&volume=130&issue=1&aid=1613.

Feldmann:1994:DSP

[1057] Anja Feldmann, Jiří Sgall, and Shang-Hua Teng. Dynamic scheduling on parallel machines. Theoretical Computer Science, 130(1):49–72, August 01, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=130&issue=1&aid=1612.

Azar:1994:LLB

[1058] Yossi Azar, Andrei Z. Broder, and Anna R. Karlin. On-line load balancing. Theoretical Computer Science, 130(1):73–84, August 01, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=130&issue=1&aid=1554.

Fiat:1994:CAW

[1059] Amos Fiat and Moty Ricklin. Competitive algorithms for the weighted server problem. Theoretical Computer Science, 130(1):85–99, August 01, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=130&issue=1&aid=1615.

dAmore:1994:LUP

[1060] Fabrizio d’Amore and Vincenzo Liberatore. The list update problem and the retrieval of sets. Theoretical Computer Science, 130(1):101–123, August 01, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=130&issue=1&aid=1651.

Kalyanasundaram:1994:CCT

[1061] Bala Kalyanasundaram and Kirk R. Pruhs. Constructing competitive tours from local information. Theoretical Computer Science, 130(1):125– 138, August 01, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.
cgi?year=1994&volume=130&issue=
1&aid=1556.

Hershberger:1994:DST

[1062] John Hershberger, Monika Rauch, and Subhash Suri. Data structures for two-edge connectivity in planar graphs. Theoretical Computer Science, 130(1):139–161, August 01, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=130&issue=1&aid=1641.

Halldorsson:1994:LBL

[1063] Magnus M. Halldórsson and Mario Szegedy. Lower bounds for on-line graph coloring. Theoretical Computer Science, 130(1):163–174, August 01, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=130&issue=1&aid=1555.

Alon:1994:LBC

[1064] Noga Alon, Gil Kalai, Moty Ricklin, and Larry Stockmeyer. Lower bounds on the competitive ratio for mobile user tracking and distributed job scheduling. Theoretical Computer Science, 130(1):175–201, August 01, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=130&issue=1&aid=1553.

Miltersen:1994:CMI

[1065] Peter Bro Miltersen, Sairam Subramanian, Jeffrey Scott Vitter, and Roberto Tamassia. Complexity models for incremental computation. Theoretical Computer Science, 130(1):203–236, August 01, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=130&issue=1&aid=1640.

Nivat:1994:GIT

[1066] M. Nivat. Gödel’s incompleteness theorem, by V. A. Uspensky. Theoretical Computer Science, 130(2):237–??, ???? 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Uspensky:1994:GIT

[1067] Vladimir A. Uspensky. Gödel’s incompleteness theorem. Theoretical Computer Science, 130(2):239–319, August 15, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/

                                        cas_sub/browse/browse.cgi?year=
                                        1994&volume=130&issue=2&aid=1608.

Murray:1994:RMP

[1068] Neil V. Murray and Erik Rosenthal. On the relative merits of path dissolution and the method of analytic tableaux. Theoretical Computer Science, 131(1):1–28, August 29, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1994&volume=131&issue=1&aid=1577.

Banach:1994:TGR

[1069] R. Banach. Term graph rewriting and garbage collection using opfibrations. Theoretical Computer Science, 131(1):29–94, August 29, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1994&volume=131&issue=1&aid=1558.

Compton:1994:SLF

[1070] Kevin J. Compton. Stratified least fixpoint logic. Theoretical Computer Science, 131(1):95–120, August 29, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1994&volume=131&issue=1&aid=1579.

Kobayashi:1994:RIG

[1071] Satoshi Kobayashi and Makoto Tatsuta. Realizability interpretation of generalized inductive definitions. Theoretical Computer Science, 131(1):

121–138, August 29, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.
cgi?year=1994&volume=131&issue=
1&aid=1629.

Fix:1994:PCU

[1072] Limor Fix, Nissim Francez, and Orna Grumberg. Program composition via unification. Theoretical Computer Science, 131(1):139–179, August 29, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=131&issue=1&aid=1593.

Aceto:1994:GFL

[1073] Luca Aceto. GSOS and finite labelled transition systems. Theoretical Computer Science, 131(1):181–195, August 29, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=131&issue=1&aid=1621.

Mathieu:1994:KLC

[1074] Philippe Mathieu and Jean-Paul Delahaye. A kind of logical compilation for knowledge bases. Theoretical Computer Science, 131(1):197–218, August 29, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=131&issue=1&aid=1569.

Scholefield:1994:SOS

[1075] David Scholefield, Hussein Zedan, and He Jifeng. A specification-oriented semantics for the refinement of realtime systems. Theoretical Computer Science, 131(1):219–241, August 29, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1994&volume=131&issue=1&aid=1522.

Etzion-Petruschka:1994:SDS

[1076] Yael Etzion-Petruschka, David Harel, and Dale Myers. On the solvability of domino snake problems. Theoretical Computer Science, 131(2):243– 269, September 12, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/tcs/cas_sub/browse/browse.
                                        cgi?year=1994&volume=131&issue=
                                        2&aid=1580.

Squier:1994:FCR

[1077] Craig C. Squier, Friedrich Otto, and Yuji Kobayashi. A finiteness condition for rewriting systems. Theoretical Computer Science, 131(2):271–294, September 12, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1994&volume=131&issue=2&aid=1669.

Idury:1994:DDM

[1078] Ramana M. Idury and Alejandro A. Schäffer. Dynamic dictionary matching with failure functions. Theoretical Computer Science, 131(2):295–310, September 12, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=131&issue=2&aid=1619.

Leiss:1994:LEO

[1079] Ernst L. Leiss. Language equations over a one-letter alphabet with union, concatenation and star: a complete solution. Theoretical Computer Science, 131(2):311–330, September 12, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=131&issue=2&aid=1571.

Siegelmann:1994:ACN

[1080] Hava T. Siegelmann and Eduardo D.

Sontag. Analog computation via neural networks. Theoretical Computer Science, 131(2):331–360, September 12, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=131&issue=2&aid=1576.

Eades:1994:DGT

[1081] Peter Eades and Sue Whitesides. Drawing graphs in two layers. Theoretical Computer Science, 131(2):361– 374, September 12, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.

                                        cgi?year=1994&volume=131&issue=
                                        2&aid=1573.

Gardy:1994:JSU

[1082] Danièle Gardy. Join sizes, urn models, and normal limiting distributions. Theoretical Computer Science, 131(2):375–414, September 12, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1994&volume=131&issue=2&aid=1497.

Spencer:1994:RDA

[1083] J. Spencer. Randomization, derandomization and antirandomization: three games. Theoretical Computer Science, 131(2):415–429, September 12, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1994&volume=131&issue=2&aid=1674.

Eriksson:1994:RDN

[1084] K. Eriksson. Reachability is decidable in the numbers game. Theoretical Computer Science, 131(2):431–439, September 12, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1994&volume=131&issue=2&aid=1609.

Huynh:1994:NCD

[1085] Dung T. Huynh and Lu Tian. A note on the complexity of deciding bisimilarity of normed unary processes. Theoretical Computer Science, 131(2):441–448, September 12, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=131&issue=2&aid=1660.

Bergey:1994:OPT

[1086] A. Bergey and R. Cori. On the orbits of the product of two permutations. Theoretical Computer Science, 131(2):449–461, September 12, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=131&issue=2&aid=1665.

Picouleau:1994:CHC

[1087] C. Picouleau. Complexity of the Hamiltonian cycle in regular graph problem. Theoretical Computer Science, 131(2):463–473, September 12, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=131&issue=2&aid=1639.

Anderson:1994:MEW

[1088] N. Anderson and D. Manley. A matrix extension of Winograd’s inner product algorithm. Theoretical Computer Science, 131(2):475–477, September 12, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=131&issue=2&aid=1681.

Flajolet:1994:CRG

[1089] Philippe Flajolet, Paul Zimmerman, and Bernard Van Cutsem. A calculus for the random generation of labelled combinatorial structures. Theoretical Computer Science, 132(1–2):1–35, September 26, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1994&volume=132&issue=1-2&aid=1632.

Juedes:1994:CDR

[1090] David W. Juedes, James I. Lathrop, and Jack H. Lutz. Computational depth and reducibility. Theoretical Computer Science, 132(1–2): 37–70, September 26, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/tcs/cas_sub/browse/browse.
                                        cgi?year=1994&volume=132&issue=
                                        1-2&aid=1655.

Leiss:1994:UCL

[1091] E. L. Leiss. Unrestricted complementation in language equations over a one-letter alphabet. Theoretical Computer Science, 132(1–2):71–84, September 26, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1994&volume=132&issue=1-2&aid=1623.

Kim:1994:RBP

[1092] Changwook Kim. Retreat bounded picture languages. Theoretical Computer Science, 132(1–2):85–112, September 26, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=132&issue=1-2&aid=1617.

Koiran:1994:CLD

[1093] Pascal Koiran, Michel Cosnard, and Max Garzon. Computability with low-dimensional dynamical systems. Theoretical Computer Science, 132(1–2):113–128, September 26, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=132&issue=1-2&aid=1585.

Kari:1994:LEI

[1094] Lila Kari. On language equations with invertible operations. Theoretical Computer Science, 132(1–2):129– 150, September 26, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.
cgi?year=1994&volume=132&issue=
1-2&aid=1584.

Bonizzoni:1994:PSP

[1095] Paola Bonizzoni. Primitive 2-structures with the (n − 2)-property. Theoretical Computer Science, 132(1–2):151– 178, September 26, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.
cgi?year=1994&volume=132&issue=
1-2&aid=1586.

Pighizzini:1994:AAV

[1096] Giovanni Pighizzini. Asynchronous automata versus asynchronous cellular automata. Theoretical Computer Science, 132(1–2):179–207, September 26, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1994&volume=132&issue=1-2&aid=1587.

Ehrenfeucht:1994:SGT

[1097] A. Ehrenfeucht and R. McConnell. A k-structure generalization of the theory of 2-structures. Theoretical Computer Science, 132(1–2):209–227, September 26, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1994&volume=132&issue=1-2&aid=1582.

Ambos-Spies:1994:MPC

[1098] Klaus Ambos-Spies, Steven Homer, and Robert I. Soare. Minimal pairs and complete problems. Theoretical Computer Science, 132(1–2):229– 241, September 26, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/tcs/cas_sub/browse/browse.
                                        cgi?year=1994&volume=132&issue=
                                        1-2&aid=1622.

Birget:1994:FLD

[1099] Jean-Camille Birget and Joseph B. Stephen. Formal languages defined by uniform substitutions. Theoretical Computer Science, 132(1–2):243– 258, September 26, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.
cgi?year=1994&volume=132&issue=
1-2&aid=1594.

Roka:1994:OWC

[1100] Zsuzsanna Róka. One-way cellular automata on Cayley graphs. Theoretical Computer Science, 132(1–2):259– 290, September 26, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.
cgi?year=1994&volume=132&issue=
1-2&aid=1581.

DeSantis:1994:KCQ

[1101] Alfredo De Santis, Giovanni Di Crescenzo, and Guiseppe Persiano. The knowledge complexity of quadratic residuosity languages. Theoretical Computer Science, 132(1–2):291–317, September 26, 1994. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.
cgi?year=1994&volume=132&issue=
1-2&aid=1595.

Hromkovic:1994:DVN

[1102] Juraj Hromkovič, Branislav Rovan, and Anna Slobodova. Deterministic versus nondeterministic space in terms of synchronized alternating machines. Theoretical Computer Science, 132(1–2):319–336, September 26, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/

                                        cas_sub/browse/browse.cgi?year=
                                        1994&volume=132&issue=1-2&aid=1590.

Diestel:1994:DGI

[1103] R. Diestel and I. Leader. Domination games on infinite graphs. Theoretical Computer Science, 132(1–2): 337–345, September 26, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/tcs/cas_sub/browse/browse.
                                        cgi?year=1994&volume=132&issue=
                                        1-2&aid=1673.

Shufelt:1994:GHC

[1104] Jefferey A. Shufelt and Hans J. Berliner. Generating Hamiltonian circuits without backtracking from errors. Theoretical Computer Science, 132(1–2):347–375, September 26, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1994&volume=132&issue=1-2&aid=1592.

Ben-Amram:1994:UCP

[1105] Amir M. Ben-Amram. Unit-cost pointers versus logarithmic-cost addresses. Theoretical Computer Science, 132(1–2):377–385, September 26, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1994&volume=132&issue=1-2&aid=1720.

Bridges:1994:RBE

[1106] Douglas Bridges and Cristian Calude. On recursive bounds for the exceptional values in speed-up. Theoretical Computer Science, 132(1–2):387– 394, September 26, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.
cgi?year=1994&volume=132&issue=
1-2&aid=1691.

Lescanne:1994:TOR

[1107] Pierre Lescanne. On termination of one rule rewrite systems. Theoretical Computer Science, 132(1–2):395– 401, September 26, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.
cgi?year=1994&volume=132&issue=
1-2&aid=1670.

Crochemore:1994:TDP

[1108] Maxime Crochemore and Wojciech Rytter. On two-dimensional pattern matching by optimal parallel algorithms. Theoretical Computer Science, 132(1–2):403–414, September 26, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=132&issue=1-2&aid=1653.

Berstel:1994:ACD

[1109] J. Berstel and M. Pocchiola. Average cost of Duval’s algorithm for generating Lyndon words. Theoretical Computer Science, 132(1–2):415–425, September 26, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/

                                        cas_sub/browse/browse.cgi?year=
                                        1994&volume=132&issue=1-2&aid=1654.

Ilie:1994:CAS

[1110] Lucian Ilie. On a conjecture about slender context-free languages. Theoretical Computer Science, 132(1–2): 427–434, September 26, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/tcs/cas_sub/browse/browse.
                                        cgi?year=1994&volume=132&issue=
                                        1-2&aid=1683.

Duris:1994:NDO

[1111] Pavol Ďuriš and José D. P. Rolim. A note on the density of oracle decreasing time-space complexity. Theoretical Computer Science, 132(1–2):435–444, September 26, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1994&volume=132&issue=1-2&aid=1713.

Anonymous:1994:WCA

[1112] Anonymous. Workshop on Continuous Algorithms and Complexity. Theoretical Computer Science, 133(1):??, October 11, 1994. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cucker:1994:SCC

[1113] F. Cucker, M. Shub, and S. Smale. Separation of complexity classes in Koiran’s weak model. Theoretical Computer Science, 133(1):3–14, October 11, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.

com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=133&issue=1&aid=1710.

Emerson:1994:RPN

[1114] T. Emerson. Relativizations of the P =? NP question over the reals (and other ordered rings). Theoretical Computer Science, 133(1):15–22, October 11, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=133&issue=1&aid=1709.

Grigoriev:1994:DTS

[1115] D. Yu. Grigoriev. Deviation theorems for solutions of differential equations and applications to lower bounds on parallel complexity of sigmoids. Theoretical Computer Science, 133(1):23–33, October 11, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=133&issue=1&aid=1702.

Koiran:1994:CRA

[1116] P. Koiran. Computing over the reals with addition and order (complexity theory). Theoretical Computer Science, 133(1):35–47, October 11, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=133&issue=1&aid=1704.

Kuorka:1994:RUS

[1117] Petr Kůrka. Regular unimodal systems and factors of finite automata. Theoretical Computer Science, 133(1):49–64, October 11, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1994&volume=133&issue=1&aid=1707.

Malajovich:1994:GNA

[1118] Gregorio Malajovich. On generalized Newton algorithms: quadratic convergence, path-following and error analysis. Theoretical Computer Science, 133(1):65–84, October 11, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1994&volume=133&issue=1&aid=1706.

Meer:1994:CQP

[1119] K. Meer. On the complexity of quadratic programming in real number models of computation. Theoretical Computer Science, 133(1):85–94, October 11, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1994&volume=133&issue=1&aid=1711.

Michaux:1994:PNN

[1120] Christian Michaux. P≠NP over the nonstandard reals implies P≠NP over R. Theoretical Computer Science, 133(1):95–104, October 11, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=133&issue=1&aid=1708.

Rojas:1994:CGA

[1121] J. Maurice Rojas. A convex geometric approach to counting the roots of a polynomial system. Theoretical Computer Science, 133(1):105–140, October 11, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=133&issue=1&aid=1703.

Shub:1994:CBT

[1122] M. Shub and S. Smale. Complexity of Bezout’s theorem V: Polynomial time. Theoretical Computer Science, 133(1):141–164, October 11, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=133&issue=1&aid=1701.

Verschelde:1994:HSP

[1123] Jan Verschelde and Ann Haegemans. Homotopies for solving polynomial systems within a bounded domain. Theoretical Computer Science, 133(1):165–185, October 11, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=133&issue=1&aid=1705.

Bonner:1994:OTL

[1124] Anthony J. Bonner and Michael Kifer. An overview of transaction logic. Theoretical Computer Science, 133(2):205–265, October 24, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1994&volume=133&issue=2&aid=1694.

Dong:1994:IID

[1125] Fangqing Dong and Laks V. S. Lakshmanan. Intuitionistic interpretation of deductive databases with incomplete information. Theoretical Computer Science, 133(2):267–306, October 24, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1994&volume=133&issue=2&aid=1695.

Kapur:1994:OTP

[1126] D. Kapur, X. Nie, and D. R. Musser. An overview of the Tecton proof system. Theoretical Computer Science, 133(2):307–340, October 24, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1994&volume=133&issue=2&aid=1712.

Ramanathan:1994:RED

[1127] Geetha Ramanathan. Refinement of events in the development of real-time distributed systems. Theoretical Computer Science, 133(2):341–359, October 24, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=133&issue=2&aid=1696.

Han:1994:TEI

[1128] Jiawei Han. Towards efficient induction mechanisms in database systems. Theoretical Computer Science, 133(2):361–385, October 24, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=133&issue=2&aid=1697.

Godin:1994:ICF

[1129] Robert Godin and Rokia Missaoui.

An incremental concept formation approach for learning from databases. Theoretical Computer Science, 133(2): 387–419, October 24, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.
cgi?year=1994&volume=133&issue=
2&aid=1698.

Sadri:1994:AOI

[1130] Fereidoon Sadri. Aggregate operations in the information source tracking method. Theoretical Computer Science, 133(2):421–442, October 24, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=133&issue=2&aid=1699.

Anonymous:1994:SIC

[1131] Anonymous. Second International Colloquium on Words, Languages and Combinatorics. Theoretical Computer Science, 134(1):??, November 07, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anderson:1994:SFM

[1132] J. A. Anderson. Semiretracts of a free monoid. Theoretical Computer Science, 134(1):3–11, November 07, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1994&volume=134&issue=1&aid=1733.

Beauquier:1994:RTA

[1133] Danièle Beauquier and Andreas Podelski. Rabin tree automata and finite monoids. Theoretical Computer Science, 134(1):13–25, November 07, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1994&volume=134&issue=1&aid=1643.

Bonnier-Rigny:1994:CSI

[1134] Agnès Bonnier-Rigny and Daniel Krob. A complete system of identities for one-letter rational expressions with multiplicities in the tropical semiring. Theoretical Computer Science, 134(1):27–50, November 07, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/

cas_sub/browse/browse.cgi?year=
1994&volume=134&issue=1&aid=1756.

Botelho:1994:BNN

[1135] Fernando Botelho and Max Garzon. Boolean neural nets are observable. Theoretical Computer Science, 134(1):51–61, November 07, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=134&issue=1&aid=1731.

See erratum [1201].

Capocelli:1994:FAU

[1136] Renato M. Capocelli, Luisa Gargano, and Ugo Vaccaro. A fast algorithm for the unique decipherability of multivalued encodings. Theoretical Computer Science, 134(1):63–78, November 07, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=134&issue=1&aid=1644.

Crvenkovic:1994:DA

[1137] Siniša Crvenković and Rozália Sz Madarász. On dynamic algebras. Theoretical Computer Science, 134(1):79–86, November 07, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=134&issue=1&aid=1645.

Diekert:1994:PTS

[1138] V. Diekert. A partial trace semantics for Petri nets. Theoretical Computer Science, 134(1):87–105, November 07, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=134&issue=1&aid=1646.

Jurgensen:1994:TDI

[1139] H. Jürgensen, K. Salomaa, and S. Yu. Transducers and the decidability of independence in free monoids. Theoretical Computer Science, 134(1):107–117, November 07, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=134&issue=1&aid=1679.

Kelemenova:1994:LC

[1140] Alica Kelemenová and Erzsébet Csuhaj-Varjú. Languages of colonies. Theoretical Computer Science, 134(1): 119–130, November 07, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.
cgi?year=1994&volume=134&issue=
1&aid=1647.

Mora:1994:ICN

[1141] Teo Mora. An introduction to commutative and non-commutative Gröbner bases. Theoretical Computer Science, 134(1):131–173, November 07, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=134&issue=1&aid=1735.

Otto:1994:CMF

[1142] Friedrich Otto and Paliath Narendran.

Codes modulo finite monadic stringrewriting systems. Theoretical Computer Science, 134(1):175–188, November 07, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=134&issue=1&aid=1734.

Reilly:1994:BVG

[1143] Norman R. Reilly. Bounds on the variety generated by completely regular syntactic monoids from finite prefix codes. Theoretical Computer Science, 134(1):189–208, November 07, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=134&issue=1&aid=1648.

Spehner:1994:BBC

[1144] J.-C. Spehner. A bijection between cliques in graphs and factorizations in free monoids. Theoretical Computer Science, 134(1):209–223, November 07, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=134&issue=1&aid=1732.

Weber:1994:FVD

[1145] Andreas Weber. Finite-valued distance automata. Theoretical Computer Science, 134(1):225–251, November 07, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1994&volume=134&issue=1&aid=1649.

Weber:1994:EUL

[1146] Andreas Weber. Exponential upper and lower bounds for the order of a regular language. Theoretical Computer Science, 134(1):253–262, November 07, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1994&volume=134&issue=1&aid=1650.

Kobler:1994:LPP

[1147] J. Köbler. Locating P/poly optimally in the extended low hierarchy. Theoretical Computer Science, 134(2):263–285, November 21, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1994&volume=134&issue=2&aid=1657.

Chan:1994:TSC

[1148] Edward P. F. Chan. Testing satisfiability of a class of object-oriented conjunctive queries. Theoretical Computer Science, 134(2):287–309, November 21, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1994&volume=134&issue=2&aid=1656.

Fulop:1994:UPD

[1149] Z. Fülöp. Undecidable properties of deterministic top-down tree transducers. Theoretical Computer Science, 134(2):311–328, November 21, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=134&issue=2&aid=1637.

Kaminski:1994:FMA

[1150] Michael Kaminski and Nissim Francez.

Finite-memory automata. Theoretical Computer Science, 134(2):329– 363, November 21, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.
cgi?year=1994&volume=134&issue=
2&aid=1631.

Tiplea:1994:SDP

[1151] Ferucio Laurentiu Tiplea, Cristian Ene, Cecilia Magdalena Ionescu, and Octavian Procopiuc. Some decision problems for parallel communicating grammar systems. Theoretical Computer Science, 134(2):365–385, November 21, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=134&issue=2&aid=1636.

Durand:1994:ICA

[1152] B. Durand. Inversion of 2D cellular automata: some complexity results. Theoretical Computer Science, 134(2):387–401, November 21, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/

                                        cas_sub/browse/browse.cgi?year=
                                        1994&volume=134&issue=2&aid=1618.

Harju:1994:RRF

[1153] T. Harju, H. C. M. Kleijn, M. Latteux, and A. Terlutte. Representation of rational functions with prefix and suffix codings. Theoretical Computer Science, 134(2):403–413, November 21, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1994&volume=134&issue=2&aid=1625.

Remila:1994:TTT

[1154] Eric Remila. On the tiling of a torus with two bars. Theoretical Computer Science, 134(2):415–426, November 21, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1994&volume=134&issue=2&aid=1628.

Mak:1994:SDA

[1155] Louis Mak. Speedup of determinism by alternation for multidimensional Turing machines. Theoretical Computer Science, 134(2):427–453, November 21, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1994&volume=134&issue=2&aid=1635.

Matus:1994:SIA

[1156] František Matúš. Stochastic independence, algebraic independence and abstract connectedness. Theoretical Computer Science, 134(2):455–471, November 21, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=134&issue=2&aid=1624.

Jiang:1994:ASS

[1157] Tao Jiang and Ming Li. Approximating shortest superstrings with constraints. Theoretical Computer Science, 134(2):473–491, November 21, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=134&issue=2&aid=1627.

Dahlhaus:1994:EPA

[1158] Elias Dahlhaus and Marek Karpinski.

An efficient parallel algorithm for the minimal elimination ordering (MEO) of an arbitrary graph. Theoretical Computer Science, 134(2):493–528, November 21, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=134&issue=2&aid=1642.

Alonso:1994:UGM

[1159] Laurent Alonso. Uniform generation of a Motzkin word. Theoretical Computer Science, 134(2):529–536, November 21, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=134&issue=2&aid=1727.

Harrison:1994:MCD

[1160] John Harrison. Morphic congruences and D0L languages. Theoretical Computer Science, 134(2):537–544, November 21, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1994&volume=134&issue=2&aid=1751.

Fortnow:1994:PMP

[1161] Lance Fortnow, John Rompel, and Michael Sipser. On the power of multi-prover interactive protocols. Theoretical Computer Science, 134(2):545–557, November 21, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1994&volume=134&issue=2&aid=1718.

Gu:1994:AOH

[1162] Xunrang Gu and Yuzhang Zhu. Asymptotic optimal HEAPSORT algorithm. Theoretical Computer Science, 134(2):559–565, November 21, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1994&volume=134&issue=2&aid=1754.

Anonymous:1994:MFP

[1163] Anonymous. Mathematical foundations of programming semantics. Theoretical Computer Science, 135(1):??, December 05, 1994. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Abramsky:1994:PP

[1164] Samson Abramsky. Proofs as processes. Theoretical Computer Science, 135(1):5–9, December 05, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=135&issue=1&aid=1744.

Bellin:1994:CLL

[1165] G. Bellin and P. J. Scott. On the π-calculus and linear logic. Theoretical Computer Science, 135(1):11– 65, December 05, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.
cgi?year=1994&volume=135&issue=
1&aid=1745.

Galmiche:1994:PNL

[1166] Didier Galmiche and Guy Perrier.

On proof normalization in linear logic. Theoretical Computer Science, 135(1):67–110, December 05, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=135&issue=1&aid=1746.

See corrigendum [1202].

Mascari:1994:HLR

[1167] G. F. Mascari and M. Pedicini. Head linear reduction and pure proof net extraction. Theoretical Computer Science, 135(1):111–137, December 05, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=135&issue=1&aid=1747.

Lincoln:1994:FOL

[1168] P. Lincoln and A. Scedrov. Firstorder linear logic without modalities is NEXPTIME-hard. Theoretical Computer Science, 135(1):139–153, December 05, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=135&issue=1&aid=1748.

Lincoln:1994:COM

[1169] Patrick Lincoln and Timothy Winkler. Constant-only multiplicative linear logic is NP-complete. Theoretical Computer Science, 135(1):155–169, December 05, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=135&issue=1&aid=1749.

Baier:1994:DSC

[1170] Christel Baier and Mila E. Majster-Cederbaum. Denotational semantics in the cpo and metric approach. Theoretical Computer Science, 135(2):171–220, December 12, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=135&issue=2&aid=1687.

Ehrig:1994:FTP

[1171] Hartmut Ehrig and Martin Große-Rhode. Functorial theory of parameterized specifications in a general specification framework. Theoretical Computer Science, 135(2):221–266, December 12, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=135&issue=2&aid=1620.

Lastovetsky:1994:AAS

[1172] Alexey L. Lastovetsky and Sergey S.

Gaissaryan. An algebraic approach to semantics of programming languages. Theoretical Computer Science, 135(2):267–288, December 12, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=135&issue=2&aid=1663.

Bracho:1994:LDA

[1173] Felipe Bracho and Manfred Droste.

Labelled domains and automata with concurrency. Theoretical Computer Science, 135(2):289–318, December 12, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=135&issue=2&aid=1671.

Manoury:1994:ATP

[1174] Pascal Manoury and Marianne Simonot. Automatizing termination proofs of recursively defined functions. Theoretical Computer Science, 135(2):319–343, December 12, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=135&issue=2&aid=1662.

Scott:1994:WTD

[1175] E. A. Scott. Weights for total division orderings on strings. Theoretical Computer Science, 135(2):345–359, December 12, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=135&issue=2&aid=1666.

Hiraishi:1994:SCR

[1176] Kunihiko Hiraishi. Some complexity results on transition systems and elementary net systems. Theoretical Computer Science, 135(2):361–376, December 12, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=135&issue=2&aid=1633.

Li:1994:FPN

[1177] Jianan Li, Ichiro Suzuki, and Masafumi Yamashita. Fair Petri nets and structural induction for rings of processes. Theoretical Computer Science, 135(2):377–404, December 12, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=135&issue=2&aid=1630.

Trigg:1994:CAU

[1178] Peter Trigg, J. Roger Hindley, and Martin W. Bunder. Combinatory abstraction using B, B′ and friends. Theoretical Computer Science, 135(2):405–422, December 12, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=135&issue=2&aid=1610.

David:1994:IFS

[1179] R. David. The Inf function in the system F. Theoretical Computer Science, 135(2):423–431, December 12, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=135&issue=2&aid=1729.

Johnstone:1994:VBT

[1180] P. T. Johnstone. Variations on the bagdomain theme. Theoretical Computer Science, 136(1):3–20, December 19, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=136&issue=1&aid=1761.

Heckmann:1994:SPD

[1181] Reinhold Heckmann. Stable power domains. Theoretical Computer Science, 136(1):21–56, December 19, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/

                                        cas_sub/browse/browse.cgi?year=
                                        1994&volume=136&issue=1&aid=1762.

Fiech:1994:ADN

[1182] Adrian Fiech and Michael Huth. Algebraic domains of natural transformations. Theoretical Computer Science, 136(1):57–78, December 19, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1994&volume=136&issue=1&aid=1763.

Melton:1994:LCC

[1183] Austin Melton, Bernd S. W. Schröder, and George E. Strecker. Lagois connections — a counterpart to Galois connections. Theoretical Computer Science, 136(1):79–107, December 19, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1994&volume=136&issue=1&aid=1764.

Mulry:1994:PMC

[1184] Philip S. Mulry. Partial map classifiers and partial Cartesian closed categories. Theoretical Computer Science, 136(1):109–123, December 19, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1994&volume=136&issue=1&aid=1765.

Ritter:1994:CAM

[1185] Eike Ritter. Categorical abstract machines for higher-order typed lambda — calculi. Theoretical Computer Science, 136(1):125–162, December 19, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=136&issue=1&aid=1766.

Robinson:1994:PI

[1186] Edmund Robinson. Parametricity as isomorphism. Theoretical Computer Science, 136(1):163–181, December 19, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=136&issue=1&aid=1767.

Kamareddine:1994:UAT

[1187] Fairouz Kamareddine and Rob Nederpelt. A unified approach to type theory through a refined lambda — calculus. Theoretical Computer Science, 136(1):183–216, December 19, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=136&issue=1&aid=1768.

Crole:1994:CAF

[1188] Roy L. Crole. Computational adequacy of the FIX-logic. Theoretical Computer Science, 136(1):217–242, December 19, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=136&issue=1&aid=1769.

Blaaberg:1994:ALF

[1189] Bettina Blaaberg and Christian Clausen. Adequacy for a lazy functional language with recursive and polymorphic types. Theoretical Computer Science, 136(1):243–275, December 19, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1994&volume=136&issue=1&aid=1770.

Wolfram:1994:SP

[1190] D. A. Wolfram. A semantics for λProlog. Theoretical Computer Science, 136(1):277–289, December 19, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1994&volume=136&issue=1&aid=1771.

Remila:1994:RGA

[1191] Eric Remila. Recognition of graphs by automata. Theoretical Computer Science, 136(2):291–332, December 29, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1994&volume=136&issue=2&aid=1686.

Ohlebusch:1994:MTT

[1192] Enno Ohlebusch. On the modularity of termination of term rewriting systems. Theoretical Computer Science, 136(2):333–360, December 29, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.

com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=136&issue=2&aid=1680.

deLuca:1994:SCP

[1193] Aldo de Luca and Filippo Mignosi.

Some combinatorial properties of Sturmian words. Theoretical Computer Science, 136(2):361–385, December 29, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=136&issue=2&aid=1676.

Senoussi:1994:QAT

[1194] H. Senoussi and A. Saoudi. A quadtree algorithm for template matching on a pyramid computer. Theoretical Computer Science, 136(2):387–417, December 29, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=136&issue=2&aid=1661.

Stetsenko:1994:ABB

[1195] Vladimir Stetsenko. On almost bad Boolean bases. Theoretical Computer Science, 136(2):419–469, December 29, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=136&issue=2&aid=1689.

Ito:1994:CIC

[1196] M. Ito and G. Thierrin. Congruences, infix and cohesive prefix codes. Theoretical Computer Science, 136(2):471–485, December 29, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=136&issue=2&aid=1659.

Mayordomo:1994:AES

[1197] Elvira Mayordomo. Almost every set in exponential time is P-bi-immune. Theoretical Computer Science, 136(2):487–506, December 29, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=136&issue=2&aid=1664.

Chrobak:1994:TRL

[1198] Marek Chrobak and Wojciech Rytter. Two results on linear embeddings of complete binary trees. Theoretical Computer Science, 136(2):507– 526, December 29, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.
cgi?year=1994&volume=136&issue=
2&aid=1677.

Kiwi:1994:NPB

[1199] M. A. Kiwi, R. Ndoundam, M. Tchuente, and E. Goles. No polynomial bound for the period of the parallel chip firing game on graphs. Theoretical Computer Science, 136(2):527–532, December 29, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=136&issue=2&aid=1772.

Petersen:1995:RPB

[1200] H. Petersen. A remark on a paper by A. B. Matos. Theoretical Computer Science, 141(1–2):329–330, April 17, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=141&issue=1-2&aid=1851.

See [1018].

Botelho:1997:EBN

[1201] Fernanda Botelho and Max Garzon.

Erratum to “Boolean neural nets are observable” [Theoret. Comput. Sci. 134 (1994) 51–61]. Theoretical Computer Science, 182(1–2):257, August 15, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=182&issue=1-2&aid=2649.

See [1135].

Perrier:1998:CGP

[1202] Guy Perrier. Corrigendum to Galmiche’s and Perrier’s “On proof normalization in linear logic” [Theoret. Comput. Sci 135(1) (1994) 67– 110]. Theoretical Computer Science, 206(1–2):353–354, October 06, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1998/
206/1-2/2965.pdf. See [1166].

DeAgostino:2000:EPC

[1203] Sergio De Agostino. Erratum to “P-complete Problems in Data Compression” [Theoret. Comput. Sci.

127 (1994) 181–186]. Theoretical Computer Science, 234(1–2):325– 326, March 6, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        170/21/37/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/170/21/37/article.pdf.  See

[1012].